the GNU GRUB manual

The GRand Unified Bootloader, version 2.12, 5 December 2023.

Gordon Matzigkeit Yoshinori K. Okuji Colin Watson Colin D. Bennett

This manual is for GNU GRUB (version 2.12, 5 December 2023). Copyright \begin{matrix}c \\ ◯ \end{matrix} 1999,2000,2001,2002,2004,2006,2008,2009,2010,2011,2012,2013 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.2 or any later version published by the Free Software Foundation; with no Invariant Sections.

Table of Contents

1 Introduction to GRUB . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 History of GRUB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Differences from previous versions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.4 GRUB features. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.5 The role of a boot loader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

2 Naming convention . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

3 OS-specific notes about grub tools . . . . . . . . . . . . 9

4 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4.1 Installing GRUB using grub-install . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.2 Making a GRUB bootable CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.3 The map between BIOS drives and OS devices . . . . . . . . . . . . . . . . 13 4.4 BIOS installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

5 Booting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

5.1 How to boot operating systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5.1.1 How to boot an OS directly with GRUB . . . . . . . . . . . . . . . . . . 15 5.1.2 Kexec with grub2-emu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5.1.3 Chain-loading an OS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

5.2 Loopback booting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 5.3 Booting from LVM cache logical volume . . . . . . . . . . . . . . . . . . . . . . . 16 5.4 Some caveats on OS-specific issues . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

5.4.1 GNU/Hurd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 5.4.2 GNU/Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 5.4.3 NetBSD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 5.4.4 DOS/Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

6 Writing your own configuration file . . . . . . . . . . 21

6.1 Simple configuration handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 6.2 Root Identifcation Heuristics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 6.3 Writing full configuration files directly . . . . . . . . . . . . . . . . . . . . . . . . . 28 6.4 Multi-boot manual config . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 6.5 Embedding a configuration file into GRUB . . . . . . . . . . . . . . . . . . . . 33

7 Theme file format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

7.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 7.2 Theme Elements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

7.2.1 Colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 7.2.2 Fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 7.2.3 Progress Bar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 7.2.4 Circular Progress Indicator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 7.2.5 Labels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 7.2.6 Boot Menu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 7.2.7 Styled Boxes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 7.2.8 Creating Styled Box Images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

7.3 Theme File Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 7.3.1 Global Properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 7.3.2 Format. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 7.3.3 Global Property List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 7.3.4 Component Construction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 7.3.5 Component List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 7.3.6 Common properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

8 Booting GRUB from the network. . . . . . . . . . . . 45

9 Using GRUB via a serial line. . . . . . . . . . . . . . . . . 49

10 Using GRUB with vendor power-on keys . . 51

11 GRUB image files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

12 Core image size limitation . . . . . . . . . . . . . . . . . . . 55

13 Filesystem syntax and semantics. . . . . . . . . . . . 57

13.1 How to specify devices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 13.2 How to specify files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 13.3 How to specify block lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

14 GRUB’s user interface . . . . . . . . . . . . . . . . . . . . . . . 61

14.1 The flexible command-line interface . . . . . . . . . . . . . . . . . . . . . . . . . . 61 14.2 The simple menu interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 14.3 Editing a menu entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

15 GRUB environment variables . . . . . . . . . . . . . . . 63

15.1 Special environment variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 15.1.1 biosnum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 15.1.2 check signatures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 15.1.3 chosen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 15.1.4 cmdpath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 15.1.5 color highlight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 15.1.6 color normal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 15.1.7 config directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 15.1.8 config file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 15.1.9 debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 15.1.10 default. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 15.1.11 fallback . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 15.1.12 gfxmode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 15.1.13 gfxpayload . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 15.1.14 gfxterm font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 15.1.15 grub cpu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 15.1.16 grub platform. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 15.1.17 icondir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 15.1.18 lang . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 15.1.19 locale dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 15.1.20 lockdown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 15.1.21 menu color highlight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 15.1.22 menu color normal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 15.1.23 net <interface> boot file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 15.1.24 net <interface> clientid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 15.1.25 net <interface> clientuuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 15.1.26 net <interface> dhcp server name . . . . . . . . . . . . . . . . . . . . . . 68 15.1.27 net <interface> domain . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 15.1.28 net <interface> extensionspath . . . . . . . . . . . . . . . . . . . . . . . . . 68 15.1.29 net <interface> hostname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 15.1.30 net <interface> ip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 15.1.31 net <interface> mac. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 15.1.32 net <interface> next server. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 15.1.33 net <interface> rootpath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 15.1.34 net default interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 15.1.35 net default ip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 15.1.36 net default mac. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 15.1.37 net default server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 15.1.38 pager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 15.1.39 prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 15.1.40 pxe blksize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 15.1.41 pxe default gateway . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 15.1.42 pxe default server. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 15.1.43 root . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 15.1.44 shim lock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 15.1.45 superusers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 15.1.46 theme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 iv GNU GRUB Manual 2.12

15.1.47 timeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 15.1.48 timeout style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 15.1.49 tpm fail fatal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

15.2 The GRUB environment block . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

16 Modules. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

16.1 bli . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

17 Available commands . . . . . . . . . . . . . . . . . . . . . . . . . 73

17.1 Commands for the menu only . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 17.1.1 menuentry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 17.1.2 submenu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

17.2 Various loader commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 17.2.1 chainloader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 17.2.2 initrd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 17.2.3 initrd16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 17.2.4 linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 17.2.5 linux16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 17.2.6 xen hypervisor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 17.2.7 xen module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

17.3 General commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 17.3.1 serial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 17.3.2 terminal input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 17.3.3 terminal output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 17.3.4 terminfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

17.4 Command-line commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 17.4.1 [ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 17.4.2 acpi. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 17.4.3 authenticate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 17.4.4 background color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 17.4.5 background image . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 17.4.6 badram . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 17.4.7 blocklist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 17.4.8 boot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 17.4.9 cat. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 17.4.10 clear . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 17.4.11 cmosclean. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 17.4.12 cmosdump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 17.4.13 cmostest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 17.4.14 cmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 17.4.15 configfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 17.4.16 cpuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 17.4.17 crc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 17.4.18 cryptomount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 17.4.19 cutmem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 17.4.20 date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 17.4.21 devicetree. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 17.4.22 distrust . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 17.4.23 drivemap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 17.4.24 echo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 17.4.25 efitextmode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 17.4.26 eval. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 17.4.27 export . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 17.4.28 false . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 17.4.29 fwsetup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 17.4.30 gdbinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 17.4.31 gettext. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 17.4.32 gptsync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 17.4.33 halt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 17.4.34 hashsum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 17.4.35 help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 17.4.36 hexdump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 17.4.37 insmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 17.4.38 keystatus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 17.4.39 list env . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 17.4.40 list trusted . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 17.4.41 load env . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 17.4.42 loadfont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 17.4.43 loopback . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 17.4.44 ls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 17.4.45 lsfonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 17.4.46 lsmod. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 17.4.47 md5sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 17.4.48 module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 17.4.49 multiboot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 17.4.50 nativedisk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 17.4.51 normal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 17.4.52 normal exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 17.4.53 parttool. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 17.4.54 password . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 17.4.55 password pbkdf2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 17.4.56 plainmount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 17.4.57 play . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 17.4.58 probe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 17.4.59 rdmsr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 17.4.60 read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 17.4.61 reboot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 17.4.62 regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 17.4.63 rmmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 17.4.64 save env . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 17.4.65 search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 17.4.66 sendkey . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 17.4.67 set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 17.4.68 sha1sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 17.4.69 sha256sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 17.4.70 sha512sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 17.4.71 sleep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 17.4.72 smbios . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 17.4.73 source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 17.4.74 test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 17.4.75 true . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 17.4.76 trust. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 17.4.77 unset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 17.4.78 verify detached . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 17.4.79 videoinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 17.4.80 wrmsr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

17.5 Networking commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 17.5.1 net add addr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 17.5.2 net add dns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 17.5.3 net add route. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 17.5.4 net bootp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 17.5.5 net del addr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 17.5.6 net del dns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 17.5.7 net del route . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 17.5.8 net dhcp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 17.5.9 net get dhcp option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 17.5.10 net ipv6 autoconf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 17.5.11 net ls addr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 17.5.12 net ls cards. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 17.5.13 net ls dns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 17.5.14 net ls routes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 17.5.15 net nslookup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 17.5.16 net set vlan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

17.6 Commands currently undocumented. . . . . . . . . . . . . . . . . . . . . . . . . 104

18 Internationalisation . . . . . . . . . . . . . . . . . . . . . . . . . 107

18.1 Charset. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 18.2 Filesystems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 18.3 Output terminal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 18.4 Input terminal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 18.5 Gettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 18.6 Regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 18.7 Other. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108

19 Security. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

19.1 Authentication and authorisation in GRUB . . . . . . . . . . . . . . . . . 111 19.2 Using digital signatures in GRUB . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 19.3 UEFI secure boot and shim support. . . . . . . . . . . . . . . . . . . . . . . . . 113 19.4 Embedded information for generation number based revocation

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 19.5 Measuring boot components . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 19.6 Lockdown when booting on a secure setup . . . . . . . . . . . . . . . . . . 114

20 Platform limitations . . . . . . . . . . . . . . . . . . . . . . . . 117

21 Outline. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

22 Supported boot targets. . . . . . . . . . . . . . . . . . . . . 123

22.1 Boot tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

23 Error messages produced by GRUB. . . . . . . 127

23.1 GRUB only offers a rescue shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 23.2 Firmware stalls instead of booting GRUB . . . . . . . . . . . . . . . . . . . 128

24 User-space utilities. . . . . . . . . . . . . . . . . . . . . . . . . . 129

24.1 Invoking grub-install . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 24.2 Invoking grub-mkconfig . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 24.3 Invoking grub-mkpasswd-pbkdf2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 24.4 Invoking grub-mkrelpath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 24.5 Invoking grub-mkrescue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 24.6 Invoking grub-mount. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 24.7 Invoking grub-probe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 24.8 Invoking grub-script-check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134

Appendix A How to obtain and build GRUB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

Appendix B Reporting bugs . . . . . . . . . . . . . . . . . . 137

Appendix C Where GRUB will go . . . . . . . . . . . 139

Appendix D Copying This Manual . . . . . . . . . . . 141

D.1 GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

D.1.1 ADDENDUM: How to use this License for your documents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147

Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149

1 Introduction to GRUB

1.1 Overview

Briefly, a boot loader is the first software program that runs when a computer starts. It is responsible for loading and transferring control to an operating system kernel software (such as Linux or GNU Mach). The kernel, in turn, initializes the rest of the operating system (e.g. a GNU system).

GNU GRUB is a very powerful boot loader, which can load a wide variety of free operating systems, as well as proprietary operating systems with chain-loading^{1}. GRUB is designed to address the complexity of booting a personal computer; both the program and this manual are tightly bound to that computer platform, although porting to other platforms may be addressed in the future.

One of the important features in GRUB is flexibility; GRUB understands filesystems and kernel executable formats, so you can load an arbitrary operating system the way you like, without recording the physical position of your kernel on the disk. Thus you can load the kernel just by specifying its file name and the drive and partition where the kernel resides.

When booting with GRUB, you can use either a command-line interface (see Section 14.1 [Command-line interface], page 61), or a menu interface (see Section 14.2 [Menu interface], page 62). Using the command-line interface, you type the drive specification and file name of the kernel manually. In the menu interface, you just select an OS using the arrow keys. The menu is based on a configuration file which you prepare beforehand (see Chapter 6 [Configuration], page 21). While in the menu, you can switch to the commandline mode, and vice-versa. You can even edit menu entries before using them.

In the following chapters, you will learn how to specify a drive, a partition, and a file name (see Chapter 2 [Naming convention], page 7) to GRUB, how to install GRUB on your drive (see Chapter 4 [Installation], page 11), and how to boot your OSes (see Chapter 5 [Booting], page 15), step by step.

1.2 History of GRUB

GRUB originated in 1995 when Erich Boleyn was trying to boot the GNU Hurd with the University of Utah’s Mach 4 microkernel (now known as GNU Mach). Erich and Brian Ford designed the Multiboot Specification (see Section “Motivation” in The Multiboot Specification), because they were determined not to add to the large number of mutuallyincompatible PC boot methods.

Erich then began modifying the FreeBSD boot loader so that it would understand Multiboot. He soon realized that it would be a lot easier to write his own boot loader from scratch than to keep working on the FreeBSD boot loader, and so GRUB was born.

Erich added many features to GRUB, but other priorities prevented him from keeping up with the demands of its quickly-expanding user base. In 1999, Gordon Matzigkeit and Yoshinori K. Okuji adopted GRUB as an official GNU package, and opened its development by making the latest sources available via anonymous CVS. See Appendix A [Obtaining and Building GRUB], page 135, for more information.

1 chain-load is the mechanism for loading unsupported operating systems by loading another boot

loader. It is typically used for loading DOS or Windows.

Over the next few years, GRUB was extended to meet many needs, but it quickly became clear that its design was not keeping up with the extensions being made to it, and we reached the point where it was very difficult to make any further changes without breaking existing features. Around 2002, Yoshinori K. Okuji started work on PUPA (Preliminary Universal Programming Architecture for GNU GRUB), aiming to rewrite the core of GRUB to make it cleaner, safer, more robust, and more powerful. PUPA was eventually renamed to GRUB 2, and the original version of GRUB was renamed to GRUB Legacy. Small amounts of maintenance continued to be done on GRUB Legacy, but the last release (0.97) was made in 2005 and at the time of writing it seems unlikely that there will be another.

By around 2007, GNU/Linux distributions started to use GRUB 2 to limited extents, and by the end of 2009 multiple major distributions were installing it by default.

1.3 Differences from previous versions

GRUB 2 is a rewrite of GRUB (see Section 1.2 [History], page 1), although it shares many characteristics with the previous version, now known as GRUB Legacy. Users of GRUB Legacy may need some guidance to find their way around this new version.

• The configuration file has a new name (‘grub.cfg’ rather than ‘menu.lst’ or ‘grub.conf’), new syntax (see Chapter 6 [Configuration], page 21) and many new commands (see Chapter 17 [Commands], page 73). Configuration cannot be copied over directly, although most GRUB Legacy users should not find the syntax too surprising.

• ‘grub.cfg’ is typically automatically generated by grub-mkconfig (see Section 6.1 [Simple configuration], page 21). This makes it easier to handle versioned kernel upgrades.

• Partition numbers in GRUB device names now start at 1, not 0 (see Chapter 2 [Naming convention], page 7).

• The configuration file is now written in something closer to a full scripting language: variables, conditionals, and loops are available.

• A small amount of persistent storage is available across reboots, using the save_env and load_env commands in GRUB and the grub-editenv utility. This is not available in all configurations (see Section 15.2 [Environment block], page 70).

• GRUB 2 has more reliable ways to find its own files and those of target kernels on multiple-disk systems, and has commands (see Section 17.4.65 [search], page 93) to find devices using file system labels or Universally Unique Identifiers (UUIDs).

• GRUB 2 is available for several other types of system in addition to the PC BIOS systems supported by GRUB Legacy: PC EFI, PC coreboot, PowerPC, SPARC, and MIPS Lemote Yeeloong are all supported.

• Many more file systems are supported, including but not limited to ext4, HFS+, and NTFS.

• GRUB 2 can read files directly from LVM and RAID devices.

• A graphical terminal and a graphical menu system are available.

• GRUB 2’s interface can be translated, including menu entry names.

• The image files (see Chapter 11 [Images], page 53) that make up GRUB have been reorganised; Stage 1, Stage 1.5, and Stage 2 are no more.

• GRUB 2 puts many facilities in dynamically loaded modules, allowing the core image to be smaller, and allowing the core image to be built in more flexible ways.

1.4 GRUB features

The primary requirement for GRUB is that it be compliant with the Multiboot Specification, which is described in Section “Motivation” in The Multiboot Specification.

The other goals, listed in approximate order of importance, are:

• Basic functions must be straightforward for end-users.

• Rich functionality to support kernel experts and designers.

• Backward compatibility for booting FreeBSD, NetBSD, OpenBSD, and Linux. Proprietary kernels (such as DOS, Windows NT, and OS/2) are supported via a chain-loading function.

Except for specific compatibility modes (chain-loading and the Linux piggyback format), all kernels will be started in much the same state as in the Multiboot Specification. Only kernels loaded at 1 megabyte or above are presently supported. Any attempt to load below that boundary will simply result in immediate failure and an error message reporting the problem.

In addition to the requirements above, GRUB has the following features (note that the Multiboot Specification doesn’t require all the features that GRUB supports):

Recognize multiple executable formats Support many of the a.out variants plus ELF. Symbol tables are also loaded.

Support non-Multiboot kernels Support many of the various free 32-bit kernels that lack Multiboot compliance (primarily FreeBSD, NetBSD^{2}, OpenBSD, and Linux). Chain-loading of other boot loaders is also supported.

Load multiples modules Fully support the Multiboot feature of loading multiple modules.

Load a configuration file Support a human-readable text configuration file with preset boot commands. You can also load another configuration file dynamically and embed a preset configuration file in a GRUB image file. The list of commands (see Chapter 17 [Commands], page 73) are a superset of those supported on the command-line. An example configuration file is provided in Chapter 6 [Configuration], page 21.

Provide a menu interface A menu interface listing preset boot commands, with a programmable timeout, is available. There is no fixed limit on the number of boot entries, and the current implementation has space for several hundred.

2 The NetBSD/i386 kernel is Multiboot-compliant, but lacks support for Multiboot modules.

Have a flexible command-line interface A fairly flexible command-line interface, accessible from the menu, is available to edit any preset commands, or write a new boot command set from scratch. If no configuration file is present, GRUB drops to the command-line. The list of commands (see Chapter 17 [Commands], page 73) are a subset of those supported for configuration files. Editing commands closely resembles the Bash command-line (see Section “Command Line Editing” in Bash Features), with TAB-completion of commands, devices, partitions, and files in a directory depending on context.

Support multiple filesystem types Support multiple filesystem types transparently, plus a useful explicit blocklist notation. The currently supported filesystem types are Amiga Fast FileSystem (AFFS), AtheOS fs, BeFS, BtrFS (including raid0, raid1, raid10, gzip and lzo), cpio (little- and big-endian bin, odc and newc variants), Linux ext2/ext3/ext4, DOS FAT12/FAT16/FAT32, exFAT, F2FS, HFS, HFS+, ISO9660 (including Joliet, Rock-ridge and multi-chunk files), JFS, Minix fs (versions 1, 2 and 3), nilfs2, NTFS (including compression), ReiserFS, ROMFS, Amiga Smart FileSystem (SFS), Squash4, tar, UDF, BSD UFS/UFS2, XFS, and ZFS (including lzjb, gzip, zle, mirror, stripe, raidz1/2/3 and encryption in AES-CCM and AES-GCM). See Chapter 13 [Filesystem], page 57, for more information.

Support automatic decompression Can decompress files which were compressed by gzip or xz^{3}. This function is both automatic and transparent to the user (i.e. all functions operate upon the uncompressed contents of the specified files). This greatly reduces a file size and loading time, a particularly great benefit for floppies.^{4} It is conceivable that some kernel modules should be loaded in a compressed state, so a different module-loading command can be specified to avoid uncompressing the modules.

Access data on any installed device Support reading data from any or all floppies or hard disk(s) recognized by the BIOS, independent of the setting of the root device.

Be independent of drive geometry translations Unlike many other boot loaders, GRUB makes the particular drive translation irrelevant. A drive installed and running with one translation may be converted to another translation without any adverse effects or changes in GRUB’s configuration.

Detect all installed ram GRUB can generally find all the installed ram on a PC-compatible machine. It uses an advanced BIOS query technique for finding all memory regions. As described on the Multiboot Specification (see Section “Motivation” in The Multiboot Specification), not all kernels make use of this information, but GRUB provides it for those who do.

3 Only CRC32 data integrity check is supported (xz default is CRC64 so one should use –check=crc32

option). LZMA BCJ filters are supported.

4 There are a few pathological cases where loading a very badly organized ELF kernel might take

longer, but in practice this never happen.

Support Logical Block Address mode In traditional disk calls (called CHS mode), there is a geometry translation problem, that is, the BIOS cannot access over 1024 cylinders, so the accessible space is limited to at least 508 MB and to at most 8GB. GRUB can’t universally solve this problem, as there is no standard interface used in all machines. However, several newer machines have the new interface, Logical Block Address (LBA) mode. GRUB automatically detects if LBA mode is available and uses it if available. In LBA mode, GRUB can access the entire disk.

Support network booting GRUB is basically a disk-based boot loader but also has network support. You can load OS images from a network by using the TFTP protocol.

Support remote terminals To support computers with no console, GRUB provides remote terminal support, so that you can control GRUB from a remote host. Only serial terminal support is implemented at the moment.

1.5 The role of a boot loader

The following is a quotation from Gordon Matzigkeit, a GRUB fanatic:

Some people like to acknowledge both the operating system and kernel when they talk about their computers, so they might say they use “GNU/Linux” or “GNU/Hurd”. Other people seem to think that the kernel is the most important part of the system, so they like to call their GNU operating systems “Linux systems.” I, personally, believe that this is a grave injustice, because the boot loader is the most important software of all. I used to refer to the above systems as either “LILO”^{5} or “GRUB” systems. Unfortunately, nobody ever understood what I was talking about; now I just use the word “GNU” as a pseudonym for GRUB. So, if you ever hear people talking about their alleged “GNU” systems, remember that they are actually paying homage to the best boot loader around… GRUB! We, the GRUB maintainers, do not (usually) encourage Gordon’s level of fanaticism,

but it helps to remember that boot loaders deserve recognition. We hope that you enjoy using GNU GRUB as much as we did writing it.

5 The LInux LOader, a boot loader that everybody uses, but nobody likes.

2 Naming convention

The device syntax used in GRUB is a wee bit different from what you may have seen before in your operating system(s), and you need to know it so that you can specify a drive/partition.

Look at the following examples and explanations:

     (fd0)

First of all, GRUB requires that the device name be enclosed with ‘(’ and ‘)’. The ‘fd’ part means that it is a floppy disk. The number ‘0’ is the drive number, which is counted from zero. This expression means that GRUB will use the whole floppy disk.

     (hd0,msdos2)

Here, ‘hd’ means it is a hard disk drive. The first integer ‘0’ indicates the drive number, that is, the first hard disk, the string ‘msdos’ indicates the partition scheme, while the second integer, ‘2’, indicates the partition number (or the pc slice number in the BSD terminology). The partition numbers are counted from one, not from zero (as was the case in previous versions of GRUB). This expression means the second partition of the first hard disk drive. In this case, GRUB uses one partition of the disk, instead of the whole disk.

     (hd0,msdos5)

This specifies the first extended partition of the first hard disk drive. Note that the partition numbers for extended partitions are counted from ‘5’, regardless of the actual number of primary partitions on your hard disk.

     (hd1,msdos1,bsd1)

This means the BSD ‘a’ partition on first pc slice number of the second hard disk. Of course, to actually access the disks or partitions with GRUB, you need to use

the device specification in a command, like ‘set root=(fd0)’ or ‘parttool (hd0,msdos3) hidden-’. To help you find out which number specifies a partition you want, the GRUB command-line (see Section 14.1 [Command-line interface], page 61) options have argument completion. This means that, for example, you only need to type

     set root=(

followed by a TAB, and GRUB will display the list of drives, partitions, or file names. So it should be quite easy to determine the name of your target partition, even with minimal knowledge of the syntax.

Note that GRUB does not distinguish IDE from SCSI - it simply counts the drive numbers from zero, regardless of their type. Normally, any IDE drive number is less than any SCSI drive number, although that is not true if you change the boot sequence by swapping IDE and SCSI drives in your BIOS.

Now the question is, how to specify a file? Again, consider an example:

     (hd0,msdos1)/vmlinuz

This specifies the file named ‘vmlinuz’, found on the first partition of the first hard disk drive. Note that the argument completion works with file names, too.

That was easy, admit it. Now read the next chapter, to find out how to actually install GRUB on your drive.

3 OS-specific notes about grub tools

On OS which have device nodes similar to Unix-like OS GRUB tools use the OS name. E.g. for GNU/Linux:

# grub-install /dev/sda

On AROS we use another syntax. For volumes:

//:<volume name>

E.g.

//:DH0

For disks we use syntax:

//:<driver name>/unit/flags

E.g.

# grub-install //:ata.device/0/0

On Windows we use UNC path. For volumes it’s typically

\\?\Volume{<GUID>}
\\?\<drive letter>:

E.g.

\\?\Volume{17f34d50-cf64-4b02-800e-51d79c3aa2ff}
\\?\C:

For disks it’s

\\?\PhysicalDrive<number>

E.g.

# grub-install \\?\PhysicalDrive0

Beware that you may need to further escape the backslashes depending on your shell. When compiled with cygwin support then cygwin drive names are automatically

when needed. E.g.

# grub-install /dev/sda

4 Installation

In order to install GRUB as your boot loader, you need to first install the GRUB system and utilities under your UNIX-like operating system (see Appendix A [Obtaining and Building GRUB], page 135). You can do this either from the source tarball, or as a package for your OS.

After you have done that, you need to install the boot loader on a drive (floppy or hard disk) by using the utility grub-install (see Section 24.1 [Invoking grub-install], page 129) on a UNIX-like OS.

GRUB comes with boot images, which are normally put in the directory ‘/usr/lib/grub/<cpu>-<platform>’ (for BIOS-based machines ‘/usr/lib/grub/i386-pc’). Hereafter, the directory where GRUB images are initially placed (normally ‘/usr/lib/grub/<cpu>-<platform>’) will be called the image directory, and the directory where the boot loader needs to find them (usually ‘/boot’) will be called the boot directory.

4.1 Installing GRUB using grub-install

For information on where GRUB should be installed on PC BIOS platforms, see Section 4.4 [BIOS installation], page 13.

In order to install GRUB under a UNIX-like OS (such as gnu), invoke the program grub-install (see Section 24.1 [Invoking grub-install], page 129) as the superuser (root).

The usage is basically very simple. You only need to specify one argument to the program, namely, where to install the boot loader. The argument has to be either a device file (like ‘/dev/hda’). For example, under Linux the following will install GRUB into the MBR of the first IDE disk:

     # grub-install /dev/sda

Likewise, under GNU/Hurd, this has the same effect:

     # grub-install /dev/hd0

But all the above examples assume that GRUB should put images under the ‘/boot’ directory. If you want GRUB to put images under a directory other than ‘/boot’, you need to specify the option ‘--boot-directory’. The typical usage is that you create a GRUB boot floppy with a filesystem. Here is an example:

     # mke2fs /dev/fd0
     # mount -t ext2 /dev/fd0 /mnt
     # mkdir /mnt/boot
     # grub-install --boot-directory=/mnt/boot /dev/fd0
     # umount /mnt

Some BIOSes have a bug of exposing the first partition of a USB drive as a floppy instead of exposing the USB drive as a hard disk (they call it “USB-FDD” boot). In such cases, you need to install like this:

     # losetup /dev/loop0 /dev/sdb1
     # mount /dev/loop0 /mnt/usb
     # grub-install --boot-directory=/mnt/usb/bugbios --force --allow-floppy /dev/loop0

This install doesn’t conflict with standard install as long as they are in separate directories.

On EFI systems for fixed disk install you have to mount EFI System Partition. If you mount it at ‘/boot/efi’ then you don’t need any special arguments:

     # grub-install

Otherwise you need to specify where your EFI System partition is mounted:

     # grub-install --efi-directory=/mnt/efi

For removable installs you have to use ‘--removable’ and specify both

‘--boot-directory’ and ‘--efi-directory’:
     # grub-install --efi-directory=/mnt/usb --boot-directory=/mnt/usb/boot -
     -removable

4.2 Making a GRUB bootable CD-ROM

GRUB supports the no emulation mode in the El Torito specification^{1}. This means that you can use the whole CD-ROM from GRUB and you don’t have to make a floppy or hard disk image file, which can cause compatibility problems.

For booting from a CD-ROM, GRUB uses a special image called ‘cdboot.img’, which is concatenated with ‘core.img’. The ‘core.img’ used for this should be built with at least the ‘iso9660’ and ‘biosdisk’ modules. Your bootable CD-ROM will usually also need to include a configuration file ‘grub.cfg’ and some other GRUB modules.

To make a simple generic GRUB rescue CD, you can use the grub-mkrescue program (see Section 24.5 [Invoking grub-mkrescue], page 131):

     $ grub-mkrescue -o grub.iso

You will often need to include other files in your image. To do this, first make a top directory for the bootable image, say, ‘iso’:

     $ mkdir iso

Make a directory for GRUB:

     $ mkdir -p iso/boot/grub

If desired, make the config file ‘grub.cfg’ under ‘iso/boot/grub’ (see Chapter 6 [Configuration], page 21), and copy any files and directories for the disc to the directory

‘iso/’.

Finally, make the image:

     $ grub-mkrescue -o grub.iso iso

This produces a file named ‘grub.iso’, which then can be burned into a CD (or a DVD), or written to a USB mass storage device.

The root device will be set up appropriately on entering your ‘grub.cfg’ configuration file, so you can refer to file names on the CD without needing to use an explicit device name. This makes it easier to produce rescue images that will work on both optical drives and USB mass storage devices.

1 El Torito is a specification for bootable CD using BIOS functions.

4.3 The map between BIOS drives and OS devices

If the device map file exists, the GRUB utilities (grub-probe, etc.) read it to map BIOS drives to OS devices. This file consists of lines like this:

     (device) file

device is a drive specified in the GRUB syntax (see Section 13.1 [Device syntax], page 57), and file is an OS file, which is normally a device file.

Historically, the device map file was used because GRUB device names had to be used in the configuration file, and they were derived from BIOS drive numbers. The map between BIOS drives and OS devices cannot always be guessed correctly: for example, GRUB will get the order wrong if you exchange the boot sequence between IDE and SCSI in your BIOS.

Unfortunately, even OS device names are not always stable. Modern versions of the Linux kernel may probe drives in a different order from boot to boot, and the prefix (‘/dev/hd*’ versus ‘/dev/sd*’) may change depending on the driver subsystem in use. As a result, the device map file required frequent editing on some systems.

GRUB avoids this problem nowadays by using UUIDs or file system labels when generating ‘grub.cfg’, and we advise that you do the same for any custom menu entries you write. If the device map file does not exist, then the GRUB utilities will assume a temporary device map on the fly. This is often good enough, particularly in the common case of single-disk systems.

However, the device map file is not entirely obsolete yet, and it is used for overriding when current environment is different from the one on boot. Most common case is if you use a partition or logical volume as a disk for virtual machine. You can put any comments in the file if needed, as the GRUB utilities assume that a line is just a comment if the first character is ‘#’.

4.4 BIOS installation

MBR

The partition table format traditionally used on PC BIOS platforms is called the Master Boot Record (MBR) format; this is the format that allows up to four primary partitions and additional logical partitions. With this partition table format, there are two ways to install GRUB: it can be embedded in the area between the MBR and the first partition (called by various names, such as the "boot track", "MBR gap", or "embedding area", and which is usually at least 1000 KiB), or the core image can be installed in a file system and a list of the blocks that make it up can be stored in the first sector of that partition.

Modern tools usually leave MBR gap of at least 1023 KiB. This amount is sufficient to cover most configurations. Hence this value is recommended by the GRUB team.

Historically many tools left only 31 KiB of space. This is not enough to parse reliably difficult structures like Btrfs, ZFS, RAID or LVM, or to use difficult disk access methods like ahci. Hence GRUB will warn if attempted to install into small MBR gap except in a small number of configurations that were grandfathered. The grandfathered config must:

• use biosdisk as disk access module for ‘/boot’ • not use any additional partition maps to access ‘/boot’

• ‘/boot’ must be on one of following filesystems: AFFS, AFS, BFS, cpio, newc, odc, ext2/3/4, FAT, exFAT, F2FS, HFS, uncompressed HFS+, ISO9660, JFS, Minix, Minix2, Minix3, NILFS2, NTFS, ReiserFS, ROMFS, SFS, tar, UDF, UFS1, UFS2, XFS

MBR gap has few technical problems. There is no way to reserve space in the embedding area with complete safety, and some proprietary software is known to use it to make it difficult for users to work around licensing restrictions. GRUB works around it by detecting sectors by other software and avoiding them and protecting its own sectors using Reed-Solomon encoding.

GRUB team recommends having MBR gap of at least 1000 KiB.

Should it not be possible, GRUB has support for a fallback solution which is heavily recommended against. Installing to a filesystem means that GRUB is vulnerable to its blocks being moved around by filesystem features such as tail packing, or even by aggressive fsck implementations, so this approach is quite fragile; and this approach can only be used if the ‘/boot’ filesystem is on the same disk that the BIOS boots from, so that GRUB does not have to rely on guessing BIOS drive numbers.

The GRUB development team generally recommends embedding GRUB before the first partition, unless you have special requirements. You must ensure that the first partition starts at least 1000 KiB (2000 sectors) from the start of the disk; on modern disks, it is often a performance advantage to align partitions on larger boundaries anyway, so the first partition might start 1 MiB from the start of the disk.

GPT

Some newer systems use the GUID Partition Table (GPT) format. This was specified as part of the Extensible Firmware Interface (EFI), but it can also be used on BIOS platforms if system software supports it; for example, GRUB and GNU/Linux can be used in this configuration. With this format, it is possible to reserve a whole partition for GRUB, called the BIOS Boot Partition. GRUB can then be embedded into that partition without the risk of being overwritten by other software and without being contained in a filesystem which might move its blocks around.

When creating a BIOS Boot Partition on a GPT system, you should make sure that it is at least 31 KiB in size. (GPT-formatted disks are not usually particularly small, so we recommend that you make it larger than the bare minimum, such as 1 MiB, to allow plenty of room for growth.) You must also make sure that it has the proper partition type. Using GNU Parted, you can set this using a command such as the following:

     # parted /dev/disk set partition-number bios_grub on

If you are using gdisk, set the partition type to ‘0xEF02’. With partitioning programs that require setting the GUID directly, it should be

‘21686148-6449-6e6f-744e656564454649’.

Caution: Be very careful which partition you select! When GRUB finds a BIOS Boot Partition during installation, it will automatically overwrite part of it. Make sure that the partition does not contain any other data.

5 Booting

GRUB can load Multiboot-compliant kernels in a consistent way, but for some free operating systems you need to use some OS-specific magic.

5.1 How to boot operating systems

GRUB has three distinct boot methods: loading an operating system directly, using kexec from userspace, and chainloading another bootloader. Generally speaking, the first two are more desirable because you don’t need to install or maintain other boot loaders and GRUB is flexible enough to load an operating system from an arbitrary disk/partition. However, chainloading is sometimes required, as GRUB doesn’t support all existing operating systems natively.

5.1.1 How to boot an OS directly with GRUB

Multiboot (see Section “Motivation” in The Multiboot Specification) is the native format supported by GRUB. For the sake of convenience, there is also support for Linux, FreeBSD, NetBSD and OpenBSD. If you want to boot other operating systems, you will have to chain-load them (see Section 5.1.3 [Chain-loading], page 15).

FIXME: this section is incomplete.

1. Run the command boot (see Section 17.4.8 [boot], page 80).

However, DOS and Windows have some deficiencies, so you might have to use more complicated instructions. See Section 5.4.4 [DOS/Windows], page 18, for more information.

5.1.2 Kexec with grub2-emu

GRUB can be run in userspace by invoking the grub2-emu tool. It will read all configuration scripts as if booting directly (see Section 5.1.1 [Loading an operating system directly], page 15). With the --kexec flag, and kexec(8) support from the operating system, the linux command will directly boot the target image. For systems that lack working systemctl(1) support for kexec, passing the --kexec flag twice will fallback to invoking kexec(8) directly; note however that this fallback may be unsafe outside read-only environments, as it does not invoke shutdown machinery.

5.1.3 Chain-loading an OS

Operating systems that do not support Multiboot and do not have specific support in GRUB (specific support is available for Linux, FreeBSD, NetBSD and OpenBSD) must be chain-loaded, which involves loading another boot loader and jumping to it in real mode or via the firmware.

The chainloader command (see Section 17.2.1 [chainloader], page 75) is used to set this up. It is normally also necessary to load some GRUB modules and set the appropriate root device. Putting this together, we get something like this, for a Windows system on the first partition of the first hard disk:

menuentry "Windows" {
        insmod chain
        insmod ntfs

        set root=(hd0,1)
        chainloader +1
}

On systems with multiple hard disks, an additional workaround may be required. See Section 5.4.4 [DOS/Windows], page 18.

Chain-loading is only supported on PC BIOS and EFI platforms.

5.2 Loopback booting

GRUB is able to read from an image (be it one of CD or HDD) stored on any of its accessible storages (refer to see Section 17.4.43 [loopback], page 88 command). However the OS itself should be able to find its root. This usually involves running a userspace program running before the real root is discovered. This is achieved by GRUB loading a specially made small image and passing it as ramdisk to the kernel. This is achieved by commands

kfreebsd_module, knetbsd_module_elf, kopenbsd_ramdisk, initrd (see Section 17.2.2

[initrd], page 75), initrd16 (see Section 17.2.3 [initrd16], page 76), multiboot_module, multiboot2_module or xnu_ramdisk depending on the loader. Note that for knetbsd the image must be put inside miniroot.kmod and the whole miniroot.kmod has to be loaded. In kopenbsd payload this is disabled by default. Additionally, behaviour of initial ramdisk depends on command line options. Several distributors provide the image for this purpose or it’s integrated in their standard ramdisk and activated by special option. Consult your kernel and distribution manual for more details. Other loaders like appleloader,

chainloader (BIOS, EFI, coreboot), freedos, ntldr, plan9 and truecrypt provide no

possibility of loading initial ramdisk and as far as author is aware the payloads in question don’t support either initial ramdisk or discovering loopback boot in other way and as such not bootable this way. Please consider alternative boot methods like copying all files from the image to actual partition. Consult your OS documentation for more details.

5.3 Booting from LVM cache logical volume

The LVM cache logical volume is the logical volume consisting of the original and the cache pool logical volume. The original is usually on a larger and slower storage device while the cache pool is on a smaller and faster one. The performance of the original volume can be improved by storing the frequently used data on the cache pool to utilize the greater performance of faster device.

GRUB boots from LVM cache logical volume merely by reading it’s original logical volume so that dirty data in cache pool volume is disregarded. This is not a problem for "writethrough" cache mode as it ensures that any data written will be stored both on the cache and the origin LV. For the other cache mode "writeback", which delays writing from the cache pool back to the origin LV to boost performance, GRUB may fail to boot in the wake of accidental power outage due to it’s inability to assemble the cache device for reading the required dirty data left behind. The situation will be improved after adding full support to the LVM cache logical volume in the future.

5.4 Some caveats on OS-specific issues

Here, we describe some caveats on several operating systems.

5.4.1 GNU/Hurd

Since GNU/Hurd is Multiboot-compliant, it is easy to boot it; there is nothing special about it. But do not forget that you have to specify a root partition to the kernel.

1. Set GRUB’s root device to the same drive as GNU/Hurd’s. The command search --set=root --file /boot/gnumach.gz or similar may help you (see Section 17.4.65 [search], page 93).

2. Load the kernel and the modules, like this:

     grub> multiboot /boot/gnumach.gz root=device:hd0s1
     grub> module  /hurd/ext2fs.static ext2fs --readonly \
                        --multiboot-command-line='${kernel-command-line}' \
                        --host-priv-port='${host-port}' \
                        --device-master-port='${device-port}' \
                        --exec-server-task='${exec-task}' -T typed '${root}' \
                        '$(task-create)' '$(task-resume)'
     grub> module /lib/ld.so.1 exec /hurd/exec '$(exec-task=task-create)'

3. Finally, run the command boot (see Section 17.4.8 [boot], page 80).

5.4.2 GNU/Linux

It is relatively easy to boot GNU/Linux from GRUB, because it somewhat resembles to boot a Multiboot-compliant OS.

1. Set GRUB’s root device to the same drive as GNU/Linux’s. The command search --set=root --file /vmlinuz or similar may help you (see Section 17.4.65 [search], page 93).

2. Load the kernel using the command linux (see Section 17.2.4 [linux], page 76):

     grub> linux /vmlinuz root=/dev/sda1

If you need to specify some kernel parameters, just append them to the command. For example, to set ‘acpi’ to ‘off’, do this:

     grub> linux /vmlinuz root=/dev/sda1 acpi=off

See the documentation in the Linux source tree for complete information on the available options. With linux GRUB uses 32-bit protocol. Some BIOS services like APM or EDD aren’t available with this protocol. In this case you need to use linux16

     grub> linux16 /vmlinuz root=/dev/sda1 acpi=off

3. If you use an initrd, execute the command initrd (see Section 17.2.2 [initrd], page 75)

after linux:

     grub> initrd /initrd

If you used linux16 you need to use initrd16:

     grub> initrd16 /initrd

4. Finally, run the command boot (see Section 17.4.8 [boot], page 80).

5.4.3 NetBSD

Booting a NetBSD kernel from GRUB is also relatively easy: first set GRUB’s root device, then load the kernel and the modules, and finally run boot.

1. Set GRUB’s root device to the partition holding the NetBSD root file system. For a disk with a NetBSD disk label, this is usually the first partition (a:). In that case, and assuming that the partition is on the first hard disk, set GRUB’s root device as follows:

         grub> insmod part_bsd
         grub> set root=(hd0,netbsd1)

For a disk with a GUID Partition Table (GPT), and assuming that the NetBSD root partition is the third GPT partition, do this:

         grub> insmod part_gpt
         grub> set root=(hd0,gpt3)

2. Load the kernel using the command knetbsd:

         grub> knetbsd /netbsd

Various options may be given to knetbsd. These options are, for the most part, the same as in the NetBSD boot loader. For instance, to boot the system in single-user mode and with verbose messages, do this:

         grub> knetbsd /netbsd -s -v

3. If needed, load kernel modules with the command knetbsd_module_elf. A typical

example is the module for the root file system:

         grub> knetbsd_module_elf /stand/amd64/6.0/modules/ffs/ffs.kmod

4. Finally, run the command boot (see Section 17.4.8 [boot], page 80).

5.4.4 DOS/Windows

GRUB cannot boot DOS or Windows directly, so you must chain-load them (see Section 5.1.3 [Chain-loading], page 15). However, their boot loaders have some critical deficiencies, so it may not work to just chain-load them. To overcome the problems, GRUB provides you with two helper functions.

If you have installed DOS (or Windows) on a non-first hard disk, you have to use the disk swapping technique, because that OS cannot boot from any disks but the first one. The workaround used in GRUB is the command drivemap (see Section 17.4.23 [drivemap], page 83), like this:

     drivemap -s (hd0) (hd1)

This performs a virtual swap between your first and second hard drive.

Caution: This is effective only if DOS (or Windows) uses BIOS to access the swapped disks. If that OS uses a special driver for the disks, this probably won’t work.

Another problem arises if you installed more than one set of DOS/Windows onto one disk, because they could be confused if there are more than one primary partitions for DOS/Windows. Certainly you should avoid doing this, but there is a solution if you do want to do so. Use the partition hiding/unhiding technique.

If GRUB hides a DOS (or Windows) partition (see Section 17.4.53 [parttool], page 90), DOS (or Windows) will ignore the partition. If GRUB unhides a DOS (or Windows) partition, DOS (or Windows) will detect the partition. Thus, if you have installed DOS (or Windows) on the first and the second partition of the first hard disk, and you want to boot the copy on the first partition, do the following:

parttool (hd0,1) hidden-
parttool (hd0,2) hidden+
set root=(hd0,1)
chainloader +1
parttool ${root} boot+
boot

6 Writing your own configuration file

GRUB is configured using ‘grub.cfg’, usually located under ‘/boot/grub’. This file is quite flexible, but most users will not need to write the whole thing by hand.

6.1 Simple configuration handling

The program grub-mkconfig (see Section 24.2 [Invoking grub-mkconfig], page 130) generates ‘grub.cfg’ files suitable for most cases. It is suitable for use when upgrading a distribution, and will discover available kernels and attempt to generate menu entries for them.

grub-mkconfig does have some limitations. While adding extra custom menu entries to the end of the list can be done by editing ‘/etc/grub.d/40_custom’ or creating ‘/boot/grub/custom.cfg’, changing the order of menu entries or changing their titles may require making complex changes to shell scripts stored in ‘/etc/grub.d/’. This may be improved in the future. In the meantime, those who feel that it would be easier to write ‘grub.cfg’ directly are encouraged to do so (see Chapter 5 [Booting], page 15, and Section 6.3 [Shell-like scripting], page 28), and to disable any system provided by their distribution to automatically run grub-mkconfig.

The file ‘/etc/default/grub’ controls the operation of grub-mkconfig. It is sourced by a shell script, and so must be valid POSIX shell input; normally, it will just be a sequence of ‘KEY=value’ lines, but if the value contains spaces or other special characters then it must be quoted. For example:

     GRUB_TERMINAL_INPUT="console serial"

Valid keys in ‘/etc/default/grub’ are as follows:

‘GRUB_DEFAULT’

The default menu entry. This may be a number, in which case it identifies the Nth entry in the generated menu counted from zero, or the title of a menu entry, or the special string ‘saved’. Using the id may be useful if you want to set a menu entry as the default even though there may be a variable number of entries before it. For example, if you have:

          menuentry 'Example GNU/Linux distribution' --class gnu-linux --id example-gnu-linux {
                  ...
          }

then you can make this the default using:

               GRUB_DEFAULT=example-gnu-linux

Previously it was documented the way to use entry title. While this still works it’s not recommended since titles often contain unstable device names and may be translated If you set this to ‘saved’, then the default menu entry will be that saved by

          ‘GRUB_SAVEDEFAULT’ or grub-set-default. This relies on the environment

block, which may not be available in all situations (see Section 15.2 [Environ-

ment block], page 70). The default is ‘0’.

‘GRUB_SAVEDEFAULT’

If this option is set to ‘true’, then, when an entry is selected, save it as a new default entry for use by future runs of GRUB. This is only useful if

‘GRUB_DEFAULT=saved’; it is a separate option because ‘GRUB_DEFAULT=saved’

is useful without this option, in conjunction with grub-set-default. Unset by default. This option relies on the environment block, which may not be available in all situations (see Section 15.2 [Environment block], page 70).

‘GRUB_TIMEOUT’

Boot the default entry this many seconds after the menu is displayed, unless a key is pressed. The default is ‘5’. Set to ‘0’ to boot immediately without displaying the menu, or to ‘-1’ to wait indefinitely.

If ‘GRUB_TIMEOUT_STYLE’ is set to ‘countdown’ or ‘hidden’, the timeout is in-

stead counted before the menu is displayed.

‘GRUB_TIMEOUT_STYLE’

If this option is unset or set to ‘menu’, then GRUB will display the menu and then wait for the timeout set by ‘GRUB_TIMEOUT’ to expire before booting the default entry. Pressing a key interrupts the timeout. If this option is set to ‘countdown’ or ‘hidden’, then, before displaying the menu, GRUB will wait for the timeout set by ‘GRUB_TIMEOUT’ to expire. If ESC or F4 are pressed, or SHIFT is held down during that time, it will display the menu and wait for input. If a hotkey associated with a menu entry is pressed, it will boot the associated menu entry immediately. If the timeout expires before either of these happens, it will boot the default entry. In the ‘countdown’ case, it will show a one-line indication of the remaining time.

‘GRUB_DEFAULT_BUTTON’
‘GRUB_TIMEOUT_BUTTON’
‘GRUB_TIMEOUT_STYLE_BUTTON’
‘GRUB_BUTTON_CMOS_ADDRESS’

Variants of the corresponding variables without the ‘_BUTTON’ suffix, used to support vendor-specific power buttons. See Chapter 10 [Vendor power-on keys], page 51.

‘GRUB_DISTRIBUTOR’

Set by distributors of GRUB to their identifying name. This is used to generate more informative menu entry titles.

‘GRUB_TERMINAL_INPUT’

Select the terminal input device. You may select multiple devices here, separated by spaces. Valid terminal input names depend on the platform, but may include ‘console’ (native platform console), ‘serial’ (serial terminal), ‘serial_<port>’ (serial terminal with explicit port selection), ‘at_keyboard’ (PC AT keyboard), or ‘usb_keyboard’ (USB keyboard using the HID Boot Protocol, for cases where the firmware does not handle this). The default is to use the platform’s native terminal input.

‘GRUB_TERMINAL_OUTPUT’

Select the terminal output device. You may select multiple devices here, separated by spaces.

Valid terminal output names depend on the platform, but may include ‘console’ (native platform console), ‘serial’ (serial terminal), ‘serial_<port>’ (serial terminal with explicit port selection), ‘gfxterm’ (graphics-mode output), ‘vga_text’ (VGA text output), ‘mda_text’ (MDA text output), ‘morse’ (Morse-coding using system beeper) or ‘spkmodem’ (simple data protocol using system speaker).

‘spkmodem’ is useful when no serial port is available. Connect the output of sending system (where GRUB is running) to line-in of receiving system (usually developer machine). On receiving system compile ‘spkmodem-recv’ from

‘util/spkmodem-recv.c’ and run:
     parecord --channels=1 --rate=48000 --format=s16le | ./spkmodem-recv

The default is to use the platform’s native terminal output.

‘GRUB_TERMINAL’

If this option is set, it overrides both ‘GRUB_TERMINAL_INPUT’ and

‘GRUB_TERMINAL_OUTPUT’ to the same value.
‘GRUB_SERIAL_COMMAND’

A command to configure the serial port when using the serial console. See Section 17.3.1 [serial], page 77. Defaults to ‘serial’.

‘GRUB_CMDLINE_LINUX’

Command-line arguments to add to menu entries for the Linux kernel.

‘GRUB_CMDLINE_LINUX_DEFAULT’

Unless ‘GRUB_DISABLE_RECOVERY’ is set to ‘true’, two menu entries will be generated for each Linux kernel: one default entry and one entry for recovery mode. This option lists command-line arguments to add only to the default menu entry, after those listed in ‘GRUB_CMDLINE_LINUX’.

‘GRUB_CMDLINE_LINUX_RECOVERY’

Unless ‘GRUB_DISABLE_RECOVERY’ is set to ‘true’, two menu entries will be generated for each Linux kernel: one default entry and one entry for recovery mode. This option lists command-line arguments to add only to the recovery menu entry, before those listed in ‘GRUB_CMDLINE_LINUX’. The default is

‘single’.
‘GRUB_CMDLINE_NETBSD’
‘GRUB_CMDLINE_NETBSD_DEFAULT’
As ‘GRUB_CMDLINE_LINUX’ and ‘GRUB_CMDLINE_LINUX_DEFAULT’, but for

NetBSD.

‘GRUB_CMDLINE_GNUMACH’

As ‘GRUB_CMDLINE_LINUX’, but for GNU Mach.

‘GRUB_CMDLINE_XEN’
‘GRUB_CMDLINE_XEN_DEFAULT’

The values of these options are passed to Xen hypervisor Xen menu entries, for all respectively normal entries.

‘GRUB_CMDLINE_LINUX_XEN_REPLACE’
‘GRUB_CMDLINE_LINUX_XEN_REPLACE_DEFAULT’

The values of these options replace the values of ‘GRUB_CMDLINE_LINUX’ and ‘GRUB_CMDLINE_LINUX_DEFAULT’ for Linux and Xen menu entries.

‘GRUB_TOP_LEVEL’
‘GRUB_TOP_LEVEL_XEN’

This option should be an absolute path to a kernel image. If provided, the image specified will be made the top-level entry if it is found in the scan.

‘GRUB_TOP_LEVEL_OS_PROBER’

This option should be a line of output from os-prober. As ‘GRUB_TOP_LEVEL’, if provided, the image specified will be made the top-level entry if it is found in the scan.

‘GRUB_EARLY_INITRD_LINUX_CUSTOM’
‘GRUB_EARLY_INITRD_LINUX_STOCK’

List of space-separated early initrd images to be loaded from ‘/boot’. This is for loading things like CPU microcode, firmware, ACPI tables, crypto keys, and so on. These early images will be loaded in the order declared, and all will be loaded before the actual functional initrd image. ‘GRUB_EARLY_INITRD_LINUX_STOCK’ is for your distribution to declare images that are provided by the distribution. It should not be modified without understanding the consequences. They will be loaded first. ‘GRUB_EARLY_INITRD_LINUX_CUSTOM’ is for your custom created images. The default stock images are as follows, though they may be overridden by your distribution:

     intel-uc.img intel-ucode.img amd-uc.img amd-ucode.img early_ucode.cpio microcode.cpio
‘GRUB_DISABLE_LINUX_UUID’

Normally, grub-mkconfig will generate menu entries that use universallyunique identifiers (UUIDs) to identify the root filesystem to the Linux kernel, using a ‘root=UUID=...’ kernel parameter. This is usually more reliable, but in some cases it may not be appropriate. To disable the use of UUIDs, set this option to ‘true’.

‘GRUB_DISABLE_LINUX_PARTUUID’

If grub-mkconfig cannot identify the root filesystem via its universally-unique indentifier (UUID), grub-mkconfig can use the UUID of the partition containing the filesystem to identify the root filesystem to the Linux kernel via a ‘root=PARTUUID=...’ kernel parameter. This is not as reliable as using the filesystem UUID, but is more reliable than using the Linux device names. When ‘GRUB_DISABLE_LINUX_PARTUUID’ is set to ‘false’, the Linux kernel version must be 2.6.37 (3.10 for systems using the MSDOS partition scheme) or newer.

This option defaults to ‘true’. To enable the use of partition UUIDs, set this option to ‘false’.

‘GRUB_DISABLE_RECOVERY’

If this option is set to ‘true’, disable the generation of recovery mode menu entries.

‘GRUB_DISABLE_UUID’

Normally, grub-mkconfig will generate menu entries that use universallyunique identifiers (UUIDs) to identify various filesystems to search for files. This is usually more reliable, but in some cases it may not be appropriate. To disable this use of UUIDs, set this option to ‘true’. Setting this option to ‘true’, will also set the options ‘GRUB_DISABLE_LINUX_UUID’ and ‘GRUB_DISABLE_LINUX_PARTUUID’ to ‘true’, unless they have been explicilty set to ‘false’.

‘GRUB_VIDEO_BACKEND’

If graphical video support is required, either because the ‘gfxterm’ graphical terminal is in use or because ‘GRUB_GFXPAYLOAD_LINUX’ is set, then grubmkconfig will normally load all available GRUB video drivers and use the one most appropriate for your hardware. If you need to override this for some reason, then you can set this option. After grub-install has been run, the available video drivers are listed in

‘/boot/grub/video.lst’.
‘GRUB_GFXMODE’

Set the resolution used on the ‘gfxterm’ graphical terminal. Note that you can only use modes which your graphics card supports via VESA BIOS Extensions (VBE), so for example native LCD panel resolutions may not be available. The default is ‘auto’, which tries to select a preferred resolution. See Section 15.1.12 [gfxmode], page 66.

‘GRUB_BACKGROUND’

Set a background image for use with the ‘gfxterm’ graphical terminal. The value of this option must be a file readable by GRUB at boot time, and it must end with ‘.png’, ‘.tga’, ‘.jpg’, or ‘.jpeg’. The image will be scaled if necessary to fit the screen. Image height and width will be restricted by an artificial limit of 16384.

‘GRUB_THEME’

Set a theme for use with the ‘gfxterm’ graphical terminal.

‘GRUB_GFXPAYLOAD_LINUX’

Set to ‘text’ to force the Linux kernel to boot in normal text mode, ‘keep’ to preserve the graphics mode set using ‘GRUB_GFXMODE’, ‘widthxheight’[‘xdepth’] to set a particular graphics mode, or a sequence of these separated by commas or semicolons to try several modes in sequence. See Section 15.1.13 [gfxpayload], page 66. Depending on your kernel, your distribution, your graphics card, and the phase of the moon, note that using this option may cause GNU/Linux to suffer from various display problems, particularly during the early part of the boot sequence. If you have problems, set this option to ‘text’ and GRUB will tell Linux to boot in normal text mode.

‘GRUB_DISABLE_OS_PROBER’

The grub-mkconfig has a feature to use the external os-prober program to discover other operating systems installed on the same machine and generate appropriate menu entries for them. It is disabled by default since automatic and silent execution of os-prober, and creating boot entries based on that data, is a potential attack vector. Set this option to ‘false’ to enable this feature in

the grub-mkconfig command.
‘GRUB_OS_PROBER_SKIP_LIST’

List of space-separated FS UUIDs of filesystems to be ignored from os-prober output. For efi chainloaders it’s <UUID>@<EFI FILE>

‘GRUB_DISABLE_SUBMENU’

Normally, grub-mkconfig will generate top level menu entry for the kernel with highest version number and put all other found kernels or alternative menu entries for recovery mode in submenu. For entries returned by os-prober first entry will be put on top level and all others in submenu. If this option is set to ‘true’, flat menu with all entries on top level will be generated instead. Changing this option will require changing existing values of ‘GRUB_DEFAULT’, ‘fallback’ (see Section 15.1.11 [fallback], page 66) and ‘default’ (see Section 15.1.10 [default], page 65) environment variables as well as saved default

entry using grub-set-default and value used with grub-reboot.
‘GRUB_ENABLE_CRYPTODISK’

If set to ‘y’, grub-mkconfig and grub-install will check for encrypted disks and generate additional commands needed to access them during boot. Note that in this case unattended boot is not possible because GRUB will wait for passphrase to unlock encrypted container.

‘GRUB_INIT_TUNE’

Play a tune on the speaker when GRUB starts. This is particularly useful for users unable to see the screen. The value of this option is passed directly to Section 17.4.57 [play], page 92.

‘GRUB_BADRAM’

If this option is set, GRUB will issue a Section 17.4.6 [badram], page 80 command to filter out specified regions of RAM.

‘GRUB_PRELOAD_MODULES’

This option may be set to a list of GRUB module names separated by spaces. Each module will be loaded as early as possible, at the start of ‘grub.cfg’.

The following options are still accepted for compatibility with existing configurations, but have better replacements:

‘GRUB_HIDDEN_TIMEOUT’

Wait this many seconds before displaying the menu. If ESC or F4 are pressed, or SHIFT is held down during that time, display the menu and wait for input according to ‘GRUB_TIMEOUT’. If a hotkey associated with a menu entry is pressed, boot the associated menu entry immediately. If the timeout expires before either of these happens, display the menu for the number of seconds specified in ‘GRUB_TIMEOUT’ before booting the default entry.

          If you set ‘GRUB_HIDDEN_TIMEOUT’, you should also set ‘GRUB_TIMEOUT=0’ so

that the menu is not displayed at all unless ESC or F4 are pressed, or SHIFT is held down. This option is unset by default, and is deprecated in favour of the less confusing

          ‘GRUB_TIMEOUT_STYLE=countdown’ or ‘GRUB_TIMEOUT_STYLE=hidden’.
‘GRUB_HIDDEN_TIMEOUT_QUIET’

In conjunction with ‘GRUB_HIDDEN_TIMEOUT’, set this to ‘true’ to suppress the verbose countdown while waiting for a key to be pressed before displaying the menu. This option is unset by default, and is deprecated in favour of the less confusing

          ‘GRUB_TIMEOUT_STYLE=countdown’.
‘GRUB_HIDDEN_TIMEOUT_BUTTON’

Variant of ‘GRUB_HIDDEN_TIMEOUT’, used to support vendor-specific power buttons. See Chapter 10 [Vendor power-on keys], page 51. This option is unset by default, and is deprecated in favour of the less confusing

          ‘GRUB_TIMEOUT_STYLE=countdown’ or ‘GRUB_TIMEOUT_STYLE=hidden’.

For more detailed customisation of grub-mkconfig’s output, you may edit the scripts in ‘/etc/grub.d’ directly. ‘/etc/grub.d/40_custom’ is particularly useful for adding entire custom menu entries; simply type the menu entries you want to add at the end of that file, making sure to leave at least the first two lines intact.

6.2 Root Identifcation Heuristics

If the target operating system uses the Linux kernel, grub-mkconfig attempts to identify the root file system via a heuristic algoirthm. This algorithm selects the identification method of the root file system by considering three factors. The first is if an initrd for the target operating system is also present. The second is ‘GRUB_DISABLE_LINUX_UUID’ and if set to ‘true’, prevents grub-mkconfig from identifying the root file system by its UUID.

The third is ‘GRUB_DISABLE_LINUX_PARTUUID’ and if set to ‘true’, prevents grub-mkconfig

from identifying the root file system via the UUID of its enclosing partition. If the variables are assigned any other value, that value is considered equivalent to ‘false’. The variables are also considered to be set to ‘false’ if they are not set.

When booting, the Linux kernel will delegate the task of mounting the root filesystem to the initrd. Most initrd images determine the root file system by checking the Linux kernel’s command-line for the ‘root’ key and use its value as the identification method of the root file system. To improve the reliability of booting, most initrd images also allow the root file system to be identified by its UUID. Because of this behavior, the grub-mkconfig command will set ‘root’ to ‘root=UUID=...’ to provide the initrd with the filesystem UUID of the root file system.

If no initrd is detected or ‘GRUB_DISABLE_LINUX_UUID’ is set to ‘true’ then grubcommand will identify the root filesystem by setting the kernel command-line variable ‘root’

to ‘root=PARTUUID=...’ unless ‘GRUB_DISABLE_LINUX_PARTUUID’ is also set to ‘true’. If
‘GRUB_DISABLE_LINUX_PARTUUID’ is also set to ‘true’, grub-command will identify by its

Linux device name.

The following table summarizes the behavior of the grub-mkconfig command.

Initrd detected GRUB DISABLE LINUX PARTUUID Set To GRUB DISABLE LINUX UUID Set To

Linux Root ID Method

false  false  false  part UUID
false  false  true  part UUID
false  true  false  dev name
false  true  true  dev name
true  false  false  fs UUID
true  false  true  part UUID
true  true  false  fs UUID
true  true  true  dev name

      Remember, ‘GRUB_DISABLE_LINUX_PARTUUID’ and ‘GRUB_DISABLE_LINUX_UUID’ are

also considered to be set to ‘true’ and ‘false’, respectively, when they are unset.

6.3 Writing full configuration files directly

‘grub.cfg’ is written in GRUB’s built-in scripting language, which has a syntax quite similar to that of GNU Bash and other Bourne shell derivatives.

Words

A word is a sequence of characters considered as a single unit by GRUB. Words are separated by metacharacters, which are the following plus space, tab, and newline:

     { } | & $ ; < >

Quoting may be used to include metacharacters in words; see below.

Reserved words

Reserved words have a special meaning to GRUB. The following words are recognised as reserved when unquoted and either the first word of a simple command or the third word of a for command:

     ! [[ ]] { }
     case do done elif else esac fi for function
     if in menuentry select then time until while

Not all of these reserved words have a useful purpose yet; some are reserved for future expansion.

Quoting

Quoting is used to remove the special meaning of certain characters or words. It can be used to treat metacharacters as part of a word, to prevent reserved words from being recognised as such, and to prevent variable expansion.

There are three quoting mechanisms: the escape character, single quotes, and double quotes.

A non-quoted backslash (\) is the escape character. It preserves the literal value of the next character that follows, with the exception of newline.

Enclosing characters in single quotes preserves the literal value of each character within the quotes. A single quote may not occur between single quotes, even when preceded by a backslash.

Enclosing characters in double quotes preserves the literal value of all characters within the quotes, with the exception of ‘$’ and ‘\’. The ‘$’ character retains its special meaning within double quotes. The backslash retains its special meaning only when followed by one of the following characters: ‘$’, ‘"’, ‘\’, or newline. A backslash-newline pair is treated as a line continuation (that is, it is removed from the input stream and effectively ignored^{1}). A double quote may be quoted within double quotes by preceding it with a backslash.

Variable expansion

The ‘$’ character introduces variable expansion. The variable name to be expanded may be enclosed in braces, which are optional but serve to protect the variable to be expanded from characters immediately following it which could be interpreted as part of the name.

Normal variable names begin with an alphabetic character, followed by zero or more alphanumeric characters. These names refer to entries in the GRUB environment (see Chapter 15 [Environment], page 63).

Positional variable names consist of one or more digits. They represent parameters passed to function calls, with ‘$1’ representing the first parameter, and so on.

The special variable name ‘?’ expands to the exit status of the most recently executed command. When positional variable names are active, other special variable names ‘@’, ‘*’ and ‘#’ are defined and they expand to all positional parameters with necessary quoting, positional parameters without any quoting, and positional parameter count respectively.

Comments

A word beginning with ‘#’ causes that word and all remaining characters on that line to be ignored.

Simple commands

A simple command is a sequence of words separated by spaces or tabs and terminated by a semicolon or a newline. The first word specifies the command to be executed. The remaining words are passed as arguments to the invoked command.

The return value of a simple command is its exit status. If the reserved word ! precedes the command, then the return value is instead the logical negation of the command’s exit status.

1 Currently a backslash-newline pair within a variable name is not handled properly, so use this feature

with some care.

Compound commands

A compound command is one of the following:

for name in word …; do list; done The list of words following in is expanded, generating a list of items. The variable name is set to each element of this list in turn, and list is executed each time. The return value is the exit status of the last command that executes. If the expansion of the items following in results in an empty list, no commands are executed, and the return status is 0.

if list; then list; [elif list; then list;] … [else list;] fi The if list is executed, where list is a series of simple commands separated by a ";". If its exit status of the last command is zero, the then list is executed. Otherwise, each elif list is executed in turn, and if its last command’s exit status is zero, the corresponding then list is executed and the command completes. Otherwise, the else list is executed, if present. The exit status is the exit status of the last command executed, or zero if no condition tested true.

while cond; do list; done until cond; do list; done

The while command continuously executes the do list as long as the last command in cond returns an exit status of zero, where cond is a list of simple commands separated by a ";". The until command is identical to the while command, except that the test is negated; the do list is executed as long as the last command in cond returns a non-zero exit status. The exit status of the while and until commands is the exit status of the last do list command executed, or zero if none was executed.

function name { command; … } This defines a function named name. The body of the function is the list of commands within braces, each of which must be terminated with a semicolon or a newline. This list of commands will be executed whenever name is specified as the name of a simple command. Function definitions do not affect the exit status in $?. When executed, the exit status of a function is the exit status of the last command executed in the body.

menuentry title [‘--class=class’ . . .] [‘--users=users’] [‘--unrestricted’]
[‘--hotkey=key’] [‘--id=id’] { command; . . . }

See Section 17.1.1 [menuentry], page 73.

Built-in Commands

Some built-in commands are also provided by GRUB script to help script writers perform actions that are otherwise not possible. For example, these include commands to jump out of a loop without fully completing it, etc.

break [n] Exit from within a for, while, or until loop. If n is specified, break n levels. n must be greater than or equal to 1. If n is greater than the number of enclosing loops, all enclosing loops are exited. The return value is 0 unless n is not greater than or equal to 1. continue [n]

Resume the next iteration of the enclosing for, while or until loop. If n is specified, resume at the nth enclosing loop. n must be greater than or equal to 1. If n is greater than the number of enclosing loops, the last enclosing loop (the top-level loop) is resumed. The return value is 0 unless n is not greater than or equal to 1.

return [n] Causes a function to exit with the return value specified by n. If n is omitted, the return status is that of the last command executed in the function body. If used outside a function the return status is false.

setparams [arg] … Replace positional parameters starting with $1 with arguments to setparams.

shift [n] The positional parameters from n+1 … are renamed to $1. . .. Parameters represented by the numbers $# down to $#-n+1 are unset. n must be a nonnegative number less than or equal to $#. If n is 0, no parameters are changed. If n is not given, it is assumed to be 1. If n is greater than $#, the positional parameters are not changed. The return status is greater than zero if n is greater than $# or less than zero; otherwise 0.

6.4 Multi-boot manual config

Currently autogenerating config files for multi-boot environments depends on os-prober and has several shortcomings. Due to that it is disabled by default. It is advised to use the power of GRUB syntax and do it yourself. A possible configuration is detailed here, feel free to adjust to your needs.

First create a separate GRUB partition, big enough to hold GRUB. Some of the following entries show how to load OS installer images from this same partition, for that you obviously need to make the partition large enough to hold those images as well. Mount this partition on/mnt/boot and disable GRUB in all OSes and manually install self-compiled latest GRUB with:

      grub-install --boot-directory=/mnt/boot /dev/sda

In all the OSes install GRUB tools but disable installing GRUB in bootsector, so you’ll have menu.lst and grub.cfg available for use. Also disable os-prober use by setting:

      GRUB_DISABLE_OS_PROBER=true

in /etc/default/grub

Then write a grub.cfg (/mnt/boot/grub/grub.cfg):

     menuentry "OS using grub2" {
        insmod xfs
        search --set=root --label OS1 --hint hd0,msdos8
        configfile /boot/grub/grub.cfg
     }
     menuentry "OS using grub2-legacy" {
        insmod ext2

search --set=root --label OS2 --hint hd0,msdos6
legacy_configfile /boot/grub/menu.lst
}
menuentry "Windows XP" {
insmod ntfs
search --set=root --label WINDOWS_XP --hint hd0,msdos1
ntldr /ntldr
}
menuentry "Windows 7" {
insmod ntfs
search --set=root --label WINDOWS_7 --hint hd0,msdos2
ntldr /bootmgr
}
menuentry "FreeBSD" {
insmod zfs
search --set=root --label freepool --hint hd0,msdos7
kfreebsd /freebsd@/boot/kernel/kernel
kfreebsd_module_elf /freebsd@/boot/kernel/opensolaris.ko
kfreebsd_module_elf /freebsd@/boot/kernel/zfs.ko
kfreebsd_module /freebsd@/boot/zfs/zpool.cache type=/boot/zfs/zpool.cache
set kFreeBSD.vfs.root.mountfrom=zfs:freepool/freebsd
set kFreeBSD.hw.psm.synaptics_support=1
}
menuentry "experimental GRUB" {
search --set=root --label GRUB --hint hd0,msdos5
multiboot /experimental/grub/i386-pc/core.img
}
menuentry "Fedora 16 installer" {
search --set=root --label GRUB --hint hd0,msdos5
linux /fedora/vmlinuz lang=en_US keymap=sg resolution=1280x800
initrd /fedora/initrd.img
}
menuentry "Fedora rawhide installer" {
search --set=root --label GRUB --hint hd0,msdos5
linux /fedora/vmlinuz repo=ftp://mirror.switch.ch/mirror/fedora/linux/development/rawhide/x86_64 lang=en_US keymap=sg resolution=1280x800
initrd /fedora/initrd.img
}
menuentry "Debian sid installer" {
search --set=root --label GRUB --hint hd0,msdos5
linux /debian/dists/sid/main/installer-amd64/current/images/hd-media/vmlinuz

               initrd /debian/dists/sid/main/installer-amd64/current/images/hd-media/initrd.gz
     }

Notes:

• Argument to search after –label is FS LABEL. You can also use UUIDs with –fs-uuid UUID instead of –label LABEL. You could also use direct root=hd0,msdosX but this is not recommended due to device name instability.

6.5 Embedding a configuration file into GRUB

GRUB supports embedding a configuration file directly into the core image, so that it is loaded before entering normal mode. This is useful, for example, when it is not straightforward to find the real configuration file, or when you need to debug problems with loading that file. grub-install uses this feature when it is not using BIOS disk functions or when installing to a different disk from the one containing ‘/boot/grub’, in which case it needs to use the search command (see Section 17.4.65 [search], page 93) to find ‘/boot/grub’.

To embed a configuration file, use the ‘-c’ option to grub-mkimage. The file is copied into the core image, so it may reside anywhere on the file system, and may be removed after

running grub-mkimage.

After the embedded configuration file (if any) is executed, GRUB will load the ‘normal’ module (see Section 17.4.51 [normal], page 89), which will then read the real configuration file from ‘$prefix/grub.cfg’. By this point, the root variable will also have been set to the root device name. For example, prefix might be set to ‘(hd0,1)/boot/grub’, and root might be set to ‘hd0,1’. Thus, in most cases, the embedded configuration file only needs to set the prefix and root variables, and then drop through to GRUB’s normal processing. A typical example of this might look like this:

     search.fs_uuid 01234567-89ab-cdef-0123-456789abcdef root
     set prefix=($root)/boot/grub

(The ‘search_fs_uuid’ module must be included in the core image for this example to work.)

In more complex cases, it may be useful to read other configuration files directly from the embedded configuration file. This allows such things as reading files not called ‘grub.cfg’, or reading files from a directory other than that where GRUB’s loadable modules are installed. To do this, include the ‘configfile’ and ‘normal’ modules in the core image, and embed a configuration file that uses the configfile command to load another file. The following example of this also requires the echo, search_label, and test modules to be included in the core image:

search.fs_label grub root
if [ -e /boot/grub/example/test1.cfg ]; then
set prefix=($root)/boot/grub
configfile /boot/grub/example/test1.cfg
else
if [ -e /boot/grub/example/test2.cfg ]; then
    set prefix=($root)/boot/grub
    configfile /boot/grub/example/test2.cfg
else
    echo "Could not find an example configuration file!"
fi
fi

The embedded configuration file may not contain menu entries directly, but may only read them from elsewhere using configfile.

7 Theme file format

7.1 Introduction

The GRUB graphical menu supports themes that can customize the layout and appearance of the GRUB boot menu. The theme is configured through a plain text file that specifies the layout of the various GUI components (including the boot menu, timeout progress bar, and text messages) as well as the appearance using colors, fonts, and images. Example is available in docs/example theme.txt

7.2 Theme Elements

7.2.1 Colors

Colors can be specified in several ways:

• HTML-style “#RRGGBB” or “#RGB” format, where *R*, *G*, and *B* are hexadecimal digits (e.g., “#8899FF”)

• as comma-separated decimal RGB values (e.g., “128, 128, 255”)

• with “SVG 1.0 color names” (e.g., “cornflowerblue”) which must be specified in lowercase.

7.2.2 Fonts

The fonts GRUB uses “PFF2 font format” bitmap fonts. Fonts are specified with full font names. Currently there is no provision for a preference list of fonts, or deriving one font from another. Fonts are loaded with the “loadfont” command in GRUB (Section 17.4.42 [loadfont], page 88). To see the list of loaded fonts, execute the “lsfonts” command (Section 17.4.45 [lsfonts], page 88). If there are too many fonts to fit on screen, do “set pager=1” before executing “lsfonts”.

7.2.3 Progress Bar

Figure 7.1

Figure 7.2

Progress bars are used to display the remaining time before GRUB boots the default menu entry. To create a progress bar that will display the remaining time before automatic boot, simply create a “progress bar” component with the id “ timeout ”. This indicates to GRUB that the progress bar should be updated as time passes, and it should be made invisible if the countdown to automatic boot is interrupted by the user.

Progress bars may optionally have text displayed on them. This text is controlled by variable “text” which contains a printf template with the only argument %d is the number of seconds remaining. Additionally special values “@TIMEOUT NOTIFICATION SHORT@”, “@TIMEOUT NOTIFICATION MIDDLE@”, “@TIMEOUT NOTIFICATION LONG@” are replaced with standard and translated templates.

7.2.4 Circular Progress Indicator

The circular progress indicator functions similarly to the progress bar. When given an id of “ timeout ”, GRUB updates the circular progress indicator’s value to indicate the time remaining. For the circular progress indicator, there are two images used to render it: the *center* image, and the *tick* image. The center image is rendered in the center of the component, while the tick image is used to render each mark along the circumference of the indicator.

7.2.5 Labels

Text labels can be placed on the boot screen. The font, color, and horizontal alignment can be specified for labels. If a label is given the id “ timeout ”, then the “text” property for that label is also updated with a message informing the user of the number of seconds remaining until automatic boot. This is useful in case you want the text displayed somewhere else instead of directly on the progress bar.

7.2.6 Boot Menu

The boot menu where GRUB displays the menu entries from the “grub.cfg” file. It is a list of items, where each item has a title and an optional icon. The icon is selected based on the *classes* specified for the menu entry. If there is a PNG file named “myclass.png” in the “grub/themes/icons” directory, it will be displayed for items which have the class *myclass*. The boot menu can be customized in several ways, such as the font and color used for the menu entry title, and by specifying styled boxes for the menu itself and for the selected item highlight.

7.2.7 Styled Boxes

One of the most important features for customizing the layout is the use of *styled boxes*. A styled box is composed of 9 rectangular (and potentially empty) regions, which are used to seamlessly draw the styled box on screen:

Northwest (nw)  North (n)  Northeast (ne)
West (w)  Center (c)  East (e)
Southwest (sw)  South (s)  Southeast (se)

To support any size of box on screen, the center slice and the slices for the top, bottom, and sides are all scaled to the correct size for the component on screen, using the following rules:

1. The edge slices (north, south, east, and west) are scaled in the direction of the edge they are adjacent to. For instance, the west slice is scaled vertically.

2. The corner slices (northwest, northeast, southeast, and southwest) are not scaled.

3. The center slice is scaled to fill the remaining space in the middle.

As an example of how an image might be sliced up, consider the styled box used for a terminal view.

Figure 7.3

7.2.8 Creating Styled Box Images

The Inkscape scalable vector graphics editor is a very useful tool for creating styled box images. One process that works well for slicing a drawing into the necessary image slices is:

1. Create or open the drawing you’d like use.

2. Create a new layer on the top of the layer stack. Make it visible. Select this layer as the current layer.

3. Draw 9 rectangles on your drawing where you’d like the slices to be. Clear the fill option, and set the stroke to 1 pixel wide solid stroke. The corners of the slices must meet precisely; if it is off by a single pixel, it will probably be evident when the styled box is rendered in the GRUB menu. You should probably go to File | Document Properties | Grids and enable a grid or create a guide (click on one of the rulers next to the drawing and drag over the drawing; release the mouse button to place the guide) to help place the rectangles precisely.

4. Right click on the center slice rectangle and choose Object Properties. Change the "Id" to “slice c“ and click Set. Repeat this for the remaining 8 rectangles, giving them Id values of “slice n“, “slice ne“, “slice e“, and so on according to the location.

5. Save the drawing.

6. Select all the slice rectangles. With the slice layer selected, you can simply press Ctrl+A to select all rectangles. The status bar should indicate that 9 rectangles are selected.

7. Click the layer hide icon for the slice layer in the layer palette. The rectangles will remain selected, even though they are hidden.

8. Choose File | Export Bitmap and check the *Batch export 9 selected objects* box. Make sure that *Hide all except selected* is unchecked. click *Export*. This will create PNG files in the same directory as the drawing, named after the slices. These can now be used for a styled box in a GRUB theme.

7.3 Theme File Manual

The theme file is a plain text file. Lines that begin with “#“ are ignored and considered comments. (Note: This may not be the case if the previous line ended where a value was expected.)

The theme file contains two types of statements:

1. Global properties.

2. Component construction.

7.3.1 Global Properties

7.3.2 Format

Global properties are specified with the simple format:

• name1: value1

• name2: "value which may contain spaces"

• name3: #88F

In this example, name3 is assigned a color value.

7.3.3 Global Property List

title-text Specifies the text to display at the top center of the screen as a title.

title-font Defines the font used for the title message at the top of the screen.

title-color  Defines the color of the title message.
message-font  Currently unused. Left for backward compatibility.
message-color  Currently unused. Left for backward compatibility.
message-bg-color  Currently unused. Left for backward compatibility.

desktop-image Specifies the image to use as the background. It will be scaled to fit the screen size or proportionally scaled depending on the scale method.

desktop-image-scale-method Specifies the scaling method for the *desktopimage*. Options are “stretch“, “crop“, “padding“, “fitwidth“, “fitheight“. “stretch“ for fitting the screen size. Otherwise it is proportional scaling of a part of *desktop-image* to the part of the screen. “crop“ part of the *desktop-image* will be proportionally scaled to fit the screen sizes. “padding“ the entire *desktop-image* will be contained on the screen. “fitwidth“ for fitting the *desktop-image*’s width with screen width. “fitheight“ for fitting the *desktop-image*’s height with the screen height. Default is “stretch“.

desktop-image-h-align Specifies the horizontal alignment of the *desktopimage* if *desktop-image-scale-method* isn’t equeal to “stretch“. Options are “left“, “center“, “right“. Default is “center“.

desktop-image-v-align Specifies the vertical alignment of the *desktopimage* if *desktop-image-scale-method* isn’t equeal to “stretch“. Options are “top“, “center“, “bottom“. Default is “center“.

desktop-color Specifies the color for the background if *desktopimage* is not specified.

terminal-box Specifies the file name pattern for the styled box slices used for the command line terminal window. For example, “terminal-box: terminal *.png“ will use the images “terminal c.png“ as the center area, “terminal n.png“ as the north (top) edge, “terminal nw.png“ as the northwest (upper left) corner, and so on. If the image for any slice is not found, it will simply be left empty.

terminal-border  Specifies the border width of the terminal window.
terminal-left  Specifies the left coordinate of the terminal window.
terminal-top  Specifies the top coordinate of the terminal window.

terminal-width  Specifies the width of the terminal window.
terminal-height  Specifies the height of the terminal window.

7.3.4 Component Construction

Greater customizability comes is provided by components. A tree of components forms the user interface. *Containers* are components that can contain other components, and there is always a single root component which is an instance of a *canvas* container.

Components are created in the theme file by prefixing the type of component with a ’+’ sign:

  + label { text="GRUB" font="aqui 11" color="#8FF" }

properties of a component are specified as "name = value" (whitespace surrounding tokens is optional and is ignored) where *value* may be:

• a single word (e.g., “align = center“, “color = #FF8080“),

• a quoted string (e.g., “text = "Hello, World!"“), or

• a tuple (e.g., “preferred size = (120, 80)“).

7.3.5 Component List

The following is a list of the components and the properties they support.

• label A label displays a line of text.

Properties: id Set to “ timeout “ to display the time elapsed to an automatical boot of the default entry.

text The text to display. If “id“ is set to “ timeout “ and no “text“ property is set then the amount of seconds will be shown. If set to “@KEYMAP SHORT@“, “@KEYMAP MIDDLE@“ or “@KEYMAP LONG@“ then predefined hotkey information will be shown.

font  The font to use for text display.
color  The color of the text.

align The horizontal alignment of the text within the component. Options are “left“, “center“ and “right“.

visible  Set to “false“ to hide the label.

• image A component that displays an image. The image is scaled to fit the component. Properties:

file  The full path to the image file to load.

• progress bar Displays a horizontally oriented progress bar. It can be rendered using simple solid filled rectangles, or using a pair of pixmap styled boxes. Properties: id Set to “ timeout “ to display the time elapsed to an automatical boot of the default entry.

fg color  The foreground color for plain solid color rendering.
bg color  The background color for plain solid color rendering.
border color  The border color for plain solid color rendering.

text color  The text color.

bar style The styled box specification for the frame of the progress bar. Example: “progress frame *.png“ If the value is equal to “highlight style“ then no styled boxes will be shown.

highlight style The styled box specification for the highlighted region of the progress bar. This box will be used to paint just the highlighted region of the bar, and will be increased in size as the bar nears completion. Example: “progress hl *.png“. If the value is equal to “bar style“ then no styled boxes will be shown.

highlight overlay If this option is set to “true“ then the highlight box side slices (every slice except the center slice) will overlay the frame box side slices. And the center slice of the highlight box can move all the way (from top to bottom), being drawn on the center slice of the frame box. That way we can make a progress bar with round-shaped edges so there won’t be a free space from the highlight to the frame in top and bottom scrollbar positions. Default is “false“.

font  The font to use for progress bar.

text The text to display on the progress bar. If the progress bar’s ID is set to “ timeout “ and the value of this property is set to “@TIMEOUT NOTIFICATION SHORT@“, “@TIMEOUT NOTIFICATION MIDDLE@“ or “@TIME-OUT NOTIFICATION LONG@“, then GRUB will update this property with an informative message as the timeout approaches.

• circular progress Displays a circular progress indicator. The appearance of this component is determined by two images: the *center* image and the *tick* image. The center image is generally larger and will be drawn in the center of the component. Around the circumference of a circle within the component, the tick image will be drawn a certain number of times, depending on the properties of the component.

Properties:

id Set to “ timeout “ to display the time elapsed to an automatical boot of the default entry.

center bitmap The file name of the image to draw in the center of the component.

tick bitmap The file name of the image to draw for the tick marks.

num ticks  The number of ticks that make up a full circle.

ticks disappear Boolean value indicating whether tick marks should progressively appear, or progressively disappear as *value* approaches *end*. Specify “true“ or “false“. Default is “false“. start angle The position of the first tick mark to appear or disappear. Measured in "parrots", 1 "parrot" = 1 / 256 of the full circle. Use values “xxx deg“ or “xxx \xc2\xb0“ to set the angle in degrees.

• boot menu Displays the GRUB boot menu. It allows selecting items and executing them.

Properties:

item font  The font to use for the menu item titles.

selected item font The font to use for the selected menu item, or “inherit“ (the default) to use “item font“ for the selected menu item as well.

item color  The color to use for the menu item titles.

selected item color The color to use for the selected menu item, or “inherit“ (the default) to use “item color“ for the selected menu item as well.

icon width The width of menu item icons. Icons are scaled to the specified size.

icon height  The height of menu item icons.
item height  The height of each menu item in pixels.

item padding The amount of space in pixels to leave on each side of the menu item contents.

item icon space The space between an item’s icon and the title text, in pixels.

item spacing The amount of space to leave between menu items, in pixels.

menu pixmap style The image file pattern for the menu frame styled box. Example: “menu *.png“ (this will use images such as “menu c.png“, “menu w.png“, ‘menu nw.png“, etc.)

item pixmap style The image file pattern for the item styled box.

selected item pixmap style The image file pattern for the selected item highlight styled box.

scrollbar Boolean value indicating whether the scroll bar should be drawn if the frame and thumb styled boxes are configured.

scrollbar frame The image file pattern for the entire scroll bar. Example: “scrollbar *.png“

scrollbar thumb The image file pattern for the scroll bar thumb (the part of the scroll bar that moves as scrolling occurs). Example: “scrollbar thumb *.png“ scrollbar thumb overlay If this option is set to “true“ then the scrollbar thumb side slices (every slice except the center slice) will overlay the scrollbar frame side slices. And the center slice of the scrollbar thumb can move all the way (from top to bottom), being drawn on the center slice of the scrollbar frame. That way we can make a scrollbar with roundshaped edges so there won’t be a free space from the thumb to the frame in top and bottom scrollbar positions. Default is “false“.

scrollbar slice The menu frame styled box’s slice in which the scrollbar will be drawn. Possible values are “west“, “center“, “east“ (default). “west“ - the scrollbar will be drawn in the west slice (right-aligned). “east“ - the scrollbar will be drawn in the east slice (left-aligned). “center“ - the scrollbar will be drawn in the center slice. Note: in case of “center“ slice: a) If the scrollbar should be drawn then boot menu entry’s width is decreased by the scrollbar’s width and the scrollbar is drawn at the right side of the center slice. b) If the scrollbar won’t be drawn then the boot menu entry’s width is the width of the center slice. c) We don’t necessary need the menu pixmap box to display the scrollbar.

scrollbar left pad The left scrollbar padding in pixels. Unused if “scrollbar slice“ is “west“.

scrollbar right pad The right scrollbar padding in pixels. Unused if “scrollbar slice“ is “east“.

scrollbar top pad  The top scrollbar padding in pixels.
scrollbar bottom pad  The bottom scrollbar padding in pixels.
visible  Set to “false“ to hide the boot menu.

• canvas Canvas is a container that allows manual placement of components within it. It does not alter the positions of its child components. It assigns all child components their preferred sizes.

• hbox The *hbox* container lays out its children from left to right, giving each one its preferred width. The height of each child is set to the maximum of the preferred heights of all children.

• vbox The *vbox* container lays out its children from top to bottom, giving each one its preferred height. The width of each child is set to the maximum of the preferred widths of all children.

7.3.6 Common properties

The following properties are supported by all components:

‘left’ The distance from the left border of container to left border of the object in either of three formats:

x  Value in pixels
p%  Percentage
p%+x  mixture of both

‘top’ The distance from the left border of container to left border of the object in same format.

‘width’ The width of object in same format.

‘height’ The height of object in same format.

‘id’ The identifier for the component. This can be any arbitrary string. The ID can be used by scripts to refer to various components in the GUI component tree. Currently, there is one special ID value that GRUB recognizes: “ timeout “ Component with this ID will be updated by GRUB and will indicate time elapsed to an automatical boot of the default entry. Affected components: “label“, “circular progress“, “progress bar“.

8 Booting GRUB from the network

The following instructions don’t work for *-emu, i386-qemu, i386-coreboot, i386-multiboot, mips loongson, mips-arc and mips qemu mips

To generate a netbootable directory, run:

grub-mknetdir --net-directory=/srv/tftp --subdir=/boot/grub -d /usr/lib/grub/<platform>

E.g. for i386-pc:

grub-mknetdir --net-directory=/srv/tftp --subdir=/boot/grub -d /usr/lib/grub/i386-pc

Then follow instructions printed out by grub-mknetdir on configuring your DHCP server.

The grub.cfg file is placed in the same directory as the path output by grub-mknetdir hereafter referred to as FWPATH. GRUB will search for its configuration files in order using the following rules where the appended value corresponds to a value on the client machine.

`(FWPATH)'/grub.cfg-`(UUID OF MACHINE)'
`(FWPATH)'/grub.cfg-`(MAC ADDRESS OF NIC)'
`(FWPATH)'/grub.cfg-`(IPv4 OR IPv6 ADDRESS)'
`(FWPATH)'/grub.cfg

The UUID is the Client Machine Identifier Option Definition as specified in RFC 4578. The client will only attempt to loouk up a UUID config file if it was provided by the DHCP server.

The client will only attempt to look up an IPv6 address config once, however, it will try the IPv4 multiple times. The concrete example below shows what would happen under the IPv4 case.

UUID: 7726a678-7fc0-4853-a4f6-c85ac36a120a
MAC:  52:54:00:ec:33:81
IPV4: 10.0.0.130 (0A000082)
`(FWPATH)'/grub.cfg-7726a678-7fc0-4853-a4f6-c85ac36a120a
`(FWPATH)'/grub.cfg-52-54-00-ec-33-81
`(FWPATH)'/grub.cfg-0A000082
`(FWPATH)'/grub.cfg-0A00008
`(FWPATH)'/grub.cfg-0A0000
`(FWPATH)'/grub.cfg-0A000
`(FWPATH)'/grub.cfg-0A00
`(FWPATH)'/grub.cfg-0A0
`(FWPATH)'/grub.cfg-0A
`(FWPATH)'/grub.cfg-0
`(FWPATH)'/grub.cfg

This feature is enabled by default but it can be disabled by setting the ‘feature_net_search_cfg’ to ‘n’. Since this happens before the configuration file is read by GRUB, this option has to be disabled in an embedded configuration file (see Section 6.5 [Embedded configuration], page 33).

After GRUB has started, files on the TFTP server will be accessible via the ‘(tftp)’ device.

The server IP address can be controlled by changing the ‘(tftp)’ device name to ‘(tftp,server-ip)’. Note that this should be changed both in the prefix and in any references to the device name in the configuration file.

GRUB provides several environment variables which may be used to inspect or change the behaviour of the PXE device. In the following description <interface> is placeholder for the name of network interface (platform dependent):

‘net_<interface>_ip’

The network interface’s IP address. Read-only.

‘net_<interface>_mac’

The network interface’s MAC address. Read-only.

‘net_<interface>_clientid’

The client id provided by DHCP. Read-only.

‘net_<interface>_clientuuid’

The client uuid provided by DHCP. Read-only.

‘net_<interface>_hostname’

The client host name provided by DHCP. Read-only.

‘net_<interface>_domain’

The client domain name provided by DHCP. Read-only.

‘net_<interface>_rootpath’

The path to the client’s root disk provided by DHCP. Read-only.

‘net_<interface>_extensionspath’

The path to additional DHCP vendor extensions provided by DHCP. Read-only.

‘net_<interface>_boot_file’

The boot file name provided by DHCP. Read-only.

‘net_<interface>_dhcp_server_name’

The name of the DHCP server responsible for these boot parameters. Readonly.

‘net_<interface>_next_server’

The IP address of the next (usually, TFTP) server provided by DHCP. Readonly.

‘net_default_interface’

Initially set to name of network interface that was used to load grub. Readwrite, although setting it affects only interpretation of ‘net_default_ip’ and

‘net_default_mac’
‘net_default_ip’

The IP address of default interface. Read-only. This is alias for the

‘net_${net_default_interface}_ip’.
‘net_default_mac’

The default interface’s MAC address. Read-only. This is alias for the

‘net_${net_default_interface}_mac’.

‘net_default_server’

The default server used by network drives (see Section 13.1 [Device syntax], page 57). Read-write, although setting this is only useful before opening a network device.

9 Using GRUB via a serial line

This chapter describes how to use the serial terminal support in GRUB.

If you have many computers or computers with no display/keyboard, it could be very useful to control the computers through serial communications. To connect one computer with another via a serial line, you need to prepare a null-modem (cross) serial cable, and you may need to have multiport serial boards, if your computer doesn’t have extra serial ports. In addition, a terminal emulator is also required, such as minicom. Refer to a manual of your operating system, for more information.

As for GRUB, the instruction to set up a serial terminal is quite simple. Here is an example:

     grub> serial --unit=0 --speed=9600
     grub> terminal_input serial; terminal_output serial

The command serial initializes the serial unit 0 with the speed 9600bps. The serial unit 0 is usually called ‘COM1’, so, if you want to use COM2, you must specify ‘--unit=1’ instead. This command accepts many other options, see Section 17.3.1 [serial], page 77 for more details.

Without argument or with ‘--port=auto’, GRUB will attempt to use ACPI when available to auto-detect the default serial port and its configuration.

The commands terminal_input (see Section 17.3.2 [terminal input], page 78) and terminal_output (see Section 17.3.3 [terminal output], page 78) choose which type of terminal you want to use. In the case above, the terminal will be a serial terminal, but you can also pass console to the command, as ‘terminal_input serial console’. In this case, a terminal in which you press any key will be selected as a GRUB terminal. In the example above, note that you need to put both commands on the same command line, as you will lose the ability to type commands on the console after the first command.

However, note that GRUB assumes that your terminal emulator is compatible with VT100 by default. This is true for most terminal emulators nowadays, but you should pass the option ‘--dumb’ to the command if your terminal emulator is not VT100-compatible or implements few VT100 escape sequences. If you specify this option then GRUB provides you with an alternative menu interface, because the normal menu requires several fancy features of your terminal.

10 Using GRUB with vendor power-on keys

Some laptop vendors provide an additional power-on button which boots another OS. GRUB

supports such buttons with the ‘GRUB_TIMEOUT_BUTTON’, ‘GRUB_TIMEOUT_STYLE_BUTTON’,
‘GRUB_DEFAULT_BUTTON’, and ‘GRUB_BUTTON_CMOS_ADDRESS’ variables in default/grub

(see Section 6.1 [Simple configuration], page 21). ‘GRUB_TIMEOUT_BUTTON’,

‘GRUB_TIMEOUT_STYLE_BUTTON’, and ‘GRUB_DEFAULT_BUTTON’ are used instead of the

corresponding variables without the ‘_BUTTON’ suffix when powered on using the special button. ‘GRUB_BUTTON_CMOS_ADDRESS’ is vendor-specific and partially model-specific. Values known to the GRUB team are:

DELL XPS M1330M 121:3

DELL XPS M1530 85:3

DELL LATITUDE E4300 85:3

ASUS EEEPC 1005PE 84:1 (unconfirmed)

LENOVO THINKPAD T410S (2912W1C)

101:3

To take full advantage of this function, install GRUB into the MBR (see Section 4.1 [Installing GRUB using grub-install], page 11).

If you have a laptop which has a similar feature and not in the above list could you figure your address and contribute? To discover the address do the following:

• boot normally

•

         sudo modprobe nvram
         sudo cat /dev/nvram | xxd > normal_button.txt

• boot using vendor button •

         sudo modprobe nvram
         sudo cat /dev/nvram | xxd > normal_vendor.txt

Then compare these text files and find where a bit was toggled. E.g. in case of Dell XPS it was:

     byte 0x47: 20 --> 28

It’s a bit number 3 as seen from following table:

0  01
1  02
2  04
3  08
4  10

5  20
6  40
7  80

0x47 is decimal 71. Linux nvram implementation cuts first 14 bytes of CMOS. So the real byte address in CMOS is 71+14=85 So complete address is 85:3

11 GRUB image files

GRUB consists of several images: a variety of bootstrap images for starting GRUB in various ways, a kernel image, and a set of modules which are combined with the kernel image to form a core image. Here is a short overview of them.

‘boot.img’

On PC BIOS systems, this image is the first part of GRUB to start. It is written to a master boot record (MBR) or to the boot sector of a partition. Because a PC boot sector is 512 bytes, the size of this image is exactly 512 bytes. The sole function of ‘boot.img’ is to read the first sector of the core image from a local disk and jump to it. Because of the size restriction, ‘boot.img’ cannot understand any file system structure, so grub-install hardcodes the location of the first sector of the core image into ‘boot.img’ when installing GRUB.

‘diskboot.img’

This image is used as the first sector of the core image when booting from a hard disk. It reads the rest of the core image into memory and starts the kernel. Since file system handling is not yet available, it encodes the location of the core image using a block list format.

‘cdboot.img’

This image is used as the first sector of the core image when booting from a CD-ROM drive. It performs a similar function to ‘diskboot.img’.

‘pxeboot.img’

This image is used as the start of the core image when booting from the network using PXE. See Chapter 8 [Network], page 45.

‘lnxboot.img’

This image may be placed at the start of the core image in order to make GRUB look enough like a Linux kernel that it can be booted by LILO using an ‘image=’ section.

‘kernel.img’

This image contains GRUB’s basic run-time facilities: frameworks for device and file handling, environment variables, the rescue mode command-line parser, and so on. It is rarely used directly, but is built into all core images.

‘core.img’

This is the core image of GRUB. It is built dynamically from the kernel image and an arbitrary list of modules by the grub-mkimage program. Usually, it contains enough modules to access ‘/boot/grub’, and loads everything else (including menu handling, the ability to load target operating systems, and so on) from the file system at run-time. The modular design allows the core image to be kept small, since the areas of disk where it must be installed are often as small as 32KB. See Section 4.4 [BIOS installation], page 13, for details on where the core image can be installed on PC systems.

‘*.mod’ Everything else in GRUB resides in dynamically loadable modules. These are often loaded automatically, or built into the core image if they are essential, but may also be loaded manually using the insmod command (see Section 17.4.37 [insmod], page 87).

For GRUB Legacy users

GRUB 2 has a different design from GRUB Legacy, and so correspondences with the images it used cannot be exact. Nevertheless, GRUB Legacy users often ask questions in the terms they are familiar with, and so here is a brief guide to how GRUB 2’s images relate to that.

‘stage1’ Stage 1 from GRUB Legacy was very similar to ‘boot.img’ in GRUB 2, and they serve the same function.

‘*_stage1_5’

In GRUB Legacy, Stage 1.5’s function was to include enough filesystem code to allow the much larger Stage 2 to be read from an ordinary filesystem. In this respect, its function was similar to ‘core.img’ in GRUB 2. However, ‘core.img’ is much more capable than Stage 1.5 was; since it offers a rescue shell, it is sometimes possible to recover manually in the event that it is unable to load any other modules, for example if partition numbers have changed. ‘core.img’ is built in a more flexible way, allowing GRUB 2 to support reading modules from advanced disk types such as LVM and RAID. GRUB Legacy could run with only Stage 1 and Stage 2 in some limited configurations, while GRUB 2 requires ‘core.img’ and cannot work without it.

‘stage2’ GRUB 2 has no single Stage 2 image. Instead, it loads modules from ‘/boot/grub’ at run-time.

‘stage2_eltorito’

In GRUB 2, images for booting from CD-ROM drives are now constructed using ‘cdboot.img’ and ‘core.img’, making sure that the core image contains the ‘iso9660’ module. It is usually best to use the grub-mkrescue program for this.

‘nbgrub’ There is as yet no equivalent for ‘nbgrub’ in GRUB 2; it was used by Etherboot and some other network boot loaders.

‘pxegrub’ In GRUB 2, images for PXE network booting are now constructed using ‘pxeboot.img’ and ‘core.img’, making sure that the core image contains the ‘pxe’ and ‘pxecmd’ modules. See Chapter 8 [Network], page 45.

12 Core image size limitation

Heavily limited platforms:

• i386-pc (normal and PXE): the core image size (compressed) is limited by 458240 bytes. kernel.img (.text + .data + .bss, uncompressed) is limited by 392704 bytes. module size (uncompressed) + kernel.img (.text + .data, uncompressed) is limited by the size of contiguous chunk at 1M address.

• sparc64-ieee1275: kernel.img (.text + .data + .bss) + modules + 256K (stack) + 2M (heap) is limited by space available at 0x4400. On most platforms it’s just 3 or 4M since ieee1275 maps only so much.

• i386-ieee1275: kernel.img (.text + .data + .bss) + modules is limited by memory available at 0x10000, at most 596K

Lightly limited platforms:

• *-xen: limited only by adress space and RAM size.

• i386-qemu: kernel.img (.text + .data + .bss) is limited by 392704 bytes. (core.img would be limited by ROM size but it’s unlimited on qemu

• All EFI platforms: limited by contiguous RAM size and possibly firmware bugs

• Coreboot and multiboot. kernel.img (.text + .data + .bss) is limited by 392704 bytes. module size is limited by the size of contiguous chunk at 1M address.

• mipsel-loongson (ELF), mips(el)-qemu mips (ELF): if uncompressed: kernel.img (.text + .data) + modules is limited by the space from 80200000 forward if compressed: kernel.img (.text + .data, uncompressed) + modules (uncompressed) + (modules + kernel.img (.text + .data)) (compressed) + decompressor is limited by the space from 80200000 forward

• mipsel-loongson (Flash), mips(el)-qemu mips (Flash): kernel.img (.text + .data) + modules is limited by the space from 80200000 forward core.img (final) is limited by flash size (512K on yeeloong and fulooong)

• mips-arc: if uncompressed: kernel.img (.text + .data) is limited by the space from 8bd00000 forward modules + dummy decompressor is limited by the space from 8bd00000 backward if compressed: kernel.img (.text + .data, uncompressed) is limited by the space from 8bd00000 forward modules (uncompressed) + (modules + kernel.img (.text + .data)) (compressed, aligned to 1M) + 1M (decompressor + scratch space) is limited by the space from 8bd00000 backward

• powerpc-ieee1275: kernel.img (.text + .data + .bss) + modules is limited by space available at 0x200000

13 Filesystem syntax and semantics

GRUB uses a special syntax for specifying disk drives which can be accessed by BIOS. Because of BIOS limitations, GRUB cannot distinguish between IDE, ESDI, SCSI, or others. You must know yourself which BIOS device is equivalent to which OS device. Normally, that will be clear if you see the files in a device or use the command search (see Section 17.4.65 [search], page 93).

13.1 How to specify devices

The device syntax is like this:

(device[,partmap-name1part-num1[,partmap-name2part-num2[,...]]])

‘[]’ means the parameter is optional. device depends on the disk driver in use. BIOS and EFI disks use either ‘fd’ or ‘hd’ followed by a digit, like ‘fd0’, or ‘cd’. AHCI, PATA (ata), crypto, USB use the name of driver followed by a number. Memdisk and host are limited to one disk and so it’s refered just by driver name. RAID (md), ofdisk (ieee1275 and nand), LVM (lvm), LDM, virtio (vdsk) and arcdisk (arc) use intrinsic name of disk prefixed by driver name. Additionally just “nand” refers to the disk aliased as “nand”. Conflicts are solved by suffixing a number if necessarry. Commas need to be escaped. Loopback uses whatever name specified to loopback command. Hostdisk uses names specified in device.map as long as it’s of the form [fhc]d[0-9]* or hostdisk/<OS DEVICE>. For crypto and RAID (md) additionally you can use the syntax <driver name>uuid/<uuid>. For LVM additionally you can use the syntax lvmid/<volume-group-uuid>/<volume-uuid>.

(fd0)
(hd0)
(cd)
(ahci0)
(ata0)
(crypto0)
(usb0)
(cryptouuid/123456789abcdef0123456789abcdef0)
(mduuid/123456789abcdef0123456789abcdef0)
(lvm/system-root)
(lvmid/F1ikgD-2RES-306G-il9M-7iwa-4NKW-EbV1NV/eLGuCQ-L4Ka-XUgR-sjtJ-ffch-bajr-fCNfz5)
(md/myraid)
(md/0)
(ieee1275/disk2)
(ieee1275//pci@1f\,0/ide@d/disk@2)
(nand)
(memdisk)
(host)
(myloop)
(hostdisk//dev/sda)

part-num represents the partition number of device, starting from one. partname is optional but is recommended since disk may have several top-level partmaps. Specifying third and later component you can access to subpartitions.

The syntax ‘(hd0)’ represents using the entire disk (or the MBR when installing GRUB), while the syntax ‘(hd0,1)’ represents using the first partition of the disk (or the boot sector of the partition when installing GRUB).

     (hd0,msdos1)
     (hd0,msdos1,msdos5)
     (hd0,msdos1,bsd3)
     (hd0,netbsd1)
     (hd0,gpt1)
     (hd0,1,3)

If you enabled the network support, the special drives (protocol[,server]) are also available. Supported protocols are ‘http’ and ‘tftp’. If server is omitted, value of environment variable ‘net_default_server’ is used. Before using the network drive, you must initialize the network. See Chapter 8 [Network], page 45, for more information.

When using ‘http’ or ‘tftp’, ports other than ‘80’ can be specified using a colon (‘:’) after the address. To avoid parsing conflicts, when using IPv6 addresses with custom ports, the addresses must be enclosed with square brackets (‘[]’), as is standard practice.

     (http,grub.example.com:31337)
     (http,192.0.2.1:339)
     (http,[2001:db8::1]:11235)

If you boot GRUB from a CD-ROM, ‘(cd)’ is available. See Section 4.2 [Making a GRUB bootable CD-ROM], page 12, for details.

13.2 How to specify files

There are two ways to specify files, by absolute file name and by block list.

An absolute file name resembles a Unix absolute file name, using ‘/’ for the directory separator (not ‘\’ as in DOS). One example is ‘(hd0,1)/boot/grub/grub.cfg’. This means the file ‘/boot/grub/grub.cfg’ in the first partition of the first hard disk. If you omit the device name in an absolute file name, GRUB uses GRUB’s root device implicitly. So if you set the root device to, say, ‘(hd1,1)’ by the command ‘set root=(hd1,1)’ (see Section 17.4.67 [set], page 96), then /boot/kernel is the same as (hd1,1)/boot/kernel.

On ZFS filesystem the first path component must be volume‘@’[snapshot]. So

‘/rootvol@snap-129/boot/grub/grub.cfg’ refers to file ‘/boot/grub/grub.cfg’ in

snapshot of volume ‘rootvol’ with name ‘snap-129’. Trailing ‘@’ after volume name is mandatory even if snapshot name is omitted.

13.3 How to specify block lists

A block list is used for specifying a file that doesn’t appear in the filesystem, like a chain-

loader. The syntax is [offset]+[length][,[offset]+[length]].... Here is an exam-

ple:

     0+100,200+1,300+300,800+

This represents that GRUB should read blocks 0 through 99, block 200, blocks 300 through 599, and blocks 800 until the end of the device. If you omit an offset, then GRUB assumes the offset is zero. If the length is omitted, then GRUB assumes the block list extends until the end of the device.

Like the file name syntax (see Section 13.2 [File name syntax], page 58), if a blocklist does not contain a device name, then GRUB uses GRUB’s root device. So (hd0,2)+1 is the same as +1 when the root device is ‘(hd0,2)’.

14 GRUB’s user interface

GRUB has both a simple menu interface for choosing preset entries from a configuration file, and a highly flexible command-line for performing any desired combination of boot commands.

GRUB looks for its configuration file as soon as it is loaded. If one is found, then the full menu interface is activated using whatever entries were found in the file. If you choose the command-line menu option, or if the configuration file was not found, then GRUB drops to the command-line interface.

14.1 The flexible command-line interface

The command-line interface provides a prompt and after it an editable text area much like a command-line in Unix or DOS. Each command is immediately executed after it is entered^{1}. The commands (see Chapter 17 [Commands], page 73) are a subset of those available in the configuration file, used with exactly the same syntax.

Cursor movement and editing of the text on the line can be done via a subset of the functions available in the Bash shell:

C-F PC RIGHT KEY

Move forward one character.

C-B PC LEFT KEY

Move back one character.

C-A HOME Move to the start of the line.

C-E END Move the the end of the line.

C-D DEL Delete the character underneath the cursor.

C-H BS Delete the character to the left of the cursor.

C-K Kill the text from the current cursor position to the end of the line.

C-U Kill backward from the cursor to the beginning of the line.

C-Y Yank the killed text back into the buffer at the cursor.

C-P PC UP KEY

Move up through the history list.

C-N PC DOWN KEY

Move down through the history list.

1 However, this behavior will be changed in the future version, in a user-invisible way.

When typing commands interactively, if the cursor is within or before the first word in the command-line, pressing the TAB key (or C-I) will display a listing of the available commands, and if the cursor is after the first word, the TAB will provide a completion listing of disks, partitions, and file names depending on the context. Note that to obtain a list of drives, one must open a parenthesis, as root (.

Note that you cannot use the completion functionality in the TFTP filesystem. This is because TFTP doesn’t support file name listing for the security.

14.2 The simple menu interface

The menu interface is quite easy to use. Its commands are both reasonably intuitive and described on screen.

Basically, the menu interface provides a list of boot entries to the user to choose from. Use the arrow keys to select the entry of choice, then press RET to run it. An optional timeout is available to boot the default entry (the first one if not set), which is aborted by pressing any key.

Commands are available to enter a bare command-line by pressing C (which operates exactly like the non-config-file version of GRUB, but allows one to return to the menu if desired by pressing ESC) or to edit any of the boot entries by pressing E.

If you protect the menu interface with a password (see Chapter 19 [Security], page 111), all you can do is choose an entry by pressing RET, or press P to enter the password.

Pressing CTRL-L will refresh the menu, which can be useful when connecting via serial after the menu has been drawn.

14.3 Editing a menu entry

The menu entry editor looks much like the main menu interface, but the lines in the menu are individual commands in the selected entry instead of entry names.

If an ESC is pressed in the editor, it aborts all the changes made to the configuration entry and returns to the main menu interface.

Each line in the menu entry can be edited freely, and you can add new lines by pressing RET at the end of a line. To boot the edited entry, press CTRL-X.

Although GRUB unfortunately does not support undo, you can do almost the same thing by just returning to the main menu using ESC.

15 GRUB environment variables

GRUB supports environment variables which are rather like those offered by all Unixlike systems. Environment variables have a name, which is unique and is usually a short identifier, and a value, which is an arbitrary string of characters. They may be set (see Section 17.4.67 [set], page 96), unset (see Section 17.4.77 [unset], page 100), or looked up (see Section 6.3 [Shell-like scripting], page 28) by name.

A number of environment variables have special meanings to various parts of GRUB. Others may be used freely in GRUB configuration files.

15.1 Special environment variables

These variables have special meaning to GRUB.

15.1.1 biosnum

When chain-loading another boot loader (see Section 5.1.3 [Chain-loading], page 15), GRUB may need to know what BIOS drive number corresponds to the root device (see Section 15.1.43 [root], page 69) so that it can set up registers properly. If the biosnum variable is set, it overrides GRUB’s own means of guessing this.

For an alternative approach which also changes BIOS drive mappings for the chainloaded system, see Section 17.4.23 [drivemap], page 83.

15.1.2 check signatures

This variable controls whether GRUB enforces digital signature validation on loaded files. See Section 19.2 [Using digital signatures], page 112.

15.1.3 chosen

When executing a menu entry, GRUB sets the chosen variable to the title of the entry being executed.

If the menu entry is in one or more submenus, then chosen is set to the titles of each of the submenus starting from the top level followed by the title of the menu entry itself, separated by ‘>’.

15.1.4 cmdpath

The location from which ‘core.img’ was loaded as an absolute directory name (see Section 13.2 [File name syntax], page 58). This is set by GRUB at startup based on information returned by platform firmware. Not every platform provides this information and some may return only device without path name.

15.1.5 color highlight

This variable contains the “highlight” foreground and background terminal colors, separated by a slash (‘/’). Setting this variable changes those colors. For the available color names, see Section 15.1.6 [color normal], page 64.

      The default is ‘black/light-gray’.

15.1.6 color normal

This variable contains the “normal” foreground and background terminal colors, separated by a slash (‘/’). Setting this variable changes those colors. Each color must be a name from the following list:

• black

• blue

• green

• cyan

• red

• magenta

• brown

• light-gray

• dark-gray

• light-blue

• light-green

• light-cyan

• light-red

• light-magenta

• yellow

• white

    The default is ‘light-gray/black’.

The color support support varies from terminal to terminal.

‘morse’ has no color support at all.

‘mda_text’ color support is limited to highlighting by black/white reversal.

‘console’ on ARC, EMU and IEEE1275, ‘serial_*’ and ‘spkmodem’ are governed by terminfo and support only 8 colors if in modes ‘vt100-color’ (default for console on emu), ‘arc’ (default for console on ARC), ‘ieee1275’ (default for console on IEEE1275). When in mode ‘vt100’ then the color support is limited to highlighting by black/white reversal. When in mode ‘dumb’ there is no color support.

When console supports no colors this setting is ignored. When console supports 8 colors, then the colors from the second half of the previous list are mapped to the matching colors of first half.

‘console’ on EFI and BIOS and ‘vga_text’ support all 16 colors.

‘gfxterm’ supports all 16 colors and would be theoretically extendable to support whole rgb24 palette but currently there is no compelling reason to go beyond the current 16 colors.

15.1.7 config directory

This variable is automatically set by GRUB to the directory part of current configuration file name (see Section 15.1.8 [config file], page 65).

15.1.8 config file

This variable is automatically set by GRUB to the name of configuration file that is being processed by commands configfile (see Section 17.4.15 [configfile], page 81) or normal (see Section 17.4.51 [normal], page 89). It is restored to the previous value when command completes.

15.1.9 debug

This variable may be set to enable debugging output from various components of GRUB. The value is an ordered list of debug facility names separated by whitespace or ‘,’. If the special facility named ‘all’ is present then debugging output of all facility names is enabled at the start of processing the value of this variable. A facility’s debug output can then be disabled by prefixing its name with a ‘-’. The last occurence facility name with or without a leading ‘-’ takes precendent over any previous occurence. This allows the easy enabling or disabling of facilities by appending a ‘,’ and then the facility name with or without the leading ‘-’, which will preserve the state of the rest of the facilities. The facility names are the first argument to grub dprintf. Consult the source for more details.

15.1.10 default

If this variable is set, it identifies a menu entry that should be selected by default, possibly after a timeout (see Section 15.1.47 [timeout], page 69). The entry may be identified by number (starting from 0 at each level of the hierarchy), by title, or by id.

For example, if you have:

menuentry 'Example GNU/Linux distribution' --class gnu-linux --id example-gnu-linux {
        ...
}

then you can make this the default using:

     default=example-gnu-linux

If the entry is in a submenu, then it must be identified using the number, title, or id of each of the submenus starting from the top level, followed by the number, title, or id of the menu entry itself, with each element separated by ‘>’. For example, take the following menu structure:

     GNU/Hurd --id gnu-hurd
       Standard Boot --id=gnu-hurd-std
       Rescue shell --id=gnu-hurd-rescue
     Other platforms --id=other
       Minix --id=minix
         Version 3.4.0 --id=minix-3.4.0
         Version 3.3.0 --id=minix-3.3.0
       GRUB Invaders --id=grub-invaders

The more recent release of Minix would then be identified as ‘Other

platforms>Minix>Version 3.4.0’, or as ‘1>0>0’, or as ‘other>minix>minix-3.4.0’.

This variable is often set by ‘GRUB_DEFAULT’ (see Section 6.1 [Simple configuration],

page 21), grub-set-default, or grub-reboot.

15.1.11 fallback

If this variable is set, it identifies a menu entry that should be selected if the default menu entry fails to boot. Entries are identified in the same way as for ‘default’ (see Section 15.1.10 [default], page 65).

15.1.12 gfxmode

If this variable is set, it sets the resolution used on the ‘gfxterm’ graphical terminal. Note that you can only use modes which your graphics card supports via VESA BIOS Extensions (VBE), so for example native LCD panel resolutions may not be available. The default is ‘auto’, which selects a platform-specific default that should look reasonable. Supported modes can be listed by ‘videoinfo’ command in GRUB.

The resolution may be specified as a sequence of one or more modes, separated by commas (‘,’) or semicolons (‘;’); each will be tried in turn until one is found. Each mode

should be either ‘auto’, ‘widthxheight’, or ‘widthxheightxdepth’.

15.1.13 gfxpayload

If this variable is set, it controls the video mode in which the Linux kernel starts up, replacing the ‘vga=’ boot option (see Section 17.2.4 [linux], page 76). It may be set to ‘text’ to force the Linux kernel to boot in normal text mode, ‘keep’ to preserve the graphics mode set using ‘gfxmode’, or any of the permitted values for ‘gfxmode’ to set a particular graphics mode (see Section 15.1.12 [gfxmode], page 66).

Depending on your kernel, your distribution, your graphics card, and the phase of the moon, note that using this option may cause GNU/Linux to suffer from various display problems, particularly during the early part of the boot sequence. If you have problems, set this variable to ‘text’ and GRUB will tell Linux to boot in normal text mode.

The default is platform-specific. On platforms with a native text mode (such as PC BIOS platforms), the default is ‘text’. Otherwise the default may be ‘auto’ or a specific video mode.

This variable is often set by ‘GRUB_GFXPAYLOAD_LINUX’ (see Section 6.1 [Simple configuration], page 21).

15.1.14 gfxterm font

If this variable is set, it names a font to use for text on the ‘gfxterm’ graphical terminal. Otherwise, ‘gfxterm’ may use any available font.

15.1.15 grub cpu

In normal mode (see Section 17.4.51 [normal], page 89), GRUB sets the ‘grub_cpu’ variable to the CPU type for which GRUB was built (e.g. ‘i386’ or ‘powerpc’).

15.1.16 grub platform

In normal mode (see Section 17.4.51 [normal], page 89), GRUB sets the ‘grub_platform’ variable to the platform for which GRUB was built (e.g. ‘pc’ or ‘efi’).

15.1.17 icondir

If this variable is set, it names a directory in which the GRUB graphical menu should look for icons after looking in the theme’s ‘icons’ directory. See Chapter 7 [Theme file format], page 35.

15.1.18 lang

If this variable is set, it names the language code that the gettext command (see Section 17.4.31 [gettext], page 85) uses to translate strings. For example, French would be named as ‘fr’, and Simplified Chinese as ‘zh_CN’.

grub-mkconfig (see Section 6.1 [Simple configuration], page 21) will try to set a reasonable default for this variable based on the system locale.

15.1.19 locale dir

If this variable is set, it names the directory where translation files may be found (see Section 17.4.31 [gettext], page 85), usually ‘/boot/grub/locale’. Otherwise, internationalization is disabled.

grub-mkconfig (see Section 6.1 [Simple configuration], page 21) will set a reasonable default for this variable if internationalization is needed and any translation files are available.

15.1.20 lockdown

If this variable is set to ‘y’, it means that GRUB has entered see Section 19.6 [Lockdown], page 114 mode.

15.1.21 menu color highlight

This variable contains the foreground and background colors to be used for the highlighted menu entry, separated by a slash (‘/’). Setting this variable changes those colors. For the available color names, see Section 15.1.6 [color normal], page 64.

The default is the value of ‘color_highlight’ (see Section 15.1.5 [color highlight], page 63).

15.1.22 menu color normal

This variable contains the foreground and background colors to be used for non-highlighted menu entries, separated by a slash (‘/’). Setting this variable changes those colors. For the available color names, see Section 15.1.6 [color normal], page 64.

The default is the value of ‘color_normal’ (see Section 15.1.6 [color normal], page 64).

15.1.23 net <interface> boot file

See Chapter 8 [Network], page 45.

15.1.24 net <interface> clientid

See Chapter 8 [Network], page 45.

15.1.25 net <interface> clientuuid

See Chapter 8 [Network], page 45.

15.1.26 net <interface> dhcp server name

See Chapter 8 [Network], page 45.

15.1.27 net <interface> domain

See Chapter 8 [Network], page 45.

15.1.28 net <interface> extensionspath

See Chapter 8 [Network], page 45.

15.1.29 net <interface> hostname

See Chapter 8 [Network], page 45.

15.1.30 net <interface> ip

See Chapter 8 [Network], page 45.

15.1.31 net <interface> mac

See Chapter 8 [Network], page 45.

15.1.32 net <interface> next server

See Chapter 8 [Network], page 45.

15.1.33 net <interface> rootpath

See Chapter 8 [Network], page 45.

15.1.34 net default interface

See Chapter 8 [Network], page 45.

15.1.35 net default ip

See Chapter 8 [Network], page 45.

15.1.36 net default mac

See Chapter 8 [Network], page 45.

15.1.37 net default server

See Chapter 8 [Network], page 45.

15.1.38 pager

If set to ‘1’, pause output after each screenful and wait for keyboard input. The default is not to pause output.

15.1.39 prefix

The location of the ‘/boot/grub’ directory as an absolute file name (see Section 13.2 [File name syntax], page 58). This is normally set by GRUB at startup based on information provided by grub-install. GRUB modules are dynamically loaded from this directory, so it must be set correctly in order for many parts of GRUB to work.

15.1.40 pxe blksize

See Chapter 8 [Network], page 45.

15.1.41 pxe default gateway

See Chapter 8 [Network], page 45.

15.1.42 pxe default server

See Chapter 8 [Network], page 45.

15.1.43 root

The root device name (see Section 13.1 [Device syntax], page 57). Any file names that do not specify an explicit device name are read from this device. The default is normally set by GRUB at startup based on the value of ‘prefix’ (see Section 15.1.39 [prefix], page 69).

For example, if GRUB was installed to the first partition of the first hard disk, then

‘prefix’ might be set to ‘(hd0,msdos1)/boot/grub’ and ‘root’ to ‘hd0,msdos1’.

15.1.44 shim lock

If this variable is set to ‘y’, it means that the shim lock verifier is registered (see see Section 19.3 [UEFI secure boot and shim], page 113).

15.1.45 superusers

This variable may be set to a list of superuser names to enable authentication support. See Chapter 19 [Security], page 111.

15.1.46 theme

This variable may be set to a directory containing a GRUB graphical menu theme. See Chapter 7 [Theme file format], page 35.

This variable is often set by ‘GRUB_THEME’ (see Section 6.1 [Simple configuration], page 21).

15.1.47 timeout

If this variable is set, it specifies the time in seconds to wait for keyboard input before booting the default menu entry. A timeout of ‘0’ means to boot the default entry immediately without displaying the menu; a timeout of ‘-1’ (or unset) means to wait indefinitely.

If ‘timeout_style’ (see Section 15.1.48 [timeout style], page 70) is set to ‘countdown’ or ‘hidden’, the timeout is instead counted before the menu is displayed.

This variable is often set by ‘GRUB_TIMEOUT’ (see Section 6.1 [Simple configuration], page 21).

15.1.48 timeout style

This variable may be set to ‘menu’, ‘countdown’, or ‘hidden’ to control the way in which the timeout (see Section 15.1.47 [timeout], page 69) interacts with displaying the menu. See the documentation of ‘GRUB_TIMEOUT_STYLE’ (see Section 6.1 [Simple configuration], page 21) for details.

15.1.49 tpm fail fatal

If this variable is set and true (i.e., not set to “0”, “false”, “disable”, or “no”), TPM measurements that fail will be treated as fatal. Otherwise, they will merely be debuglogged and boot will continue.

Call to EFI firmware, like hash log extend event(), can return an unknown error, i.e. due to bug present in firmware. When this variable is set and true (same values as with TPM measurements) this situation will be considered to be fatal and error-logged as “unknown TPM error”. If not set, booting the OS will be enabled.

15.2 The GRUB environment block

It is often useful to be able to remember a small amount of information from one boot to the next. For example, you might want to set the default menu entry based on what was selected the last time. GRUB deliberately does not implement support for writing files in order to minimise the possibility of the boot loader being responsible for file system corruption, so a GRUB configuration file cannot just create a file in the ordinary way. However, GRUB provides an “environment block” which can be used to save a small amount of state.

The environment block is a preallocated 1024-byte file, which normally lives in ‘/boot/grub/grubenv’ (although you should not assume this). At boot time, the load_ env command (see Section 17.4.41 [load env], page 87) loads environment variables from it, and the save_env (see Section 17.4.64 [save env], page 93) command saves environment variables to it. From a running system, the grub-editenv utility can be used to edit the environment block.

For safety reasons, this storage is only available when installed on a plain disk (no LVM or RAID), using a non-checksumming filesystem (no ZFS), and using BIOS or EFI functions (no ATA, USB or IEEE1275).

grub-mkconfig uses this facility to implement ‘GRUB_SAVEDEFAULT’ (see Section 6.1 [Simple configuration], page 21).

16 Modules

In this chapter, we list all modules that are available in GRUB (currently incomplete).

Modules can be loaded via the insmod (see Section 17.4.37 [insmod], page 87) command.

16.1 bli

This module provides basic support for the Boot Loader Interface. The Boot Loader Interface specifies a set of EFI variables that are used to communicate boot-time information between the bootloader and the operating system.

The following variables are placed under the vendor UUID 4a67b082-0a4c-41cf-b6c7-440b29bb8c4f when the module is loaded:

The GPT partition UUID of the EFI System Partition used during boot is published via the LoaderDevicePartUUID variable. The Boot Loader Interface specification requires GPT formatted drives. The bli module ignores drives/partitions in any other format. If GRUB is loaded from a non-GPT partition, e.g. from an MSDOS formatted drive or network, this variable will not be set.

A string identifying GRUB as the active bootloader including the version number is

stored in LoaderInfo.

This module is only available on UEFI platforms.

17 Available commands

In this chapter, we list all commands that are available in GRUB.

Commands belong to different groups. A few can only be used in the global section of the configuration file (or “menu”); most of them can be entered on the command-line and can be used either anywhere in the menu or specifically in the menu entries.

In rescue mode, only the insmod (see Section 17.4.37 [insmod], page 87), ls (see Section 17.4.44 [ls], page 88), set (see Section 17.4.67 [set], page 96), and unset (see Section 17.4.77 [unset], page 100) commands are normally available. If you end up in rescue mode and do not know what to do, then see Section 23.1 [GRUB only offers a rescue shell], page 127.

17.1 Commands for the menu only

The semantics used in parsing the configuration file are the following:

• The files must be in plain-text format.

• ‘#’ at the beginning of a line in a configuration file means it is only a comment.

• Options are separated by spaces.

• All numbers can be either decimal or hexadecimal. A hexadecimal number must be preceded by ‘0x’, and is case-insensitive.

These commands can only be used in the menu:

17.1.1 menuentry

[Command] menuentry title [‘--class=class’ . . .] [‘--users=users’]
   [‘--unrestricted’] [‘--hotkey=key’] [‘--id=id’] [arg . . .] { command;

… }

This defines a GRUB menu entry named title. When this entry is selected from the menu, GRUB will set the chosen environment variable to value of ‘--id’ if ‘--id’ is given, execute the list of commands given within braces, and if the last command in the list returned successfully and a kernel was loaded it will execute the boot command. The ‘--class’ option may be used any number of times to group menu entries into classes. Menu themes may display different classes using different styles. The ‘--users’ option grants specific users access to specific menu entries. See Chapter 19 [Security], page 111. The ‘--unrestricted’ option grants all users access to specific menu entries. See Chapter 19 [Security], page 111. The ‘--hotkey’ option associates a hotkey with a menu entry. key may be a single letter, or one of the aliases ‘backspace’, ‘tab’, or ‘delete’. The ‘--id’ may be used to associate unique identifier with a menu entry. id is string of ASCII aphanumeric characters, underscore and hyphen and should not start with a digit. All other arguments including title are passed as positional parameters when list of commands is executed with title always assigned to $1.

17.1.2 submenu

[Command] submenu title [‘--class=class’ . . .] [‘--users=users’]
       [‘--unrestricted’] [‘--hotkey=key’] [‘--id=id’] { menu entries . . . }

This defines a submenu. An entry called title will be added to the menu; when that entry is selected, a new menu will be displayed showing all the entries within this submenu. All options are the same as in the menuentry command (see Section 17.1.1 [menuentry], page 73).

17.2 Various loader commands

These commands are used to load necessary components to boot desired OS. Many of the loader commands are not sufficiently documented. The following is a list of commands that could use more documentation:

• appleloader - Boot BIOS-based system.

• freedos - Load FreeDOS kernel.sys.

• kfreebsd_loadenv - Load FreeBSD env.

• kfreebsd_module_elf - Load FreeBSD kernel module (ELF).

• kfreebsd_module - Load FreeBSD kernel module.

• kfreebsd - Load kernel of FreeBSD.

• knetbsd_module_elf - Load NetBSD kernel module (ELF).

• knetbsd_module - Load NetBSD kernel module.

• knetbsd - Load kernel of NetBSD.

• kopenbsd - Load kernel of OpenBSD.

• kopenbsd_ramdisk - Load kOpenBSD ramdisk.

• legacy_initrd_nounzip - Simulate grub-legacy ‘modulenounzip’ command

• legacy_initrd - Simulate grub-legacy ‘initrd’ command

• legacy_kernel - Simulate grub-legacy ‘kernel’ command

• module2 - Load a multiboot 2 module.

• module - Load a multiboot module.

• multiboot2 - Load a multiboot 2 kernel.

• multiboot - Load a multiboot kernel.

• ntldr - Load NTLDR or BootMGR.

• plan9 - Load Plan9 kernel.

• pxechainloader - Load a PXE image.

• truecrypt - Load Truecrypt ISO.

• xnu_kernel64 - Load 64-bit XNU image.

• xnu_kernel - Load XNU image.

• xnu_kextdir - Load XNU extension directory.

• xnu_kext - Load XNU extension.

• xnu_mkext - Load XNU extension package.

• xnu_ramdisk - Load XNU ramdisk. It will be available in OS as md0.

• xnu_resume - Load an image of hibernated XNU.

• xnu_splash - Load a splash image for XNU.

17.2.1 chainloader

[Command] chainloader [‘--force’] file [args...]

Load file as a chain-loader. Like any other file loaded by the filesystem code, it can use the blocklist notation (see Section 13.3 [Block list syntax], page 58) to grab the first sector of the current partition with ‘+1’. On EFI platforms, any arguments after file will be sent to the loaded image. If you specify the option ‘--force’, then load file forcibly, whether it has a correct signature or not. This is required when you want to load a defective boot loader, such as SCO UnixWare 7.1.

17.2.2 initrd

[Command] initrd file [file …]

Load, in order, all initrds for a Linux kernel image, and set the appropriate parameters in the Linux setup area in memory. This may only be used after the linux command (see Section 17.2.4 [linux], page 76) has been run. See Section 5.4.2 [GNU/Linux], page 17 for more info on booting GNU/Linux. For more information on initrds see the GNU/Linux kernel documentation

(https://docs.kernel.org/filesystems/ramfs-rootfs-initramfs.html).

A new-style initrd (for kernels newer than 2.6) containing one file with leading path components can also be generated at run time. This can be done by prefixing an argument with newc: followed by the path of the file in the new initrd, a :, and then the GRUB file path to the file data to be be included. For example:

     initrd newc:/etc/ssh/config:(hd0,2)/home/user/.ssh/config \
            newc:/etc/ssh/ssh_host_rsa_key:/etc/ssh/ssh_host_rsa_key \
            /boot/initrd.gz \
            newc:/init:/home/user/init.fixed

This command will generate two new-style initrds on the fly. The first contains the

path ‘/etc/ssh/config’ with the contents of ‘(hd0,2)/home/user/.ssh/config’

and the path ‘/etc/ssh/ssh_host_rsa_key’ with the contents of ‘/etc/ssh/ssh_host_rsa_key’ on the root device. Parent directory paths will automatically be generated as needed. This first generated initrd will then have ‘/boot/initrd.gz’ concatenated after it. Next, another new-style archive will be generated with the contents of ‘/home/user/init.fixed’ in the path ‘/init’ and appended to the previous concatenation. Finally, the result will be sent to the kernel when booted. Keep in mind that paths that come later will take precedence. So in the example above, the generated path ‘/init’ will overwrite any ‘/init’ in ‘/boot/initrd.gz’. This can be useful when changing the main initrd is undesirable or difficult.

17.2.3 initrd16

[Command] initrd16 file [file …]

Load, in order, all initrds for a Linux kernel image to be booted in 16-bit mode, and set the appropriate parameters in the Linux setup area in memory. This may only be used after the linux16 command (see Section 17.2.5 [linux16], page 76) has been run. See also Section 5.4.2 [GNU/Linux], page 17 and the initrd command (see Section 17.2.2 [initrd], page 75) for more details on arguments. This command is only available on the pc platform for x86 systems.

17.2.4 linux

[Command] linux file …

Load a Linux kernel image from file. The rest of the line is passed verbatim as the kernel command-line. Any initrd must be reloaded after using this command (see Section 17.2.2 [initrd], page 75). On x86 systems, the kernel will be booted using the 32-bit boot protocol. Note that this means that the ‘vga=’ boot option will not work; if you want to set a special video mode, you will need to use GRUB commands such as ‘set gfxpayload=1024x768’ or ‘set gfxpayload=keep’ (to keep the same mode as used in GRUB) instead. GRUB can automatically detect some uses of ‘vga=’ and translate them to appropriate settings of ‘gfxpayload’. The linux16 command (see Section 17.2.5 [linux16], page 76) avoids this restriction.

17.2.5 linux16

[Command] linux16 file …

Load a Linux kernel image from file in 16-bit mode. The rest of the line is passed verbatim as the kernel command-line. Any initrd must be reloaded after using this command (see Section 17.2.3 [initrd16], page 76). The kernel will be booted using the traditional 16-bit boot protocol. As well as bypassing problems with ‘vga=’ described in Section 17.2.4 [linux], page 76, this permits booting some other programs that implement the Linux boot protocol for the sake of convenience. This command is only available on x86 systems.

17.2.6 xen hypervisor

[Command] xen_hypervisor file [arguments] …

Load a Xen hypervisor binary from file. The rest of the line is passed verbatim as the kernel command-line. Any other binaries must be reloaded after using this command. This command is only available on AArch64 systems.

17.2.7 xen module

[Command] xen_module [–nounzip] file [arguments] Load a module for xen hypervisor at the booting process of xen. The rest of the line is passed verbatim as the module command line. Modules should be loaded in the following order: - dom0 kernel image - dom0 ramdisk if present - XSM policy if present This command is only available on AArch64 systems.

17.3 General commands

Commands usable anywhere in the menu and in the command-line.

17.3.1 serial

[Command] serial [‘--unit=unit’] [‘--port=port’] [‘--speed=speed’]
   [‘--word=word’] [‘--parity=parity’] [‘--stop=stop’]

Initialize a serial device. unit is a number in the range 0-3 specifying which serial port to use; default is 0, which corresponds to the port often called COM1.

port is the I/O port where the UART is to be found or, if prefixed with ‘mmio,’, the MMIO address of the UART. If specified it takes precedence over unit.

Additionally, an MMIO address can be suffixed with:

• ‘.b’ for bytes access (default)

• ‘.w’ for 16-bit word access

• ‘.l’ for 32-bit long word access or

• ‘.q’ for 64-bit long long word access

Also, port can be of the form ‘pci,XX:XX.X’ to indicate a serial device exposed on the PCI bus.

speed is the transmission speed; default is 9600. word and stop are the number of data bits and stop bits. Data bits must be in the range 5-8 and stop bits must be 1 or 2. Default is 8 data bits and one stop bit. parity is one of ‘no’, ‘odd’, ‘even’ and defaults to ‘no’.

If passed no unit nor port, or if port is set to ‘auto’ then GRUB will attempt to use ACPI to automatically detect the system default serial port and its configuration. If this information is not available, it will default to unit 0.

The serial port is not used as a communication channel unless the terminal_input or terminal_output command is used (see Section 17.3.2 [terminal input], page 78, see Section 17.3.3 [terminal output], page 78).

Note, valid port values, excluding IO port addresses, can be found by listing terminals with terminal_output, selecting all names prefixed by ‘serial_’ and removing that prefix.

Examples:

     serial --port=0x3f8 --speed=9600
     serial --port=mmio,fefb0000.l --speed=115200
     serial --port=pci,00:16.3 --speed=115200

See also Chapter 9 [Serial terminal], page 49.

17.3.2 terminal input

[Command] terminal_input [‘--append’|‘--remove’] [terminal1] [terminal2] …

List or select an input terminal.

With no arguments, list the active and available input terminals.

With ‘--append’, add the named terminals to the list of active input terminals; any of these may be used to provide input to GRUB.

With ‘--remove’, remove the named terminals from the active list.

With no options but a list of terminal names, make only the listed terminal names active.

17.3.3 terminal output

[Command] terminal_output [‘--append’|‘--remove’] [terminal1] [terminal2] …

List or select an output terminal.

With no arguments, list the active and available output terminals.

With ‘--append’, add the named terminals to the list of active output terminals; all of these will receive output from GRUB.

With ‘--remove’, remove the named terminals from the active list.

With no options but a list of terminal names, make only the listed terminal names active.

17.3.4 terminfo

[Command] terminfo [‘-a’|‘-u’|‘-v’] [‘-g WxH’] [term] [type]

Define the capabilities of your terminal by giving the name of an entry in the terminfo database, which should correspond roughly to a ‘TERM’ environment variable in Unix.

The currently available terminal types are ‘vt100’, ‘vt100-color’, ‘ieee1275’, and ‘dumb’. If you need other terminal types, please contact us to discuss the best way to include support for these in GRUB.

The ‘-a’ (‘--ascii’), ‘-u’ (‘--utf8’), and ‘-v’ (‘--visual-utf8’) options control how non-ASCII text is displayed. ‘-a’ specifies an ASCII-only terminal; ‘-u’ specifies logically-ordered UTF-8; and ‘-v’ specifies "visually-ordered UTF-8" (in other words, arranged such that a terminal emulator without bidirectional text support will display right-to-left text in the proper order; this is not really proper UTF-8, but a workaround).

The ‘-g’ (‘--geometry’) can be used to specify terminal geometry.

If no option or terminal type is specified, the current terminal type is printed.

17.4 Command-line commands

These commands are usable in the command-line and in menu entries. If you forget a command, you can run the command help (see Section 17.4.35 [help], page 86).

17.4.1 [

[Command] [ expression ] Alias for test expression (see Section 17.4.74 [test], page 98).

17.4.2 acpi

[Command] acpi [‘-1’|‘-2’]

   [‘--exclude=table1,...’|‘--load-only=table1,...’] [‘--oemid=id’]
   [‘--oemtable=table’] [‘--oemtablerev=rev’]
   [‘--oemtablecreator=creator’] [‘--oemtablecreatorrev=rev’]

[‘--no-ebda’] filename …

Modern BIOS systems normally implement the Advanced Configuration and Power Interface (ACPI), and define various tables that describe the interface between an ACPI-compliant operating system and the firmware. In some cases, the tables provided by default only work well with certain operating systems, and it may be necessary to replace some of them. Normally, this command will replace the Root System Description Pointer (RSDP) in the Extended BIOS Data Area to point to the new tables. If the ‘--no-ebda’ option is used, the new tables will be known only to GRUB, but may be used by GRUB’s EFI emulation. Note: The command is not allowed when lockdown is enforced (see Section 19.6 [Lockdown], page 114). Otherwise an attacker can instruct the GRUB to load an SSDT table to overwrite the kernel lockdown configuration and later load and execute unsigned code.

17.4.3 authenticate

[Command] authenticate [userlist]

Check whether user is in userlist or listed in the value of variable ‘superusers’. See see Section 15.1.45 [superusers], page 69 for valid user list format. If ‘superusers’ is empty, this command returns true. See Chapter 19 [Security], page 111.

17.4.4 background color

[Command] background_color color

Set background color for active terminal. For valid color specifications see see Chapter 7 [Colors], page 35. Background color can be changed only when using ‘gfxterm’ for terminal output. This command sets color of empty areas without text. Text background color is controlled by environment variables color normal, color highlight, menu color normal, menu color highlight. See Section 15.1 [Special environment variables], page 63.

17.4.5 background image

[Command] background_image [[‘--mode’ ‘stretch’|‘normal’] file]

Load background image for active terminal from file. Image is stretched to fill up entire screen unless option ‘--mode’ ‘normal’ is given. Without arguments remove currently loaded background image. Background image can be changed only when using ‘gfxterm’ for terminal output.

17.4.6 badram

[Command] badram addr,mask[,addr,mask...] Filter out bad RAM. This command notifies the memory manager that specified regions of RAM ought to be filtered out (usually, because they’re damaged). This remains in effect after a payload kernel has been loaded by GRUB, as long as the loaded kernel obtains its memory map from GRUB. Kernels that support this include Linux, GNU Mach, the kernel of FreeBSD and Multiboot kernels in general. Syntax is the same as provided by the Memtest86+ utility (http://www.memtest.org/): a list of address/mask pairs. Given a page-aligned address and a base address / mask pair, if all the bits of the page-aligned address that are enabled by the mask match with the base address, it means this page is to be filtered. This syntax makes it easy to represent patterns that are often result of memory damage, due to physical distribution of memory cells. The command is similar to cutmem command. Note: The command is not allowed when lockdown is enforced (see Section 19.6 [Lockdown], page 114). This prevents removing EFI memory regions to potentially subvert the security mechanisms provided by the UEFI secure boot.

17.4.7 blocklist

[Command] blocklist file

Print a block list (see Section 13.3 [Block list syntax], page 58) for file.

17.4.8 boot

[Command] boot

Boot the OS or chain-loader which has been loaded. Only necessary if running the fully interactive command-line (it is implicit at the end of a menu entry).

17.4.9 cat

[Command] cat [‘--dos’] file

Display the contents of the file file. This command may be useful to remind you of your OS’s root partition:

     grub> cat /etc/fstab

If the ‘--dos’ option is used, then carriage return / new line pairs will be displayed as a simple new line. Otherwise, the carriage return will be displayed as a control character (‘<d>’) to make it easier to see when boot problems are caused by a file formatted using DOS-style line endings. Note: cat can be used to view the contents of devices using the block list syntax (see Section 13.3 [Block list syntax], page 58). However, it is not advised to view binary data because it will try to decode UTF-8 strings, which can lead to some bytes missing or added in the output. Instead, use the hexdump command (see Section 17.4.36 [hexdump], page 87).

17.4.10 clear

[Command] clear

Clear the screen.

17.4.11 cmosclean

[Command] cmosclean byte:bit

Clear value of bit in CMOS at location byte:bit. This command is available only on platforms that support CMOS.

17.4.12 cmosdump

[Dump] CMOS contents Dump full CMOS contents as hexadecimal values. This command is available only on platforms that support CMOS.

17.4.13 cmostest

[Command] cmostest byte:bit

Test value of bit in CMOS at location byte:bit. Exit status is zero if bit is set, non zero otherwise. This command is available only on platforms that support CMOS.

17.4.14 cmp

[Command] cmp [‘-v’] file1 file2 Compare the file file1 with the file file2. If they are completely identical, $? will be set to 0. Otherwise, if the files are not identical, $? will be set to a nonzero value. By default nothing will be output. If the ‘-v’ is used, verbose mode is enabled. In this mode when when the files differ in size, print the sizes like this:

     Differ in size: 0x1234 [foo], 0x4321 [bar]

If the sizes are equal but the bytes at an offset differ, then print the bytes like this:

     Differ at the offset 777: 0xbe [foo], 0xef [bar]

17.4.15 configfile

[Command] configfile file

Load file as a configuration file. If file defines any menu entries, then show a menu containing them immediately. Any environment variable changes made by the commands in file will not be preserved after configfile returns.

17.4.16 cpuid

[Command] cpuid [-l] [-p]

Check for CPU features. This command is only available on x86 systems. With the ‘-l’ option, return true if the CPU supports long mode (64-bit).

With the ‘-p’ option, return true if the CPU supports Physical Address Extension (PAE). If invoked without options, this command currently behaves as if it had been invoked with ‘-l’. This may change in the future.

17.4.17 crc

[Command] crc arg …

Alias for hashsum --hash crc32 arg .... See command hashsum (see Section 17.4.34 [hashsum], page 86) for full description.

17.4.18 cryptomount

[Command] cryptomount [ [‘-p’ password] | [‘-k’ keyfile [‘-O’ keyoffset] [‘-S’

keysize] ] ] [‘-H’ file] device|‘-u’ uuid|‘-a’|‘-b’ Setup access to encrypted device. A passphrase will be requested interactively, if neither the ‘-p’ nor ‘-k’ options are given. The option ‘-p’ can be used to supply a passphrase (useful for scripts). Alternatively the ‘-k’ option can be used to supply a keyfile with options ‘-O’ and ‘-S’ optionally supplying the offset and size, respectively, of the key data in the given key file. The ‘-H’ options can be used to supply cryptomount backends with an alternative header file (aka detached header). Not all backends have headers nor support alternative header files (currently only LUKS1 and LUKS2 support them). Argument device configures specific grub device (see Chapter 2 [Naming convention], page 7); option ‘-u’ uuid configures device with specified uuid; option ‘-a’ configures all detected encrypted devices; option ‘-b’ configures all geli containers that have boot flag set. Devices are not allowed to be given as key files nor as detached header files. However, this limitation can be worked around by using blocklist syntax. So for instance, (hd1,gpt2) can not be used, but (hd1,gpt2)0+ will achieve the desired result. GRUB suports devices encrypted using LUKS, LUKS2 and geli. Note that necessary modules (luks, luks2 and geli) have to be loaded manually before this command can be used. For LUKS2 only the PBKDF2 key derivation function is supported, as Argon2 is not yet supported. Successfully decrypted disks are named as (cryptoX) and have increasing numeration suffix for each new decrypted disk. If the encrypted disk hosts some higher level of abstraction (like LVM2 or MDRAID) it will be created under a separate device namespace in addition to the cryptodisk namespace. Support for plain encryption mode (plain dm-crypt) is provided via separate see

Section 17.4.56 [plainmount], page 90 command.

17.4.19 cutmem

[Command] cutmem from[K|M|G] to[K|M|G]

Remove any memory regions in specified range. This command notifies the memory manager that specified regions of RAM ought to be filtered out. This remains in effect after a payload kernel has been loaded by GRUB, as long as the loaded kernel obtains its memory map from GRUB. Kernels that support this include Linux, GNU Mach, the kernel of FreeBSD and Multiboot kernels in general. The command is similar to badram command. Note: The command is not allowed when lockdown is enforced (see Section 19.6 [Lockdown], page 114). This prevents removing EFI memory regions to potentially subvert the security mechanisms provided by the UEFI secure boot.

17.4.20 date

[Command] date [[year-]month-day] [hour:minute[:second]] With no arguments, print the current date and time. Otherwise, take the current date and time, change any elements specified as arguments, and set the result as the new date and time. For example, ‘date 01-01’ will set the current month and day to January 1, but leave the year, hour, minute, and second unchanged.

17.4.21 devicetree

[Command] devicetree file

Load a device tree blob (.dtb) from a filesystem, for later use by a Linux kernel. Does not perform merging with any device tree supplied by firmware, but rather replaces it completely. Note: The command is not allowed when lockdown is enforced (see Section 19.6 [Lockdown], page 114). This is done to prevent subverting various security mechanisms.

17.4.22 distrust

[Command] distrust pubkey id

Remove public key pubkey id from GRUB’s keyring of trusted keys. pubkey id is the last four bytes (eight hexadecimal digits) of the GPG v4 key id, which is also the output of list_trusted (see Section 17.4.40 [list trusted], page 87). Outside of GRUB, the key id can be obtained using gpg --fingerprint). These keys are used to validate signatures when environment variable check_signatures is set to enforce (see Section 15.1.2 [check signatures], page 63), and by some invocations of verify_detached (see Section 17.4.78 [verify detached], page 100). See Section 19.2 [Using digital signatures], page 112, for more information.

17.4.23 drivemap

[Command] drivemap ‘-l’|‘-r’|[‘-s’] from drive to drive Without options, map the drive from drive to the drive to drive. This is necessary when you chain-load some operating systems, such as DOS, if such an OS resides at a non-first drive. For convenience, any partition suffix on the drive is ignored, so you can safely use ${root} as a drive specification. With the ‘-s’ option, perform the reverse mapping as well, swapping the two drives. With the ‘-l’ option, list the current mappings. With the ‘-r’ option, reset all mappings to the default values. For example:

     drivemap -s (hd0) (hd1)

NOTE: Only available on i386-pc.

17.4.24 echo

[Command] echo [‘-n’] [‘-e’] string …

Display the requested text and, unless the ‘-n’ option is used, a trailing new line. If there is more than one string, they are separated by spaces in the output. As usual in GRUB commands, variables may be substituted using ‘${var}’. The ‘-e’ option enables interpretation of backslash escapes. The following sequences are recognised:

\\ backslash

\a alert (BEL)

\c suppress trailing new line

\f form feed

\n new line

\r carriage return

\t horizontal tab

\v vertical tab

When interpreting backslash escapes, backslash followed by any other character will print that character.

17.4.25 efitextmode

[Command] efitextmode [min | max | <mode num> | <cols> <rows>]

When used with no arguments displays all available text output modes. The set mode determines the columns and rows of the text display when in text mode. An asterisk, ‘*’, will be at the end of the line of the currently set mode. If given a single parameter, it must be ‘min’, ‘max’, or a mode number given by the listing when run with no arguments. These arguments set the mode to the minimum, maximum, and particular mode respectively. Otherwise, the command must be given two numerical arguments specifying the columns and rows of the desired mode. Specifying a columns and rows combination that corresponds to no supported mode, will return error, but otherwise have no effect. By default GRUB will start in whatever mode the EFI firmware defaults to. There are firmwares known to set up the default mode such that output behaves strangely, for example the cursor in the GRUB shell never reaches the bottom of the screen or, when typing characters at the prompt, characters from previous command output are overwritten. Setting the mode may fix this. The EFI specification says that mode 0 must be available and have columns and rows of 80 and 25 respectively. Mode 1 may be defined and if so must have columns and rows of 80 and 50 respectively. Any other modes may have columns and rows arbitrarily defined by the firmware. This means that a mode with columns and rows of 100 and 31 on one firmware may be a different mode number on a different firmware or not exist at all. Likewise, mode number 2 on one firmware may have a different number of columns and rows than mode 2 on a different firmware. So one should not rely on a particular mode number or a mode of a certain number of columns and rows existing on all firmwares, except for mode 0. Note: This command is only available on EFI platforms and is similar to EFI shell "mode" command.

17.4.26 eval

[Command] eval string ...

Concatenate arguments together using single space as separator and evaluate result as sequence of GRUB commands.

17.4.27 export

[Command] export envvar

Export the environment variable envvar. Exported variables are visible to subsidiary configuration files loaded using configfile.

17.4.28 false

[Command] false

Do nothing, unsuccessfully. This is mainly useful in control constructs such as if and while (see Section 6.3 [Shell-like scripting], page 28).

17.4.29 fwsetup

[Command] fwsetup [‘--is-supported’]

Reboot into the firmware setup menu. If ‘--is-supported’ option is specified, instead check whether the firmware supports a setup menu and exit successfully if so.

17.4.30 gdbinfo

[Command] gdbinfo

Output text to be used as a GDB command for a GDB session using the gdb grub script and attached to a running GRUB instance. The GDB command that is output will tell GDB how to load debugging symbols to their proper runtime address. Currently this is only available for EFI platforms. See the Debugging in the developer documentation for more information.

17.4.31 gettext

[Command] gettext string

Translate string into the current language. The current language code is stored in the ‘lang’ variable in GRUB’s environment (see Section 15.1.18 [lang], page 67). Translation files in MO format are read from ‘locale_dir’ (see Section 15.1.19 [locale dir], page 67), usually

‘/boot/grub/locale’.

17.4.32 gptsync

[Command] gptsync device [partition[+/-[type]]] … Disks using the GUID Partition Table (GPT) also have a legacy Master Boot Record (MBR) partition table for compatibility with the BIOS and with older operating systems. The legacy MBR can only represent a limited subset of GPT partition entries. This command populates the legacy MBR with the specified partition entries on device. Up to three partitions may be used. type is an MBR partition type code; prefix with ‘0x’ if you want to enter this in hexadecimal. The separator between partition and type may be ‘+’ to make the partition active, or ‘-’ to make it inactive; only one partition may be active. If both the separator and type are omitted, then the partition will be inactive.

17.4.33 halt

[Command] halt [‘--no-apm’]

The command halts the computer. On the i386-pc target, the ‘--no-apm’ option, or short ‘-n’, is specified, no APM BIOS call is performed. Otherwise, the computer is shut down using APM on that target.

17.4.34 hashsum

[Command] hashsum ‘--hash’ hash ‘--keep-going’ ‘--uncompress’ ‘--check’ file

[‘--prefix’ dir]|file …

Compute or verify file hashes. Hash type is selected with option ‘--hash’. Sup-

ported hashes are: ‘adler32’, ‘crc64’, ‘crc32’, ‘crc32rfc1510’, ‘crc24rfc2440’,
‘md4’, ‘md5’, ‘ripemd160’, ‘sha1’, ‘sha224’, ‘sha256’, ‘sha512’, ‘sha384’, ‘tiger192’,

‘tiger’, ‘tiger2’, ‘whirlpool’. Option ‘--uncompress’ uncompresses files before computing hash. When list of files is given, hash of each file is computed and printed, followed by file name, each file on a new line. When option ‘--check’ is given, it points to a file that contains list of hash name pairs in the same format as used by UNIX md5sum command. Option ‘--prefix’ may be used to give directory where files are located. Hash verification stops after the first mismatch was found unless option ‘--keep-going’ was given. The exit code $? is set to 0 if hash verification is successful. If it fails, $? is set to a nonzero value.

17.4.35 help

[Command] help [pattern …] Display helpful information about builtin commands. If you do not specify pattern, this command shows short descriptions of all available commands. If you specify any patterns, it displays longer information about each of the commands whose names begin with those patterns.

17.4.36 hexdump

[Command] hexdump [–skip offset] [–length len] FILE OR DEVICE Show raw contents of a file or memory. When option ‘--skip’ is given, ‘offset’ number of bytes are skipped from the start of the device or file given. And ‘--length’ allows specifying a maximum number of bytes to be shown. If given the special device named ‘(mem)’, then the ‘offset’ given to ‘--skip’ is treated as the address of a memory location to dump from.

17.4.37 insmod

[Command] insmod module

Insert the dynamic GRUB module called module.

17.4.38 keystatus

[Command] keystatus [‘--shift’] [‘--ctrl’] [‘--alt’]

Return true if the Shift, Control, or Alt modifier keys are held down, as requested by options. This is useful in scripting, to allow some user control over behaviour without having to wait for a keypress. Checking key modifier status is only supported on some platforms. If invoked without any options, the keystatus command returns true if and only if checking key modifier status is supported.

17.4.39 list env

[Command] list_env [‘--file’ file]

List all variables in the environment block file. See Section 15.2 [Environment block], page 70. The ‘--file’ option overrides the default location of the environment block.

17.4.40 list trusted

[Command] list_trusted

List all public keys trusted by GRUB for validating signatures. The output is in GPG’s v4 key fingerprint format (i.e., the output of gpg --fingerprint). The least significant four bytes (last eight hexadecimal digits) can be used as an argument to distrust (see Section 17.4.22 [distrust], page 83). See Section 19.2 [Using digital signatures], page 112, for more information about uses for these keys.

17.4.41 load env

[Command] load_env [‘--file’ file] [‘--skip-sig’] [whitelisted variable name] …

Load all variables from the environment block file into the environment. See Section 15.2 [Environment block], page 70. The ‘--file’ option overrides the default location of the environment block. The ‘--skip-sig’ option skips signature checking even when the value of environment variable check_signatures is set to enforce (see Section 15.1.2 [check signatures], page 63).

If one or more variable names are provided as arguments, they are interpreted as a whitelist of variables to load from the environment block file. Variables set in the file but not present in the whitelist are ignored. The ‘--skip-sig’ option should be used with care, and should always be used in concert with a whitelist of acceptable variables whose values should be set. Failure to employ a carefully constructed whitelist could result in reading a malicious value into critical environment variables from the file, such as setting check_signatures=no, modifying prefix to boot from an unexpected location or not at all, etc. When used with care, ‘--skip-sig’ and the whitelist enable an administrator to configure a system to boot only signed configurations, but to allow the user to select from among multiple configurations, and to enable “one-shot” boot attempts and “savedefault” behavior. See Section 19.2 [Using digital signatures], page 112, for more information.

17.4.42 loadfont

[Command] loadfont file …

Load specified font files. Unless absolute pathname is given, file is assumed to be in directory ‘$prefix/fonts’ with suffix ‘.pf2’ appended. See Chapter 7 [Fonts], page 35.

17.4.43 loopback

[Command] loopback [‘-d’] [‘-D’] device file

Make the device named device correspond to the contents of the filesystem image in file. For example:

     loopback loop0 /path/to/image
     ls (loop0)/

Specifying the ‘-D’ option allows the loopback file to be tranparently decompressed if there is an appropriate decompressor loaded. With the ‘-d’ option, delete a device previously created using this command.

17.4.44 ls

[Command] ls [arg …]

List devices or files. With no arguments, print all devices known to GRUB. If the argument is a device name enclosed in parentheses (see Section 13.1 [Device syntax], page 57), then print the name of the filesystem of that device. If the argument is a directory given as an absolute file name (see Section 13.2 [File name syntax], page 58), then list the contents of that directory.

17.4.45 lsfonts

[Command] lsfonts

List loaded fonts.

17.4.46 lsmod

[Command] lsmod

Show list of loaded modules.

17.4.47 md5sum

[Command] md5sum arg …

Alias for hashsum --hash md5 arg .... See command hashsum (see Section 17.4.34 [hashsum], page 86) for full description.

17.4.48 module

[Command] module [–nounzip] file [arguments] Load a module for multiboot kernel image. The rest of the line is passed verbatim as the module command line.

17.4.49 multiboot

[Command] multiboot [–quirk-bad-kludge] [–quirk-modules-after-kernel] file … Load a multiboot kernel image from file. The rest of the line is passed verbatim as the kernel command-line. Any module must be reloaded after using this command (see Section 17.4.48 [module], page 89). Some kernels have known problems. You need to specify –quirk-* for those. –quirkbad-kludge is a problem seen in several products that they include loading kludge information with invalid data in ELF file. GRUB prior to 0.97 and some custom builds preferred ELF information while 0.97 and GRUB 2 use kludge. Use this option to ignore kludge. Known affected systems: old Solaris, SkyOS. –quirk-modules-after-kernel is needed for kernels which load at relatively high address e.g. 16MiB mark and can’t cope with modules stuffed between 1MiB mark and beginning of the kernel. Known afftected systems: VMWare.

17.4.50 nativedisk

[Command] nativedisk

Switch from firmware disk drivers to native ones. Really useful only on platforms where both firmware and native disk drives are available. Currently i386-pc, i386-efi, i386-ieee1275 and x86 64-efi.

17.4.51 normal

[Command] normal [file]

Enter normal mode and display the GRUB menu. In normal mode, commands, filesystem modules, and cryptography modules are automatically loaded, and the full GRUB script parser is available. Other modules may be explicitly loaded using insmod (see Section 17.4.37 [insmod], page 87). If a file is given, then commands will be read from that file. Otherwise, they will be read from ‘$prefix/grub.cfg’ if it exists. normal may be called from within normal mode, creating a nested environment. It is more usual to use configfile (see Section 17.4.15 [configfile], page 81) for this.

17.4.52 normal exit

[Command] normal_exit

Exit normal mode (see Section 17.4.51 [normal], page 89). If this instance of normal mode was not nested within another one, then return to rescue mode.

17.4.53 parttool

[Command] parttool partition commands Make various modifications to partition table entries. Each command is either a boolean option, in which case it must be followed with ‘+’ or ‘-’ (with no intervening space) to enable or disable that option, or else it takes a value in the form ‘command=value’. Currently, parttool is only useful on DOS partition tables (also known as Master Boot Record, or MBR). On these partition tables, the following commands are available:

‘boot’ (boolean)

When enabled, this makes the selected partition be the active (bootable) partition on its disk, clearing the active flag on all other partitions. This command is limited to primary partitions.

‘type’ (value)

Change the type of an existing partition. The value must be a number in the range 0-0xFF (prefix with ‘0x’ to enter it in hexadecimal).

‘hidden’ (boolean)

When enabled, this hides the selected partition by setting the hidden bit in its partition type code; when disabled, unhides the selected partition by clearing this bit. This is useful only when booting DOS or Windows and multiple primary FAT partitions exist in one disk. See also Section 5.4.4 [DOS/Windows], page 18.

17.4.54 password

[Command] password user clear-password Define a user named user with password clear-password. See Chapter 19 [Security], page 111.

17.4.55 password pbkdf2

[Command] password_pbkdf2 user hashed-password Define a user named user with password hash hashed-password. Use grub-mkpasswd-pbkdf2 (see Section 24.3 [Invoking grub-mkpasswd-pbkdf2], page 130) to generate password hashes. See Chapter 19 [Security], page 111.

17.4.56 plainmount

[Command] plainmount device ‘-c’ cipher ‘-s’ key size [‘-h’ hash] [‘-S’ sector size] [‘-p’ password] [‘-u’ uuid] [[‘-d’ keyfile] [‘-O’ keyfile offset]] Setup access to the encrypted device in plain mode. Offset of the encrypted data at the device is specified in terms of 512 byte sectors using the blocklist syntax and loopback device. The following example shows how to specify 1MiB offset:

     loopback node (hd0,gpt1)2048+
     plainmount node ...

The plainmount command can be used to open LUKS encrypted volume if its master key and parameters (key size, cipher, offset, etc) are known.

There are two ways to specify a password: a keyfile and a secret passphrase. The keyfile path parameter has higher priority than the secret passphrase parameter and is specified with the option ‘-d’. Password data obtained from keyfiles is not hashed and is used directly as a cipher key. An optional offset of password data in the keyfile can be specified with the option ‘-O’ or directly with the option ‘-d’ and GRUB blocklist syntax, if the keyfile data can be accessed from a device and is 512 byte aligned. The following example shows both methods to specify password data in the keyfile at offset 1MiB:

     plainmount -d (hd0,gpt1)2048+ ...
     plainmount -d (hd0,gpt1)+ -O 1048576 ...

If no keyfile is specified then the password is set to the string specified by option ‘-p’ or is requested interactively from the console. In both cases the provided password is hashed with the algorithm specified by the option ‘-h’. This option is mandatory if no keyfile is specified, but it can be set to ‘plain’ which means that no hashing is done and such password is used directly as a key.

Cipher ‘-c’ and keysize ‘-s’ options specify the cipher algorithm and the key size respectively and are mandatory options. Cipher must be specified with the mode separated by a dash (for example, ‘aes-xts-plain64’). Key size option ‘-s’ is the key size of the cipher in bits, not to be confused with the offset of the key data in a keyfile specified with the ‘-O’ option. It must not exceed 1024 bits, so a 32 byte key would be specified as 256 bits

The optional parameter ‘-S’ specifies encrypted device sector size. It must be at least 512 bytes long (default value) and a power of 2.^{1}. Disk sector size is configured when creating the encrypted volume. Attempting to decrypt volumes with a different sector size than it was created with will not result in an error, but will decrypt to random bytes and thus prevent accessing the volume (in some cases the filesystem driver can detect the presence of a filesystem, but nevertheless will refuse to mount it).

By default new plainmount devices will be given a UUID starting with ’109fea84a6b7-34a8-4bd1-1c506305a401’ where the last digits are incremented by one for each plainmounted device beyond the first up to 2^10 devices.

All encryption arguments (cipher, hash, key size, disk offset and disk sector size) must match the parameters used to create the volume. If any of them does not match the actual arguments used during the initial encryption, plainmount will create virtual device with the garbage data and GRUB will report unknown filesystem for such device.

1 Current implementation of cryptsetup supports only 512/1024/2048/4096 byte sectors

17.4.57 play

[Command] play file | tempo [pitch1 duration1] [pitch2 duration2] … Plays a tune If the argument is a file name (see Section 13.2 [File name syntax], page 58), play the tune recorded in it. The file format is first the tempo as an unsigned 32bit little-endian number, then pairs of unsigned 16bit little-endian numbers for pitch and duration pairs. If the arguments are a series of numbers, play the inline tune. The tempo is the base for all note durations. 60 gives a 1-second base, 120 gives a half-second base, etc. Pitches are Hz. Set pitch to 0 to produce a rest.

17.4.58 probe

[Command] probe [‘--set’ var]

   ‘--driver’|‘--partmap’|‘--fs’|‘--fs-uuid’|‘--label’|‘--part-uuid’

device Retrieve device information. If option ‘--set’ is given, assign result to variable var, otherwise print information on the screen. The option ‘--part-uuid’ is currently only implemented for MSDOS and GPT formatted disks.

17.4.59 rdmsr

[Command:] rdmsr 0xADDR [-v VARNAME] Read a model-specific register at address 0xADDR. If the parameter ‘-v’ is used and an environment variable VARNAME is given, set that environment variable to the value that was read. Please note that on SMP systems, reading from a MSR that has a scope per hardware thread, implies that the value that is returned only applies to the particular cpu/core/thread that runs the command. Also, if you specify a reserved or unimplemented MSR address, it will cause a general protection exception (which is not currently being handled) and the system will reboot.

17.4.60 read

[Command] read [-s] [var]

Read a line of input from the user. If an environment variable var is given, set that environment variable to the line of input that was read, with no terminating newline. If the parameter ‘-s’ is used, enable silent mode where input is not printed to the terminal.

17.4.61 reboot

[Command] reboot

Reboot the computer.

17.4.62 regexp

[Command] regexp [‘--set’ [number:]var] regexp string Test if regular expression regexp matches string. Supported regular expressions are POSIX.2 Extended Regular Expressions. If option ‘--set’ is given, store numberth matched subexpression in variable var. Subexpressions are numbered in order of their opening parentheses starting from ‘1’. number defaults to ‘1’.

17.4.63 rmmod

[Command] rmmod module Remove a loaded module.

17.4.64 save env

[Command] save_env [‘--file’ file] var …

Save the named variables from the environment to the environment block file. See Section 15.2 [Environment block], page 70. The ‘--file’ option overrides the default location of the environment block. This command will operate successfully even when environment variable check_ signatures is set to enforce (see Section 15.1.2 [check signatures], page 63), since it writes to disk and does not alter the behavior of GRUB based on any contents of disk that have been read. It is possible to modify a digitally signed environment block file from within GRUB using this command, such that its signature will no longer be valid on subsequent boots. Care should be taken in such advanced configurations to avoid rendering the system unbootable. See Section 19.2 [Using digital signatures], page 112, for more information.

17.4.65 search

[Command] search [‘--file’|‘--label’|‘--fs-uuid’] [‘--set’ [var]]
   [‘--no-floppy’|‘--efidisk-only’] name

Search devices by file (‘-f’, ‘--file’), filesystem label (‘-l’, ‘--label’), or filesystem UUID (‘-u’, ‘--fs-uuid’). If the (‘-s’, ‘--set’) option is used, the first device found is set as the value of environment variable var. The default variable is ‘root’. The (‘-n’, ‘--no-floppy’) option prevents searching floppy devices, which can be slow. The (‘--efidisk-only’) option prevents searching any other devices then EFI disks. This is typically used when chainloading to local EFI partition.

The ‘search.file’, ‘search.fs_label’, and ‘search.fs_uuid’ commands are aliases
for ‘search --file’, ‘search --label’, and ‘search --fs-uuid’ respectively.

Also hints as to which device may be the most likely to contain the item searched for may be given via the (‘-h’, ‘--hint’) option with a device name as an argument. If the argument ends with a comma, then partitions on the device are also searched. Furthermore, platform specific hints may be given via the options ‘--hint-ieee1275’,

‘--hint-bios’, ‘--hint-baremetal’, ‘--hint-efi’, and ‘--hint-arc’. When spec-

ified, these options take an argument and operate like ‘--hint’, but only on the specified platform.

17.4.66 sendkey

[Command] sendkey [‘--num’|‘--caps’|‘--scroll’|‘--insert’|
   ‘--pause’|‘--left-shift’|‘--right-shift’|
   ‘--sysrq’|‘--numkey’|‘--capskey’|‘--scrollkey’|
   ‘--insertkey’|‘--left-alt’|‘--right-alt’|
   ‘--left-ctrl’|‘--right-ctrl’ ‘on’|‘off’]. . . [‘no-led’] keystroke

Insert keystrokes into the keyboard buffer when booting. Sometimes an operating system or chainloaded boot loader requires particular keys to be pressed: for example, one might need to press a particular key to enter "safe mode", or when chainloading another boot loader one might send keystrokes to it to navigate its menu. Note: This command is currently only available on the i386-pc target. You may provide up to 16 keystrokes (the length of the BIOS keyboard buffer). Keystroke names may be upper-case or lower-case letters, digits, or taken from the following table:

Name  Key
escape  Escape
exclam  !
at  @
numbersign  #
dollar  $
percent  %
caret  ^
ampersand  &
asterisk  *
parenleft  (
parenright  )
minus  -

underscore

equal  =
plus  +
backspace  Backspace
tab  Tab
bracketleft  [
braceleft  {
bracketright  ]
braceright  }
enter  Enter
control  press and release Control
semicolon  ;
colon  :
quote  ’

doublequote  "
backquote  ‘
tilde  ~
shift  press and release left Shift
backslash  \
bar  |
comma  ,
less  <
period  .
greater  >
slash  /
question  ?
rshift  press and release right Shift
alt  press and release Alt
space  space bar
capslock  Caps Lock
F1  F1
F2  F2
F3  F3
F4  F4
F5  F5
F6  F6
F7  F7
F8  F8
F9  F9
F10  F10
F11  F11
F12  F12
num1  1 (numeric keypad)
num2  2 (numeric keypad)
num3  3 (numeric keypad)
num4  4 (numeric keypad)
num5  5 (numeric keypad)
num6  6 (numeric keypad)
num7  7 (numeric keypad)
num8  8 (numeric keypad)
num9  9 (numeric keypad)
num0  0 (numeric keypad)
numperiod  . (numeric keypad)
numend  End (numeric keypad)
numdown  Down (numeric keypad)
numpgdown  Page Down (numeric keypad)
numleft  Left (numeric keypad)

numcenter 5 with Num Lock inactive (numeric keypad)

numright  Right (numeric keypad)

numhome  Home (numeric keypad)
numup  Up (numeric keypad)
numpgup  Page Up (numeric keypad)
numinsert  Insert (numeric keypad)
numdelete  Delete (numeric keypad)
numasterisk  * (numeric keypad)
numminus  - (numeric keypad)
numplus  + (numeric keypad)
numslash  / (numeric keypad)
numenter  Enter (numeric keypad)
delete  Delete
insert  Insert
home  Home
end  End
pgdown  Page Down
pgup  Page Up
down  Down
up  Up
left  Left
right  Right

As well as keystrokes, the sendkey command takes various options that affect the BIOS keyboard status flags. These options take an ‘on’ or ‘off’ parameter, specifying that the corresponding status flag be set or unset; omitting the option for a given status flag will leave that flag at its initial state at boot. The ‘--num’, ‘--caps’, ‘--scroll’, and ‘--insert’ options emulate setting the corresponding mode, while

the ‘--numkey’, ‘--capskey’, ‘--scrollkey’, and ‘--insertkey’ options emulate

pressing and holding the corresponding key. The other status flag options are selfexplanatory.

If the ‘--no-led’ option is given, the status flag options will have no effect on keyboard LEDs.

If the sendkey command is given multiple times, then only the last invocation has any effect.

Since sendkey manipulates the BIOS keyboard buffer, it may cause hangs, reboots, or other misbehaviour on some systems. If the operating system or boot loader that runs after GRUB uses its own keyboard driver rather than the BIOS keyboard functions, then sendkey will have no effect.

This command is only available on PC BIOS systems.

17.4.67 set

[Command] set [envvar=value] Set the environment variable envvar to value. If invoked with no arguments, print all environment variables with their values. For the list of environment variables currently used by GRUB itself see the relevant section see Chapter 15 [Environment], page 63.

17.4.68 sha1sum

[Command] sha1sum arg …

Alias for hashsum --hash sha1 arg .... See command hashsum (see Section 17.4.34 [hashsum], page 86) for full description.

17.4.69 sha256sum

[Command] sha256sum arg …

Alias for hashsum --hash sha256 arg ....  See command hashsum (see

Section 17.4.34 [hashsum], page 86) for full description.

17.4.70 sha512sum

[Command] sha512sum arg …

Alias for hashsum --hash sha512 arg ....  See command hashsum (see

Section 17.4.34 [hashsum], page 86) for full description.

17.4.71 sleep

[Command] sleep [‘--verbose’] [‘--interruptible’] count

Sleep for count seconds. If option ‘--interruptible’ is given, allow pressing ESC, F4 or holding down SHIFT to interrupt sleep. With ‘--verbose’ show countdown of remaining seconds. Exit code is set to 0 if timeout expired and to 1 if timeout was interrupted using any of the mentioned keys.

17.4.72 smbios

[Command] smbios [‘--type’ type] [‘--handle’ handle] [‘--match’ match]
   (‘--get-byte’ | ‘--get-word’ | ‘--get-dword’ | ‘--get-qword’ |
   ‘--get-string’ | ‘--get-uuid’) offset [‘--set’ variable]

Retrieve SMBIOS information. The smbios command returns the value of a field in an SMBIOS structure. The following options determine which structure to select.

• Specifying ‘--type’ will select structures with a matching type. The type can be any integer from 0 to 255.

• Specifying ‘--handle’ will select structures with a matching handle. The handle can be any integer from 0 to 65535.

• Specifying ‘--match’ will select structure number match in the filtered list of structures; e.g. smbios --type 4 --match 2 will select the second Process Information (Type 4) structure. The list is always ordered the same as the hardware’s SMBIOS table. The match number must be a positive integer. If unspecified, the first matching structure will be selected.

The remaining options determine which field in the selected SMBIOS structure to return. Only one of these options may be specified at a time.

• When given ‘--get-byte’, return the value of the byte at offset bytes into the selected SMBIOS structure. It will be formatted as an unsigned decimal integer.

• When given ‘--get-word’, return the value of the word (two bytes) at offset bytes into the selected SMBIOS structure. It will be formatted as an unsigned decimal integer.

• When given ‘--get-dword’, return the value of the dword (four bytes) at offset bytes into the selected SMBIOS structure. It will be formatted as an unsigned decimal integer.

• When given ‘--get-qword’, return the value of the qword (eight bytes) at offset bytes into the selected SMBIOS structure. It will be formatted as an unsigned decimal integer.

• When given ‘--get-string’, return the string with its index found at offset bytes into the selected SMBIOS structure.

• When given ‘--get-uuid’, return the value of the UUID (sixteen bytes) at offset bytes into the selected SMBIOS structure. It will be formatted as lower-case hyphenated hexadecimal digits, with the first three fields as little-endian, and the rest printed byte-by-byte.

The default action is to print the value of the requested field to the console, but a variable name can be specified with ‘--set’ to store the value instead of printing it. For example, this will store and then display the system manufacturer’s name.

     smbios --type 1 --get-string 4 --set system_manufacturer
     echo $system_manufacturer

17.4.73 source

[Command] source file

Read file as a configuration file, as if its contents had been incorporated directly into the sourcing file. Unlike configfile (see Section 17.4.15 [configfile], page 81), this executes the contents of file without changing context: any environment variable changes made by the commands in file will be preserved after source returns, and the menu will not be shown immediately.

17.4.74 test

[Command] test expression Evaluate expression and return zero exit status if result is true, non zero status otherwise. expression is one of:

string1 == string2 the strings are equal

string1 != string2 the strings are not equal

string1 < string2 string1 is lexicographically less than string2

string1 <= string2 string1 is lexicographically less or equal than string2 string1 > string2

string1 is lexicographically greater than string2

string1 >= string2 string1 is lexicographically greater or equal than string2

integer1 -eq integer2 integer1 is equal to integer2

integer1 -ge integer2 integer1 is greater than or equal to integer2

integer1 -gt integer2 integer1 is greater than integer2

integer1 -le integer2 integer1 is less than or equal to integer2

integer1 -lt integer2 integer1 is less than integer2

integer1 -ne integer2 integer1 is not equal to integer2

prefixinteger1 -pgt prefixinteger2 integer1 is greater than integer2 after stripping off common non-numeric prefix.

prefixinteger1 -plt prefixinteger2 integer1 is less than integer2 after stripping off common non-numeric prefix.

file1 -nt file2 file1 is newer than file2 (modification time). Optionally numeric bias may be directly appended to -nt in which case it is added to the first file modification time.

file1 -ot file2 file1 is older than file2 (modification time). Optionally numeric bias may be directly appended to -ot in which case it is added to the first file modification time.

-d file file exists and is a directory

-e file file exists

-f file file exists and is not a directory

-s file file exists and has a size greater than zero

-n string the length of string is nonzero

string string is equivalent to -n string

-z string the length of string is zero

( expression )

expression is true ! expression

expression is false

expression1 -a expression2 both expression1 and expression2 are true

expression1 expression2 both expression1 and expression2 are true. This syntax is not POSIXcompliant and is not recommended.

expression1 -o expression2 either expression1 or expression2 is true

17.4.75 true

[Command] true

Do nothing, successfully. This is mainly useful in control constructs such as if and while (see Section 6.3 [Shell-like scripting], page 28).

17.4.76 trust

[Command] trust [‘--skip-sig’] pubkey file

Read public key from pubkey file and add it to GRUB’s internal list of trusted public keys. These keys are used to validate digital signatures when environment vari-

able check_signatures is set to enforce. Note that if check_signatures is set

to enforce when trust executes, then pubkey file must itself be properly signed. The ‘--skip-sig’ option can be used to disable signature-checking when reading pubkey file itself. It is expected that ‘--skip-sig’ is useful for testing and manual booting. See Section 19.2 [Using digital signatures], page 112, for more information.

17.4.77 unset

[Command] unset envvar Unset the environment variable envvar.

17.4.78 verify detached

[Command] verify_detached [‘--skip-sig’] file signature file [pubkey file]

Verifies a GPG-style detached signature, where the signed file is file, and the signature itself is in file signature file. Optionally, a specific public key to use can be specified using pubkey file. When environment variable check_signatures is set to enforce, then pubkey file must itself be properly signed by an already-trusted key. An unsigned pubkey file can be loaded by specifying ‘--skip-sig’. If pubkey file is omitted, then public keys from GRUB’s trusted keys (see Section 17.4.40 [list trusted], page 87, see Section 17.4.76 [trust], page 100, and see Section 17.4.22 [distrust], page 83) are tried.

Exit code $? is set to 0 if the signature validates successfully. If validation fails, it is set to a non-zero value. See Section 19.2 [Using digital signatures], page 112, for more information.

17.4.79 videoinfo

[Command] videoinfo [[WxH]xD]

List available video modes. If resolution is given, show only matching modes.

17.4.80 wrmsr

[Command:] wrmsr 0xADDR 0xVALUE Write a 0xVALUE to a model-specific register at address 0xADDR. Please note that on SMP systems, writing to a MSR that has a scope per hardware thread, implies that the value that is written only applies to the particular cpu/core/thread that runs the command. Also, if you specify a reserved or unimplemented MSR address, it will cause a general protection exception (which is not currently being handled) and the system will reboot. Note: The command is not allowed when lockdown is enforced (see Section 19.6 [Lockdown], page 114). This is done to prevent subverting various security mechanisms.

17.5 Networking commands

17.5.1 net add addr

[Command] net_add_addr interface card address

Configure additional network interface with address on a network card. address can be either IP in dotted decimal notation, or symbolic name which is resolved using DNS lookup. If successful, this command also adds local link routing entry to the default subnet of address with name interface‘:local’ via interface.

17.5.2 net add dns

[Command] net_add_dns server

Resolve server IP address and add to the list of DNS servers used during name lookup.

17.5.3 net add route

[Command] net_add_route shortname ip[/prefix] [interface | ‘gw’
   gateway]

Add route to network with address ip as modified by prefix via either local interface or gateway. prefix is optional and defaults to 32 for IPv4 address and 128 for IPv6 address. Route is identified by shortname which can be used to remove it (see Section 17.5.7 [net del route], page 102).

17.5.4 net bootp

[Command] net_bootp [card]

Alias for net dhcp, for compatibility with older Grub versions. Will perform the same DHCP handshake with potential fallback to BOOTP as the net dhcp command (see Section 17.5.8 [net dhcp], page 102).

17.5.5 net del addr

[Command] net_del_addr interface

Remove configured interface with associated address.

17.5.6 net del dns

[Command] net_del_dns address

Remove address from list of servers used during name lookup.

17.5.7 net del route

[Command] net_del_route shortname

Remove route entry identified by shortname.

17.5.8 net dhcp

[Command] net_dhcp [card]

Perform configuration of card using DHCP protocol. If no card name is specified, try to configure all existing cards. Falls back to the BOOTP protocol, if needed. If configuration was successful, interface with name card‘:dhcp’ and configured address is added to card. Additionally the following DHCP options are recognized and processed:

‘1 (Subnet Mask)’

Used to calculate network local routing entry for interface card‘:dhcp’.

‘3 (Router)’

Adds default route entry with the name card‘:dhcp:default’ via gateway from DHCP option. Note that only option with single route is accepted.

‘6 (Domain Name Server)’

Adds all servers from option value to the list of servers used during name resolution.

‘12 (Host Name)’

Sets environment variable ‘net_’<card>‘_dhcp_hostname’ (see Section 15.1.29 [net <interface> hostname], page 68) to the value of option.

‘15 (Domain Name)’

Sets environment variable ‘net_’<card>‘_dhcp_domain’ (see Section 15.1.27 [net <interface> domain], page 68) to the value of option.

‘17 (Root Path)’

Sets environment variable ‘net_’<card>‘_dhcp_rootpath’ (see Section 15.1.33 [net <interface> rootpath], page 68) to the value of option.

‘18 (Extensions Path)’

Sets environment variable ‘net_’<card>‘_dhcp_extensionspath’ (see Section 15.1.28 [net <interface> extensionspath], page 68) to the value of option.

‘66 (TFTP Server Name)’

Sets environment variable ‘net_’<card>‘_dhcp_server_name’ (see Section 15.1.26 [net <interface> dhcp server name], page 68) to the value of option.

‘67 (Filename)’

Sets environment variable ‘net_’<card>‘_boot_file’ (see Section 15.1.23 [net <interface> boot file], page 67) to the value of option.

17.5.9 net get dhcp option

[Command] net_get_dhcp_option var interface number type

Request DHCP option number of type via interface. type can be one of ‘string’, ‘number’ or ‘hex’. If option is found, assign its value to variable var. Values of types ‘number’ and ‘hex’ are converted to string representation.

17.5.10 net ipv6 autoconf

[Command] net_ipv6_autoconf [card]

Perform IPv6 autoconfiguration by adding to the card interface with name card‘:link’ and link local MAC-based address. If no card is specified, perform autoconfiguration for all existing cards.

17.5.11 net ls addr

[Command] net_ls_addr

List all configured interfaces with their MAC and IP addresses.

17.5.12 net ls cards

[Command] net_ls_cards

List all detected network cards with their MAC address.

17.5.13 net ls dns

[Command] net_ls_dns

List addresses of DNS servers used during name lookup.

17.5.14 net ls routes

[Command] net_ls_routes

List routing entries.

17.5.15 net nslookup

[Command] net_nslookup name [server]

Resolve address of name using DNS server server. If no server is given, use default list of servers.

17.5.16 net set vlan

[Command] net_set_vlan interface vlanid

Set the 802.1Q VLAN identifier on interface to vlanid. For example, to set the VLAN identifier on interface ‘efinet1’ to ‘100’:

         net_set_vlan efinet1 100

The VLAN identifier can be removed by setting it to ‘0’:

         net_set_vlan efinet1 0

17.6 Commands currently undocumented

Unfortunately, not all GRUB commands are documented at this time due to developer resource constraints. One way to contribute back to the GRUB project would be to help document these commands, and submit patches or ideas to the mailing list. The following is a (most likely incomplete) list of undocumented or poorly documented commands and not all of them are allowed for all platforms. Running the command help from within the GRUB shell may provide more information on parameters and usage.

• all_functional_test - Run all functional tests.

• backtrace - Print backtrace.

• boottime - Show boot time statistics.

• cacheinfo - Get disk cache info.

• cbmemc - Show CBMEM console content.

• cmosset - Set bit at BYTE:BIT in CMOS.

• coreboot_boottime - Show coreboot boot time statistics.

• dump - Show memory contents.

• efiemu_loadcore - Load and initialize EFI emulator.

• efiemu_prepare - Finalize loading of EFI emulator.

• efiemu_unload - Unload EFI emulator.

• exit - Exit from GRUB.

• extract_entries_configfile - Load another config file but take only menu entries.

• extract_entries_source - Load another config file without changing context but take only menu entries.

• extract_legacy_entries_configfile - Parse legacy config in new context taking only menu entries

• extract_legacy_entries_source - Parse legacy config in same context taking only menu entries

• extract_syslinux_entries_configfile - Execute syslinux config in new context taking only menu entries

• extract_syslinux_entries_source - Execute syslinux config in same context taking only menu entries

• fakebios - Create BIOS-like structures for backward compatibility with existing OS.

• file - Check if FILE is of specified type.

• fix_video - Fix video problem.

• fpswa - Display FPSWA version.

• functional_test - Run all loaded functional tests.

• gdbstub_break - Break into GDB

• gdbstub - Start GDB stub on given port

• gdbstub_stop - Stop GDB stub

• hdparm - Get/set ATA disk parameters.

• hexdump_random - Hexdump random data.

• inb - Read 8-bit value from PORT.

• inl - Read 32-bit value from PORT.

• inw - Read 16-bit value from PORT.

• jpegtest - Tests loading of JPEG bitmap.

• keymap - Load a keyboard layout.

• legacy_check_password - Simulate grub-legacy ‘password’ command in menu entry mode

• legacy_configfile - Parse legacy config in new context

• legacy_password - Simulate grub-legacy ‘password’ command

• legacy_source - Parse legacy config in same context

• loadbios - Load BIOS dump.

• lsacpi - Show ACPI information.

• lsapm - Show APM information.

• lscoreboot - List coreboot tables.

• lsdev - List devices.

• lsefi - Display EFI handles.

• lsefimmap - Display EFI memory map.

• lsefisystab - Display EFI system tables.

• lsmmap - List memory map provided by firmware.

• lspci - List PCI devices.

• lssal - Display SAL system table.

• lsspd - Print Memory information.

• macppcbless - Bless DIR of HFS or HFS+ partition for PPC macs.

• mactelbless - Bless FILE of HFS or HFS+ partition for intel macs.

• net_set_vlan - Set an interface’s vlan id.

• outb - Write 8-bit VALUE to PORT.

• outl - Write 32-bit VALUE to PORT.

• outw - Write 16-bit VALUE to PORT.

• pcidump - Show raw dump of the PCI configuration space.

• pngtest - Tests loading of PNG bitmap.

• read_byte - Read 8-bit value from ADDR.

• read_dword - Read 32-bit value from ADDR.

• read_word - Read 16-bit value from ADDR.

• setpci - Manipulate PCI devices.

• suspend - Return to IEEE1275 prompt.

• syslinux_configfile - Execute syslinux config in new context

• syslinux_source - Execute syslinux config in same context

• test_blockarg - Print and execute block argument., 0

• testload - Load the same file in multiple ways.

• testspeed - Test file read speed.

• tgatest - Tests loading of TGA bitmap.

• time - Measure time used by COMMAND

• tr - Translate SET1 characters to SET2 in STRING.

• usb - Test USB support.

• vbeinfo - List available video modes. If resolution is given show only modes matching it.

• vbetest - Test video subsystem.

• videotest - Test video subsystem in mode WxH.

• write_byte - Write 8-bit VALUE to ADDR.

• write_dword - Write 32-bit VALUE to ADDR.

• write_word - Write 16-bit VALUE to ADDR.

• xen_cat - List Xen storage.

• xen_ls - List Xen storage.

• xnu_devprop_load - Load ‘device-properties’ dump.

• xnu_uuid - Transform 64-bit UUID to format suitable for XNU. If -l is given keep it lowercase as done by blkid.

• zfs-bootfs - Print ZFS-BOOTFSOBJ or store it into VARIABLE

• zfsinfo - Print ZFS info about DEVICE.

• zfskey - Import ZFS wrapping key stored in FILE.

18 Internationalisation

18.1 Charset

GRUB uses UTF-8 internally other than in rendering where some GRUB-specific appropriate representation is used. All text files (including config) are assumed to be encoded in UTF-8.

18.2 Filesystems

NTFS, JFS, UDF, HFS+, exFAT, long filenames in FAT, Joliet part of ISO9660 are treated as UTF-16 as per specification. AFS and BFS are read as UTF-8, again according to specification. BtrFS, cpio, tar, squash4, minix, minix2, minix3, ROMFS, ReiserFS, XFS, ext2, ext3, ext4, FAT (short names), F2FS, RockRidge part of ISO9660, nilfs2, UFS1, UFS2 and ZFS are assumed to be UTF-8. This might be false on systems configured with legacy charset but as long as the charset used is superset of ASCII you should be able to access ASCII-named files. And it’s recommended to configure your system to use UTF-8 to access the filesystem, convmv may help with migration. ISO9660 (plain) filenames are specified as being ASCII or being described with unspecified escape sequences. GRUB assumes that the ISO9660 names are UTF-8 (since any ASCII is valid UTF-8). There are some old CD-ROMs which use CP437 in non-compliant way. You’re still able to access files with names containing only ASCII characters on such filesystems though. You’re also able to access any file if the filesystem contains valid Joliet (UTF-16) or RockRidge (UTF-8). AFFS, SFS and HFS never use unicode and GRUB assumes them to be in Latin1, Latin1 and MacRoman respectively. GRUB handles filesystem case-insensitivity however no attempt is performed at case conversion of international characters so e.g. a file named lowercase greek alpha is treated as different from the one named as uppercase alpha. The filesystems in questions are NTFS (except POSIX namespace), HFS+ (configurable at mkfs time, default insensitive), SFS (configurable at mkfs time, default insensitive), JFS (configurable at mkfs time, default sensitive), HFS, AFFS, FAT, exFAT and ZFS (configurable on per-subvolume basis by property “casesensitivity”, default sensitive). On ZFS subvolumes marked as case insensitive files containing lowercase international characters are inaccessible. Also like all supported filesystems except HFS+ and ZFS (configurable on per-subvolume basis by property “normalization”, default none) GRUB makes no attempt at check of canonical equivalence so a file name u-diaresis is treated as distinct from u+combining diaresis. This however means that in order to access file on HFS+ its name must be specified in normalisation form D. On normalized ZFS subvolumes filenames out of normalisation are inaccessible.

18.3 Output terminal

Firmware output console “console” on ARC and IEEE1275 are limited to ASCII. BIOS firmware console and VGA text are limited to ASCII and some pseudographics. None of above mentioned is appropriate for displaying international and any unsup-

ported character is replaced with question mark except pseudographics which we attempt to approximate with ASCII.

EFI console on the other hand nominally supports UTF-16 but actual language coverage depends on firmware and may be very limited.

The encoding used on serial can be chosen with terminfo as either ASCII, UTF-8 or “visual UTF-8”. Last one is against the specification but results in correct rendering of right-to-left on some readers which don’t have own bidi implementation.

On emu GRUB checks if charset is UTF-8 and uses it if so and uses ASCII otherwise. When using gfxterm or gfxmenu GRUB itself is responsible for rendering the text.

In this case GRUB is limited by loaded fonts. If fonts contain all required characters then bidirectional text, cursive variants and combining marks other than enclosing, half (e.g. left half tilde or combining overline) and double ones. Ligatures aren’t supported though. This should cover European, Middle Eastern (if you don’t mind lack of lam-alif ligature in Arabic) and East Asian scripts. Notable unsupported scripts are Brahmic family and derived as well as Mongolian, Tifinagh, Korean Jamo (precomposed characters have no problem) and tonal writing (2e5-2e9). GRUB also ignores deprecated (as specified in Unicode) characters (e.g. tags). GRUB also doesn’t handle so called “annotation characters” If you can complete either of two lists or, better, propose a patch to improve rendering, please contact developer team.

18.4 Input terminal

Firmware console on BIOS, IEEE1275 and ARC doesn’t allow you to enter non-ASCII characters. EFI specification allows for such but author is unaware of any actual implementations. Serial input is currently limited for latin1 (unlikely to change). Own keyboard implementations (at keyboard and usb keyboard) supports any key but work on one-char-per-keystroke. So no dead keys or advanced input method. Also there is no keymap change hotkey. In practice it makes difficult to enter any text using non-Latin alphabet. Moreover all current input consumers are limited to ASCII.

18.5 Gettext

GRUB supports being translated. For this you need to have language *.mo files in $prefix/locale, load gettext module and set “lang” variable.

18.6 Regexp

Regexps work on unicode characters, however no attempt at checking cannonical equivalence has been made. Moreover the classes like [:alpha:] match only ASCII subset.

18.7 Other

Currently GRUB always uses YEAR-MONTH-DAY HOUR:MINUTE:SECOND [WEEK-DAY] 24-hour datetime format but weekdays are translated. GRUB always uses the decimal number format with [0-9] as digits and . as descimal separator and no group separator. IEEE1275 aliases are matched case-insensitively except non-ASCII which is matched as binary. Similar behaviour is for matching OSBundleRequired. Since IEEE1275 aliases and OSBundleRequired don’t contain any non-ASCII it should never be a problem in practice. Case-sensitive identifiers are matched as raw strings, no canonical equivalence check is performed. Case-insenstive identifiers are matched as RAW but additionally [a-z] is equivalent to [A-Z]. GRUB-defined identifiers use only ASCII and so should user-defined ones. Identifiers containing non-ASCII may work but aren’t supported. Only the ASCII space characters (space U+0020, tab U+000b, CR U+000d and LF U+000a) are recognised. Other unicode space characters aren’t a valid field separator. test (see Section 17.4.74 [test], page 98) tests <, >, <=, >=, -pgt and -plt compare the strings in the lexicographical order of unicode codepoints, replicating the behaviour of test from coreutils. environment variables and commands are listed in the same order.

19 Security

19.1 Authentication and authorisation in GRUB

By default, the boot loader interface is accessible to anyone with physical access to the console: anyone can select and edit any menu entry, and anyone can get direct access to a GRUB shell prompt. For most systems, this is reasonable since anyone with direct physical access has a variety of other ways to gain full access, and requiring authentication at the boot loader level would only serve to make it difficult to recover broken systems.

However, in some environments, such as kiosks, it may be appropriate to lock down the boot loader to require authentication before performing certain operations.

The ‘password’ (see Section 17.4.54 [password], page 90) and ‘password_pbkdf2’ (see Section 17.4.55 [password pbkdf2], page 90) commands can be used to define users, each of which has an associated password. ‘password’ sets the password in plain text, requiring ‘grub.cfg’ to be secure; ‘password_pbkdf2’ sets the password hashed using the Password-Based Key Derivation Function (RFC 2898), requiring the use of grub-mkpasswd-pbkdf2 (see Section 24.3 [Invoking grub-mkpasswd-pbkdf2], page 130) to generate password hashes.

In order to enable authentication support, the ‘superusers’ environment variable must be set to a list of usernames, separated by any of spaces, commas, semicolons, pipes, or ampersands. Superusers are permitted to use the GRUB command line, edit menu entries, and execute any menu entry. If ‘superusers’ is set, then use of the command line and editing of menu entries are automatically restricted to superusers. Setting ‘superusers’ to empty string effectively disables both access to CLI and editing of menu entries. Note: The environment variable needs to be exported to also affect the section defined by the ‘submenu’ command (see Section 17.1.2 [submenu], page 74).

Other users may be allowed to execute specific menu entries by giving a list of usernames (as above) using the ‘--users’ option to the ‘menuentry’ command (see Section 17.1.1 [menuentry], page 73). If the ‘--unrestricted’ option is used for a menu entry, then that entry is unrestricted. If the ‘--users’ option is not used for a menu entry, then that only superusers are able to use it.

Putting this together, a typical ‘grub.cfg’ fragment might look like this:

     set superusers="root"
     password_pbkdf2 root grub.pbkdf2.sha512.10000.biglongstring
     password user1 insecure
     menuentry "May be run by any user" --unrestricted {
     set root=(hd0,1)
     linux /vmlinuz
     }
     menuentry "Superusers only" --users "" {
     set root=(hd0,1)
     linux /vmlinuz single
     }
     menuentry "May be run by user1 or a superuser" --users user1 {
     set root=(hd0,2)
     chainloader +1
     }

The grub-mkconfig program does not yet have built-in support for generating configuration files with authentication. You can use ‘/etc/grub.d/40_custom’ to add simple

superuser authentication, by adding set superusers= and password or password_pbkdf2

commands.

19.2 Using digital signatures in GRUB

GRUB’s ‘core.img’ can optionally provide enforcement that all files subsequently read from disk are covered by a valid digital signature. This document does not cover how to ensure that your platform’s firmware (e.g., Coreboot) validates ‘core.img’.

If environment variable check_signatures (see Section 15.1.2 [check signatures], page 63) is set to enforce, then every attempt by the GRUB ‘core.img’ to load another file ‘foo’ implicitly invokes verify_detached foo foo.sig (see Section 17.4.78 [verify detached], page 100). foo.sig must contain a valid digital signature over the contents of foo, which can be verified with a public key currently trusted by GRUB (see Section 17.4.40 [list trusted], page 87, see Section 17.4.76 [trust], page 100, and see Section 17.4.22 [distrust], page 83). If validation fails, then file ‘foo’ cannot be opened. This failure may halt or otherwise impact the boot process.

An initial trusted public key can be embedded within the GRUB ‘core.img’ using the --pubkey option to grub-install (see Section 24.1 [Invoking grub-install], page 129).

GRUB uses GPG-style detached signatures (meaning that a file ‘foo.sig’ will be produced when file ‘foo’ is signed), and currently supports the DSA and RSA signing algorithms. A signing key can be generated as follows:

     gpg --gen-key

An individual file can be signed as follows:

     gpg --detach-sign /path/to/file

For successful validation of all of GRUB’s subcomponents and the loaded OS kernel, they must all be signed. One way to accomplish this is the following (after having already produced the desired ‘grub.cfg’ file, e.g., by running grub-mkconfig (see Section 24.2 [Invoking grub-mkconfig], page 130):

     # Edit /dev/shm/passphrase.txt to contain your signing key's passphrase
     for i in `find /boot -name "*.cfg" -or -name "*.lst" -or \
       -name "*.mod" -or -name "vmlinuz*" -or -name "initrd*" -or \
       -name "grubenv"`;
     do
       gpg --batch --detach-sign --passphrase-fd 0 $i < \
         /dev/shm/passphrase.txt
     done
     shred /dev/shm/passphrase.txt

See also: Section 15.1.2 [check signatures], page 63, Section 17.4.78 [verify detached], page 100, Section 17.4.76 [trust], page 100, Section 17.4.40 [list trusted], page 87, Section 17.4.22 [distrust], page 83, Section 17.4.41 [load env], page 87, Section 17.4.64 [save env], page 93.

Note that internally signature enforcement is controlled by setting the environment variable check_signatures equal to enforce. Passing one or more --pubkey options to

grub-mkimage implicitly defines check_signatures equal to enforce in ‘core.img’ prior

to processing any configuration files.

Note that signature checking does not prevent an attacker with (serial, physical, ...) console access from dropping manually to the GRUB console and executing:

     set check_signatures=no

To prevent this, password-protection (see Section 19.1 [Authentication and authorisation], page 111) is essential. Note that even with GRUB password protection, GRUB itself cannot prevent someone with physical access to the machine from altering that machine’s firmware (e.g., Coreboot or BIOS) configuration to cause the machine to boot from a different (attacker-controlled) device. GRUB is at best only one link in a secure boot chain.

19.3 UEFI secure boot and shim support

The GRUB, except the chainloader command, works with the UEFI secure boot and the shim. This functionality is provided by the shim lock verifier. It is built into the ‘core.img’ and is registered if the UEFI secure boot is enabled. The ‘shim_lock’ variable is set to ‘y’ when shim lock verifier is registered. If it is desired to use UEFI secure boot without shim, one can disable shim lock by disabling shim verification with MokSbState UEFI variable or by building grub image with ‘--disable-shim-lock’ option.

All GRUB modules not stored in the ‘core.img’, OS kernels, ACPI tables, Device Trees, etc. have to be signed, e.g, using PGP. Additionally, the commands that can be used to subvert the UEFI secure boot mechanism, such as iorw and memrw will not be available when the UEFI secure boot is enabled. This is done for security reasons and are enforced by the GRUB Lockdown mechanism (see Section 19.6 [Lockdown], page 114).

19.4 Embedded information for generation number based revocation

The Secure Boot Advanced Targeting (SBAT) is a mechanism to allow the revocation of components in the boot path by using generation numbers embedded into the EFI binaries. The SBAT metadata is located in an .sbat data section that has set of UTF-8 strings as comma-separated values (CSV). See

https://github.com/rhboot/shim/blob/main/SBAT.md for more details.

To add a data section containing the SBAT information into the binary, the ‘--sbat’ option of grub-mkimage command should be used. The content of a CSV file, encoded with UTF-8, is copied as is to the .sbat data section into the generated EFI binary. The CSV file can be stored anywhere on the file system.

     grub-mkimage -O x86_64-efi -o grubx64.efi -p '(tftp)/grub' --sbat sbat.csv efinet tftp

19.5 Measuring boot components

If the tpm module is loaded and the platform has a Trusted Platform Module installed, GRUB will log each command executed and each file loaded into the TPM event log and extend the PCR values in the TPM correspondingly. All events will be logged into the PCR described below with a type of EV IPL and an event description as described below.

Event type  PCR Description

Command 8 All executed commands (including those from configuration files) will be logged and measured as entered with a prefix of “grub cmd: “

Kernel command line 8 Any command line passed to a kernel will be logged and measured as entered with a prefix of “kernel cmdline: ”

Module command line 8 Any command line passed to a kernel module will be logged and measured as entered with a prefix of “module cmdline: “

Files 9 Any file read by GRUB will be logged and measured with a descriptive text corresponding to the filename.

GRUB will not measure its own ‘core.img’ - it is expected that firmware will carry this out. GRUB will also not perform any measurements until the tpm module is loaded. As such it is recommended that the tpm module be built into ‘core.img’ in order to avoid a potential gap in measurement between ‘core.img’ being loaded and the tpm module being loaded.

Measured boot is currently only supported on EFI and IBM IEEE1275 PowerPC platforms.

19.6 Lockdown when booting on a secure setup

The GRUB can be locked down when booted on a secure boot environment, for example if the UEFI secure boot is enabled. On a locked down configuration, the GRUB will be restricted and some operations/commands cannot be executed.

Chapter 19: Security 115

The ‘lockdown’ variable is set to ‘y’ when the GRUB is locked down. Otherwise it does not exit.

20 Platform limitations

GRUB2 is designed to be portable and is actually ported across platforms. We try to keep all platforms at the level. Unfortunately some platforms are better supported than others. This is detailed in current and 2 following sections.

All platforms have an artificially GRUB imposed disk size restriction of 1 EiB. In some cases, larger disk sizes can be used, but access will not be allowed beyond 1 EiB.

LUKS2 devices with size larger than 16 EiB are currently not supported. They can not be created as crypto devices by cryptomount, so can not even be partially read from. LUKS have no limitations other than those imposed by the format.

ARC platform is unable to change datetime (firmware doesn’t seem to provide a function for it). EMU has similar limitation.

On EMU platform no serial port is available.

Console charset refers only to firmware-assisted console. gfxterm is always Unicode (see Internationalisation section for its limitations). Serial is configurable to UTF-8 or ASCII (see Internationalisation). In case of qemu and coreboot ports the refered console is vga text. Loongson always uses gfxterm.

Most limited one is ASCII. CP437 provides additionally pseudographics. GRUB2 doesn’t use any language characters from CP437 as often CP437 is replaced by national encoding compatible only in pseudographics. Unicode is the most versatile charset which supports many languages. However the actual console may be much more limited depending on firmware

On BIOS, network is supported only if the image is loaded through network. On sparc64, GRUB is unable to determine which server it was booted from.

Direct ATA/AHCI support allows to circumvent various firmware limitations but isn’t needed for normal operation except on baremetal ports.

AT keyboard support allows keyboard layout remapping and support for keys not available through firmware. It isn’t needed for normal operation except baremetal ports.

Speaker allows morse and spkmodem communication.

USB support provides benefits similar to ATA (for USB disks) or AT (for USB keyboards). In addition it allows USBserial.

Chainloading refers to the ability to load another bootloader through the same protocol

Hints allow faster disk discovery by already knowing in advance which is the disk in question. On some platforms hints are correct unless you move the disk between boots. On other platforms it’s just an educated guess. Note that hint failure results in just reduced performance, not a failure

BadRAM is the ability to mark some of the RAM as “bad”. Note: due to protocol limitations mips-loongson (with Linux protocol) and mips-qemu mips can use only memory up to first hole.

Bootlocation is ability of GRUB to automatically detect where it boots from. “disk” means the detection is limited to detecting the disk with partition being discovered on install time. “partition” means that disk and partiton can be automatically discovered.

“file” means that boot image file name as well as disk and partition can be discovered. For consistency, default install ignores partition and relies solely on disk detection. If no bootlocation discovery is available or boot and grub-root disks are different, UUID is used instead. On ARC if no device to install to is specified, UUID is used instead as well.

BIOS  Coreboot  Multiboot  Qemu
video  yes  yes  yes  yes
console charset  CP437  CP437  CP437  CP437
network  yes (*)  no  no  no
ATA/AHCI  yes  yes  yes  yes
AT keyboard  yes  yes  yes  yes
Speaker  yes  yes  yes  yes
USB  yes  yes  yes  yes
chainloader  local  yes  yes  no
cpuid  partial  partial  partial  partial
rdmsr  partial  partial  partial  partial
wrmsr  partial  partial  partial  partial
hints  guess  guess  guess  guess
PCI  yes  yes  yes  yes
badram  yes  yes  yes  yes
compression  always  pointless  no  no
exit  yes  no  no  no
bootlocation  disk  no  no  no
ia32 EFI  amd64 EFI  ia32 IEEE1275 Itanium
video  yes  yes  no  no
console charset  Unicode  Unicode  ASCII  Unicode
network  yes  yes  yes  yes
ATA/AHCI  yes  yes  yes  no
AT keyboard  yes  yes  yes  no
Speaker  yes  yes  yes  no
USB  yes  yes  yes  no
chainloader  local  local  no  local
cpuid  partial  partial  partial  no
rdmsr  partial  partial  partial  no
wrmsr  partial  partial  partial  no
hints  guess  guess  good  guess
PCI  yes  yes  yes  no
badram  yes  yes  no  yes
compression  no  no  no  no
exit  yes  yes  yes  yes
bootlocation  file  file  file, ignored  file
Loongson  sparc64  Powerpc  ARC
video  yes  no  yes  no
console charset  N/A  ASCII  ASCII  ASCII
network  no  yes (*)  yes  no
ATA/AHCI  yes  no  no  no
AT keyboard  yes  no  no  no

Speaker  no  no  no  no
USB  yes  no  no  no
chainloader  yes  no  no  no
cpuid  no  no  no  no
rdmsr  no  no  no  no
wrmsr  no  no  no  no
hints  good  good  good  no
PCI  yes  no  no  no
badram  yes (*)  no  no  no
compression  configurable  no  no  configurable
exit  no  yes  yes  yes
bootlocation  no  partition  file  file (*)
MIPS qemu  emu  xen
video  no  yes  no
console charset  CP437  Unicode (*)  ASCII
network  no  yes  no
ATA/AHCI  yes  no  no
AT keyboard  yes  no  no
Speaker  no  no  no
USB  N/A  yes  no
chainloader  yes  no  yes
cpuid  no  no  yes
rdmsr  no  no  yes
wrmsr  no  no  yes
hints  guess  no  no
PCI  no  no  no
badram  yes (*)  no  no
compression  configurable  no  no
exit  no  yes  no
bootlocation  no  file  no

21 Outline

Some platforms have features which allows to implement some commands useless or not implementable on others.

Quick summary: Information retrieval:

• mipsel-loongson: lsspd

• mips-arc: lsdev

• efi: lsefisystab, lssal, lsefimmap, lsefi

• i386-pc: lsapm

• i386-coreboot: lscoreboot, coreboot boottime, cbmemc

• acpi-enabled (i386-pc, i386-coreboot, i386-multiboot, *-efi): lsacpi Workarounds for platform-specific issues:

• i386-efi/x86 64-efi: loadbios, fakebios, fix video

• acpi-enabled (i386-pc, i386-coreboot, i386-multiboot, *-efi): acpi (override ACPI tables)

• i386-pc: drivemap

• i386-pc: sendkey Advanced operations for power users:

• x86: iorw (direct access to I/O ports)

Miscelaneous:

• cmos (x86-*, ieee1275, mips-qemu mips, mips-loongson): cmostest (used on some laptops to check for special power-on key), cmosclean

• i386-pc: play

22 Supported boot targets

X86 support is summarised in the following table. “Yes” means that the kernel works on the given platform, “crashes” means an early kernel crash which we hope will be fixed by concerned kernel developers. “no” means GRUB doesn’t load the given kernel on a given platform. “headless” means that the kernel works but lacks console drivers (you can still use serial or network console). In case of “no” and “crashes” the reason is given in footnote.

BIOS  Coreboot
BIOS chainloading  yes  no (1)
NTLDR  yes  no (1)
Plan9  yes  no (1)
Freedos  yes  no (1)
FreeBSD bootloader  yes  crashes (1)
32-bit kFreeBSD  yes  crashes (5)
64-bit kFreeBSD  yes  crashes (5)
32-bit kNetBSD  yes  crashes (1)
64-bit kNetBSD  yes  crashes
32-bit kOpenBSD  yes  yes
64-bit kOpenBSD  yes  yes
Multiboot  yes  yes
Multiboot2  yes  yes
32-bit Linux (legacy protocol)  yes  no (1)
64-bit Linux (legacy protocol)  yes  no (1)
32-bit Linux (modern protocol)  yes  yes
64-bit Linux (modern protocol)  yes  yes
32-bit XNU  yes  ?
64-bit XNU  yes  ?
32-bit EFI chainloader  no (2)  no (2)
64-bit EFI chainloader  no (2)  no (2)
Appleloader  no (2)  no (2)
Multiboot  Qemu
BIOS chainloading  no (1)  no (1)
NTLDR  no (1)  no (1)
Plan9  no (1)  no (1)
FreeDOS  no (1)  no (1)
FreeBSD bootloader  crashes (1)  crashes (1)
32-bit kFreeBSD  crashes (5)  crashes (5)
64-bit kFreeBSD  crashes (5)  crashes (5)
32-bit kNetBSD  crashes (1)  crashes (1)
64-bit kNetBSD  yes  yes
32-bit kOpenBSD  yes  yes
64-bit kOpenBSD  yes  yes
Multiboot  yes  yes
Multiboot2  yes  yes
32-bit Linux (legacy protocol)  no (1)  no (1)
64-bit Linux (legacy protocol)  no (1)  no (1)

32-bit Linux (modern protocol)  yes  yes
64-bit Linux (modern protocol)  yes  yes
32-bit XNU  ?  ?
64-bit XNU  ?  ?
32-bit EFI chainloader  no (2)  no (2)
64-bit EFI chainloader  no (2)  no (2)
Appleloader  no (2)  no (2)
ia32 EFI  amd64 EFI
BIOS chainloading  no (1)  no (1)
NTLDR  no (1)  no (1)
Plan9  no (1)  no (1)
FreeDOS  no (1)  no (1)
FreeBSD bootloader  crashes (1)  crashes (1)
32-bit kFreeBSD  headless  headless
64-bit kFreeBSD  headless  headless
32-bit kNetBSD  crashes (1)  crashes (1)
64-bit kNetBSD  yes  yes
32-bit kOpenBSD  headless  headless
64-bit kOpenBSD  headless  headless
Multiboot  yes  yes
Multiboot2  yes  yes
32-bit Linux (legacy protocol)  no (1)  no (1)
64-bit Linux (legacy protocol)  no (1)  no (1)
32-bit Linux (modern protocol)  yes  yes
64-bit Linux (modern protocol)  yes  yes
32-bit XNU  yes  yes
64-bit XNU  yes (4)  yes
32-bit EFI chainloader  yes  no (3)
64-bit EFI chainloader  no (3)  yes
Appleloader  yes  yes

ia32 IEEE1275

BIOS chainloading  no (1)
NTLDR  no (1)
Plan9  no (1)
FreeDOS  no (1)
FreeBSD bootloader  crashes (1)
32-bit kFreeBSD  crashes (5)
64-bit kFreeBSD  crashes (5)
32-bit kNetBSD  crashes (1)
64-bit kNetBSD  ?
32-bit kOpenBSD  ?
64-bit kOpenBSD  ?
Multiboot  ?
Multiboot2  ?
32-bit Linux (legacy protocol)  no (1)
64-bit Linux (legacy protocol)  no (1)

32-bit Linux (modern protocol)  ?
64-bit Linux (modern protocol)  ?
32-bit XNU  ?
64-bit XNU  ?
32-bit EFI chainloader  no (2)
64-bit EFI chainloader  no (2)
Appleloader  no (2)

1. Requires BIOS

2. EFI only

3. 32-bit and 64-bit EFI have different structures and work in different CPU modes so it’s not possible to chainload 32-bit bootloader on 64-bit platform and vice-versa

4. Some modules may need to be disabled

5. Requires ACPI

PowerPC, IA64 and Sparc64 ports support only Linux. MIPS port supports Linux and multiboot2.

22.1 Boot tests

As you have seen in previous chapter the support matrix is pretty big and some of the configurations are only rarely used. To ensure the quality bootchecks are available for all x86 targets except EFI chainloader, Appleloader and XNU. All x86 platforms have bootcheck facility except ieee1275. Multiboot, multiboot2, BIOS chainloader, ntldr and freebsd-bootloader boot targets are tested only with a fake kernel images. Only Linux is tested among the payloads using Linux protocols.

Following variables must be defined:

GRUB PAYLOADS DIR directory containing the required kernels

GRUB CBFSTOOL cbfstool from Coreboot package (for coreboot platform only)

GRUB COREBOOT ROM empty Coreboot ROM
GRUB QEMU OPTS  additional options to be supplied to QEMU

Required files are:

kfreebsd env.i386  32-bit kFreeBSD device hints
kfreebsd.i386  32-bit FreeBSD kernel image

kfreebsd.x86 64, kfreebsd env.x86 64

same from 64-bit kFreeBSD

knetbsd.i386  32-bit NetBSD kernel image
knetbsd.miniroot.i386  32-bit kNetBSD miniroot.kmod.

knetbsd.x86 64, knetbsd.miniroot.x86 64

same from 64-bit kNetBSD

kopenbsd.i386  32-bit OpenBSD kernel bsd.rd image
kopenbsd.x86 64  same from 64-bit kOpenBSD
linux.i386  32-bit Linux
linux.x86 64  64-bit Linux

23 Error messages produced by GRUB

23.1 GRUB only offers a rescue shell

GRUB’s normal start-up procedure involves setting the ‘prefix’ environment variable to a value set in the core image by grub-install, setting the ‘root’ variable to match, loading the ‘normal’ module from the prefix, and running the ‘normal’ command (see Section 17.4.51 [normal], page 89). This command is responsible for reading ‘/boot/grub/grub.cfg’, running the menu, and doing all the useful things GRUB is supposed to do.

If, instead, you only get a rescue shell, this usually means that GRUB failed to load the ‘normal’ module for some reason. It may be possible to work around this temporarily: for instance, if the reason for the failure is that ‘prefix’ is wrong (perhaps it refers to the wrong device, or perhaps the path to ‘/boot/grub’ was not correctly made relative to the device), then you can correct this and enter normal mode manually:

 # Inspect the current prefix (and other preset variables):
 set
 # Find out which devices are available:
 ls
 # Set to the correct value, which might be something like this:
 set prefix=(hd0,1)/grub
 set root=(hd0,1)
 insmod normal
 normal

However, any problem that leaves you in the rescue shell probably means that GRUB was not correctly installed. It may be more useful to try to reinstall it properly using grubinstall device (see Section 24.1 [Invoking grub-install], page 129). When doing this, there are a few things to remember:

• Drive ordering in your operating system may not be the same as the boot drive ordering used by your firmware. Do not assume that your first hard drive (e.g. ‘/dev/sda’) is the one that your firmware will boot from. ‘device.map’ (see Section 4.3 [Device map], page 13) can be used to override this, but it is usually better to use UUIDs or file system labels and avoid depending on drive ordering entirely.

• At least on BIOS systems, if you tell grub-install to install GRUB to a partition but GRUB has already been installed in the master boot record, then the GRUB installation in the partition will be ignored.

• If possible, it is generally best to avoid installing GRUB to a partition (unless it is a special partition for the use of GRUB alone, such as the BIOS Boot Partition used on GPT). Doing this means that GRUB may stop being able to read its core image due to a file system moving blocks around, such as while defragmenting, running checks, or even during normal operation. Installing to the whole disk device is normally more robust.

• Check that GRUB actually knows how to read from the device and file system containing ‘/boot/grub’. It will not be able to read from encrypted devices with unsupported encryption scheme, nor from file systems for which support has not yet been added to GRUB.

23.2 Firmware stalls instead of booting GRUB

The EFI implementation of some older MacBook laptops stalls when it gets presented a grub-mkrescue ISO image for x86 64-efi target on an USB stick. Affected are models of year 2010 or earlier. Workaround is to zeroize the bytes 446 to 461 of the EFI partition, where mformat has put a partition table entry which claims partition start at block 0. This change will not hamper bootability on other machines.

24 User-space utilities

24.1 Invoking grub-install

The program grub-install generates a GRUB core image using grub-mkimage and installs it on your system. You must specify the device name on which you want to install GRUB, like this:

grub-install install_device

The device name install device is an OS device name or a GRUB device name.

grub-install accepts the following options:

‘--help’ Print a summary of the command-line options and exit.

‘--version’

Print the version number of GRUB and exit.

‘--boot-directory=dir’

Install GRUB images under the directory ‘dir/grub/’ This option is useful when you want to install GRUB into a separate partition or a removable disk. If this option is not specified then it defaults to ‘/boot’, so

     grub-install /dev/sda

is equivalent to

     grub-install --boot-directory=/boot/ /dev/sda

Here is an example in which you have a separate boot partition which is mounted on ‘/mnt/boot’:

     grub-install --boot-directory=/mnt/boot /dev/sdb
‘--recheck’

Recheck the device map, even if ‘/boot/grub/device.map’ already exists. You should use this option whenever you add/remove a disk into/from your computer.

‘--no-rs-codes’

By default on x86 BIOS systems, grub-install will use some extra space in the bootloader embedding area for Reed-Solomon error-correcting codes. This enables GRUB to still boot successfully if some blocks are corrupted. The exact amount of protection offered is dependent on available space in the embedding area. R sectors of redundancy can tolerate up to R/2 corrupted sectors. This redundancy may be cumbersome if attempting to cryptographically validate the contents of the bootloader embedding area, or in more modern systems with GPT-style partition tables (see Section 4.4 [BIOS installation], page 13) where GRUB does not reside in any unpartitioned space outside of the MBR. Disable the Reed-Solomon codes with this option.

24.2 Invoking grub-mkconfig

The program grub-mkconfig generates a configuration file for GRUB (see Section 6.1 [Simple configuration], page 21).

     grub-mkconfig -o /boot/grub/grub.cfg

grub-mkconfig accepts the following options:

‘--help’ Print a summary of the command-line options and exit.

‘--version’

Print the version number of GRUB and exit.

‘-o file’
‘--output=file’

Send the generated configuration file to file. The default is to send it to standard output.

24.3 Invoking grub-mkpasswd-pbkdf2

The program grub-mkpasswd-pbkdf2 generates password hashes for GRUB (see Chapter 19 [Security], page 111).

     grub-mkpasswd-pbkdf2

grub-mkpasswd-pbkdf2 accepts the following options:

‘-c number’
‘--iteration-count=number’

Number of iterations of the underlying pseudo-random function. Defaults to 10000.

‘-l number’
‘--buflen=number’

Length of the generated hash. Defaults to 64.

‘-s number’
‘--salt=number’

Length of the salt. Defaults to 64.

24.4 Invoking grub-mkrelpath

The program grub-mkrelpath makes a file system path relative to the root of its containing file system. For instance, if ‘/usr’ is a mount point, then:

     $ grub-mkrelpath /usr/share/grub/unicode.pf2
     `/share/grub/unicode.pf2'

This is mainly used internally by other GRUB utilities such as grub-mkconfig (see Section 24.2 [Invoking grub-mkconfig], page 130), but may occasionally also be useful for debugging.

grub-mkrelpath accepts the following options:

‘--help’ Print a summary of the command-line options and exit.

‘--version’

Print the version number of GRUB and exit.

24.5 Invoking grub-mkrescue

The program grub-mkrescue generates a bootable GRUB rescue image (see Section 4.2 [Making a GRUB bootable CD-ROM], page 12).

     grub-mkrescue -o grub.iso

All arguments not explicitly listed as grub-mkrescue options are passed on directly to xorriso in mkisofs emulation mode. Options passed to xorriso will normally be interpreted as mkisofs options; if the option ‘--’ is used, then anything after that will be interpreted as native xorriso options.

Non-option arguments specify additional source directories. This is commonly used to add extra files to the image:

     mkdir -p disk/boot/grub

(add extra files to ‘disk/boot/grub’)

     grub-mkrescue -o grub.iso disk

grub-mkrescue accepts the following options:

‘--help’ Print a summary of the command-line options and exit.

‘--version’

Print the version number of GRUB and exit.

‘-o file’
‘--output=file’

Save output in file. This "option" is required.

‘--modules=modules’

Pre-load the named GRUB modules in the image. Multiple entries in modules should be separated by whitespace (so you will probably need to quote this for your shell).

‘--rom-directory=dir’

If generating images for the QEMU or Coreboot platforms, copy the resulting ‘qemu.img’ or ‘coreboot.elf’ files respectively to the dir directory as well as including them in the image.

‘--xorriso=file’

Use file as the xorriso program, rather than the built-in default.

‘--grub-mkimage=file’

Use file as the grub-mkimage program, rather than the built-in default.

24.6 Invoking grub-mount

The program grub-mount performs a read-only mount of any file system or file system image that GRUB understands, using GRUB’s file system drivers via FUSE. (It is only available if FUSE development files were present when GRUB was built.) This has a number of uses:

• It provides a convenient way to check how GRUB will view a file system at boot time. You can use normal command-line tools to compare that view with that of your operating system, making it easy to find bugs.

• It offers true read-only mounts. Linux does not have these for journalling file systems, because it will always attempt to replay the journal at mount time; while you can temporarily mark the block device read-only to avoid this, that causes the mount to fail. Since GRUB intentionally contains no code for writing to file systems, it can easily provide a guaranteed read-only mount mechanism.

• It allows you to examine any file system that GRUB understands without needing to load additional modules into your running kernel, which may be useful in constrained environments such as installers.

• Since it can examine file system images (contained in regular files) just as easily as file systems on block devices, you can use it to inspect any file system image that GRUB understands with only enough privileges to use FUSE, even if nobody has yet written a FUSE module specifically for that file system type.

Using grub-mount is normally as simple as:

     grub-mount /dev/sda1 /mnt

grub-mount must be given one or more images and a mount point as non-option arguments (if it is given more than one image, it will treat them as a RAID set), and also accepts the following options:

‘--help’ Print a summary of the command-line options and exit.

‘--version’

Print the version number of GRUB and exit.

‘-C’
‘--crypto’

Mount encrypted devices, prompting for a passphrase if necessary.

‘-d string’
‘--debug=string’

Show debugging output for conditions matching string.

‘-K prompt|file’
‘--zfs-key=prompt|file’

Load a ZFS encryption key. If you use ‘prompt’ as the argument, grub-mount will read a passphrase from the terminal; otherwise, it will read key material from the specified file.

‘-r device’
‘--root=device’

Set the GRUB root device to device. You do not normally need to set this; grub-mount will automatically set the root device to the root of the supplied file system. If device is just a number, then it will be treated as a partition number within the supplied image. This means that, if you have an image of an entire disk in ‘disk.img’, then you can use this command to mount its second partition:

               grub-mount -r 2 disk.img mount-point
‘-v’
‘--verbose’

Print verbose messages.

24.7 Invoking grub-probe

The program grub-probe probes device information for a given path or device.

grub-probe --target=fs /boot/grub
grub-probe --target=drive --device /dev/sda1

grub-probe must be given a path or device as a non-option argument, and also accepts the following options:

‘--help’ Print a summary of the command-line options and exit.

‘--version’

Print the version number of GRUB and exit.

‘-d’
‘--device’

If this option is given, then the non-option argument is a system device name (such as ‘/dev/sda1’), and grub-probe will print information about that device. If it is not given, then the non-option argument is a filesystem path (such as ‘/boot/grub’), and grub-probe will print information about the device containing that part of the filesystem.

‘-m file’
‘--device-map=file’

Use file as the device map (see Section 4.3 [Device map], page 13) rather than

the default, usually ‘/boot/grub/device.map’.
‘-t target’
‘--target=target’

Print information about the given path or device as defined by target. The available targets and their meanings are:

‘fs’ GRUB filesystem module.

‘fs_uuid’ Filesystem Universally Unique Identifier (UUID).

‘fs_label’

Filesystem label.

‘drive’ GRUB device name.

‘device’ System device name.

‘partmap’ GRUB partition map module.

‘abstraction’

GRUB abstraction module (e.g. ‘lvm’).

‘cryptodisk_uuid’

Crypto device UUID.

‘msdos_parttype’

MBR partition type code (two hexadecimal digits).

‘hints_string’

A string of platform search hints suitable for passing to the search command (see Section 17.4.65 [search], page 93).

          ‘bios_hints’

Search hints for the PC BIOS platform.

          ‘ieee1275_hints’

Search hints for the IEEE1275 platform.

          ‘baremetal_hints’

Search hints for platforms where disks are addressed directly rather than via firmware.

          ‘efi_hints’

Search hints for the EFI platform.

          ‘arc_hints’

Search hints for the ARC platform.

          ‘compatibility_hint’

A guess at a reasonable GRUB drive name for this device, which may be used as a fallback if the search command fails.

‘disk’ System device name for the whole disk.

‘-v’
‘--verbose’

Print verbose messages.

24.8 Invoking grub-script-check

The program grub-script-check takes a GRUB script file (see Section 6.3 [Shell-like scripting], page 28) and checks it for syntax errors, similar to commands such as sh -n. It may take a path as a non-option argument; if none is supplied, it will read from standard input.

     grub-script-check /boot/grub/grub.cfg

grub-script-check accepts the following options:

‘--help’ Print a summary of the command-line options and exit.

‘--version’

Print the version number of GRUB and exit.

‘-v’
‘--verbose’

Print each line of input after reading it.

Appendix A How to obtain and build GRUB

Caution: GRUB requires binutils-2.9.1.0.23 or later because the GNU assembler has been changed so that it can produce real 16bits machine code between

2.9.1 and 2.9.1.0.x. See http://sources.redhat.com/binutils/, to obtain

information on how to get the latest version. GRUB is available from the GNU alpha archive site ftp://ftp.gnu.org/gnu/grub

or any of its mirrors. The file will be named grub-version.tar.gz. The current version is 2.12, so the file you should grab is:

 ftp://ftp.gnu.org/gnu/grub/grub-2.12.tar.gz

To unbundle GRUB use the instruction:

zcat grub-2.12.tar.gz | tar xvf -

which will create a directory called ‘grub-2.12’ with all the sources. You can look at the file ‘INSTALL’ for detailed instructions on how to build and install GRUB, but you should be able to just do:

cd grub-2.12
./configure
make install
 Also, the latest version is available using Git. See http://www.gnu.org/software/grub/grub-download.html

for more information.

Appendix B Reporting bugs

These are the guideline for how to report bugs. Take a look at this list below before you submit bugs:

1. Before getting unsettled, read this manual through and through. Also, see the GNU

GRUB FAQ (http://www.gnu.org/software/grub/grub-faq.html).

2. Always mention the information on your GRUB. The version number and the configuration are quite important. If you build it yourself, write the options specified to the configure script and your operating system, including the versions of gcc and binutils.

3. If you have trouble with the installation, inform us of how you installed GRUB. Don’t omit error messages, if any. Just ‘GRUB hangs up when it boots’ is not enough. The information on your hardware is also essential. These are especially important: the geometries and the partition tables of your hard disk drives and your BIOS.

4. If GRUB cannot boot your operating system, write down everything you see on the

screen. Don’t paraphrase them, like ‘The foo OS crashes with GRUB, even though
it can boot with the bar boot loader just fine’. Mention the commands you ex-

ecuted, the messages printed by them, and information on your operating system including the version number.

5. Explain what you wanted to do. It is very useful to know your purpose and your wish, and how GRUB didn’t satisfy you.

6. If you can investigate the problem yourself, please do. That will give you and us much more information on the problem. Attaching a patch is even better. When you attach a patch, make the patch in unified diff format, and write ChangeLog entries. But, even when you make a patch, don’t forget to explain the problem, so that we can understand what your patch is for.

7. Write down anything that you think might be related. Please understand that we often need to reproduce the same problem you encountered in our environment. So your information should be sufficient for us to do the same thing—Don’t forget that we cannot see your computer directly. If you are not sure whether to state a fact or leave it out, state it! Reporting too many things is much better than omitting something important.

If you follow the guideline above, submit a report to the Bug Tracking System

(http://savannah.gnu.org/bugs/?group=grub). Alternatively, you can submit a report

via electronic mail to bug-grub@gnu.org, but we strongly recommend that you use the Bug Tracking System, because e-mail can be passed over easily.

Once we get your report, we will try to fix the bugs.

Appendix C Where GRUB will go

GRUB 2 is now quite stable and used in many production systems. We are currently working towards a 2.0 release.

If you are interested in the development of GRUB 2, take a look at the homepage

(http://www.gnu.org/software/grub/grub.html).

Appendix D Copying This Manual

D.1 GNU Free Documentation License

Version 1.2, November 2002

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000,2001,2002 Free Software Foundation, Inc. 51 Franklin St, Fifth Floor, Boston, MA 02110-1301, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

D.1.1 ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.2
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with...Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Index 149

Index

[

[ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

A

acpi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 authenticate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

B

background_color. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 background_image. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 badram . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 blocklist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 boot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

C

cat. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 chainloader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 clear . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 cmosclean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 cmostest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 CMOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 cmp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 configfile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 cpuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 crc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 cryptomount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 cutmem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

D

date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 devicetree. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 distrust . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 drivemap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

E

echo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 efitextmode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 export . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

F

false . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 FDL, GNU Free Documentation License . . . . . . 141 fwsetup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

G

gdbinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 gettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 gptsync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

H

halt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 hashsum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 hexdump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

I

initrd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 initrd16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 insmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

K

keystatus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

L

linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 linux16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 list_env . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 list_trusted . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 load_env . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 loadfont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 loopback . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 ls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 lsfonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 lsmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

M

md5sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 menuentry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 multiboot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

N

nativedisk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 net_add_addr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 net_add_dns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 net_add_route . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 net_bootp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 net_del_addr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 net_del_dns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 net_del_route . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 net_dhcp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 net_get_dhcp_option . . . . . . . . . . . . . . . . . . . . . . . . 103 net_ipv6_autoconf . . . . . . . . . . . . . . . . . . . . . . . . . . 103 net_ls_addr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

150 GNU GRUB Manual 2.12

net_ls_cards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 net_ls_dns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 net_ls_routes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 net_nslookup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 net_set_vlan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 normal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 normal_exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

P

parttool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 password . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 password_pbkdf2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 plainmount. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 play . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 probe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

R

rdmsr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 reboot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 rmmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

S

save_env . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 sendkey . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 serial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 sha1sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 sha256sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 sha512sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 sleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 smbios . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 submenu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

T

terminal_input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 terminal_output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 terminfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 true . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 trust . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

U

unset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

V

verify_detached. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 videoinfo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

W

wrmsr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

X

xen_hypervisor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 xen_module. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76